### v1

# PLM-Based Discrete Diffusion Language Models with Entropy-Adaptive Gibbs Sampling

Abstract
Recently, discrete diffusion language models have demonstrated promising results in NLP. However, there has been limited research on integrating Pretrained Language Models (PLMs) into discrete diffusion models, resulting in underwhelming performance in downstream NLP generation tasks. This integration is particularly challenging because of the discrepancy between step-wise denoising strategy of diffusion models and single-step mask prediction approach of MLM-based PLMs. In this paper, we introduce Diffusion-EAGS, a novel approach that effectively integrates PLMs with the diffusion models. Furthermore, as it is challenging for PLMs to determine where to apply denoising during the diffusion process, we integrate an entropy tracking module to assist them. Finally, we propose entropy-based noise scheduling in the forward process to improve the effectiveness of entropy-adaptive sampling throughout the generation phase. Experimental results show that Diffusion-EAGS outperforms existing diffusion baselines in downstream generation tasks, achieving high text quality and diversity with precise token-level control. We also show that our model is capable of adapting to bilingual and low-resource settings, which are common in real-world applications.
PLM-Based Discrete Diffusion Language Models with Entropy-Adaptive Gibbs Sampling
[…]
1 Introduction
Type Dataset OpenWebText Gokaslan and Cohen (2019) RocStories Mostafazadeh et al. (2016) Deontology Hendrycks et al. (2023) Question Generation Dhingra et al. (2017) QQP Wang et al. (2017) ALMA Xu et al. (2024a) ParaDetox Logacheva et al. (2022)
Open-ended Generation ? $\checkmark$ $\checkmark$ $\triangle$ $\checkmark$ $\times$ $\times$ $\times$
Conditional Generation ? $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Context Provided ? - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Content Provided ? - $\times$ $\triangle$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Format Provided ? - - $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Table 1: Categorization of NLP tasks based on constraint aspects. $\checkmark$ indicates full support, $\times$ indicates no support, and $\triangle$ indicates partial or limited support.
As diffusion models significantly enhance the quality and diversity of generated outputs in continuous domains such as images and audio (Song et al., 2021b), recent research has increasingly applied diffusion models to NLP  (Li et al., 2022; Gong et al., 2023a; Yuan et al., 2023; He et al., 2023).
To integrate the diffusion process with traditional language models, research has predominantly followed two approaches: Continuous Diffusion Language Model (CDLM) (Yuan et al., 2023; Lovelace et al., 2023; Chen et al., 2023), which embed text in a continuous latent space, and Discrete Diffusion Language Model (DDLM) (He et al., 2023; Lou et al., 2024; Zhou et al., 2024; Shi et al., 2024; Sahoo et al., 2024; Zheng et al., 2024), which operate directly on the vocab space. Recent DDLMs adeptly reflect the highly structured nature of language; this property enables sequence control in natural language, generally resulting in high performance in NLP fields such as unconditional generation and open-ended generation  (Lou et al., 2024). However, for DDLMs to be widely adopted across various NLP domains, particularly in tasks such as close-ended conditional generation — hereafter referred to as dataset-guided generation — existing DDLMs still exhibit shortcomings in their performance. Our experiments suggest that models such as SEDD falls short in such tasks, potentially limiting their practical applications.
Better performance can be achieved by integrating Pretrained Language Models (PLMs) through the use of an encoder-based PLM as the initial setup for embedding functions, leveraging its capability to capture overall semantics (He et al., 2023). However, integrating PLMs into diffusion models is non-trivial as PLMs typically predict masked elements in a single step, whereas diffusion models require step-wise denoising based on the overall semantics of each timestep sequence, and such gap yields limited results. Therefore, we need to consider such inconsistency to effectively adopt PLMs into DDLMs by a new methodology.
In this paper, we introduce Diffusion-EAGS, a novel approach that effectively integrates Mask Language Model (MLM)-based PLMs with DDLMs for dataset-guided generation. To address the gap between the step-wise nature of diffusion models and the one-step prediction strategy of PLMs, we begin by considering MLM-based PLMs as the denoising function at each step of the diffusion process. Since identifying where to apply denoising during the diffusion process is challenging for PLMs, we incorporate an entropy tracking module to support their operation. For constructing entropy tracking module, we interpret each denoising step as a constrained Markov Random Field (cMRF), which enables us to adopt adaptive Gibbs sampling, which then facilitates leveraging entropy information from conditional sequences. To train DDLMs to effectively exploit an entropy tracking module, we propose a forward process based on an entropy-based noising scheduling during training phase. Specifically, our entropy-based noising scheduling noises lower entropy tokens first, thereby learning to progressively generate sequences guided by the entropy from entropy tracking module, which enhances the quality of the generated text.
Experimental results demonstrate that Diffusion-EAGS achieves outstanding performance compared to existing autoregressive model (GPT-2), CDLMs, and DDLMs across various dataset-guided generation tasks. We demonstrate that both PLMs and entropy-based Gibbs sampling contribute to higher performance in our ablation study. Furthermore, our model exhibits higher diversity in certain tasks compared to LLMs and can facilitate token-level generation, indicating the potential applicability of our model across a wide range of dataset-guided generation tasks. In addition, we validate that our model can adapt to both bilingual and low-resource settings, which are frequently encountered in practical applications.
2 Task Setting
2.1 Dataset-Guided Generation
In sequence-to-sequence (seq2seq) conditional generation tasks, the level of constraint imposed by the dataset plays a critical role in shaping the generation process. As shown in Table 1, constraints can be categorized into three levels: the provision of context alone, the provision of specific content that should be covered in the target sequence, and the provision of content format. Each level represents an increasing degree of constraint in the task requirements. The goal is to develop a model capable of generalizing across a wide range of downstream tasks, rather than being specialized for a specific task, while adhering to the constraints of the given dataset—a task we refer to as Dataset-Guided Generation.
2.2 DDLM with BERT
Diffusion Models The process of diffusion models consists of two main stages: the forward process, where noise is added to the input over multiple steps, and the reverse process, where noise is removed to restore the information. There are two types: Continuous and Discrete diffusion models. Discrete diffusion models operate the diffusion process across categorical random variables Hoogeboom et al. (2021). As discrete diffusion models show promising results in recent studies Venkatraman et al. (2024); Sahoo et al. (2024); Lou et al. (2024), we set up our language models grounded on discrete diffusion language models.
BERT as constrained Markov Random Field (cMRF) As the pretraining paradigm has significantly contributed to the success of language models, the integration of PLMs is essential in the development of language models. Nevertheless, adopting PLMs to DDLMs is a complex task, as DDLMs require a step-wise denoising strategy, while MLM-based PLMs are trained to predict masked elements in a single step. To effectively integrate PLMs into DDLMs, it is necessary to resolve such a discrepancy. Therefore, we begin by interpreting the intrinsic characteristics of MLMs as Markov Random Field (MRF), as demonstrated in Wang and Cho (2019). Details are in Appendix A.
Figure 1: Potential Function Difference of MLM based pre-trained BERT generation and RocStories dataset-guided BERT generation
However, dataset guided generation occupy different fields in terms of randomness, compared to open-ended generation. In Figure 1, we can observe an increase of potentials for same sequence X("The quick brown fox jumps over the lazy dog.") on RocStories. That is, it is necessary to investigate whether the dataset-constrained space is still subject to MRF.
Given a sequence of fully connected random variables $X=(x_{1},x_{2},\ldots,x_{L})$ with a set of observed variables $Y=(y_{1},y_{2},\ldots,y_{K})$, $f_{\theta}(X\backslash\{x_{l}\};Y)$ can be represented by logits from MLM given the sequence $X\backslash\{x_{l}\}$ and the observed variables $Y$. Then, similarly to Equation 2, log-potential can be calculated as:
$\log\phi_{l}(X;Y)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\};Y),&\text{if }\text{[M]}\notin X_{l-1:l+1}\\ \\ 0,&\text{otherwise}\end{cases}$
By this, we can define the potential function for the clique as:
$\phi(X;Y)=\exp\left(\sum_{l=1}^{L}\log\phi_{l}(X;Y)\right)$
Then, the constrained probability of sequence X is:
$p_{\theta}(X;Y)=\frac{1}{Z(Y,\theta)}\prod_{l=1}^{L}\phi_{l}(X;Y)$
Using the softmax function, the distribution of a variable $x_{l}$ given the observed variables $Y$ is:
$p(x_{l}|X\backslash\{x_{l}\},Y)=\frac{\exp(1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\};Y))}{\sum_{m=1}^{M}\exp(1h(m)^{T}f_{\theta}(X\backslash\{x_{l}\};Y))}$ (1)
This formulation implies that the dataset-constrained generation process in BERT can still be interpreted as a MRF with constraint Y, where the log-potential functions can be derived from the output logits of MLM.
3 Methodology
Figure 2: Overview of Diffusion-EAGS. During forward steps, We schedule the noise based on ENS using PLM. In generation steps, we denoise the sequences based on EAGS.
We propose Diffusion-EAGS that integrates PLM with DDLMs in dataset-guided generation, through Entropy-based Adaptive Gibbs Sampling (EAGS) and Entropy-based Noise Scheduling (ENS). The overview of our process is in Figure 2.
3.1 Generation Process: Entropy-based Adaptive Gibbs Sampling (EAGS)
Building on the premise in Section 2.2, we propose EAGS to leverage PLM as the function of denoising process in diffusion models. As MLM-based PLMs can be interpreted as cMRF, we can apply Gibbs sampling in each denoising step in DDLM. By Gibbs sampling, we can sample each token $x_{l}$ from its constrained distribution $p(x_{l}|X\backslash\{x_{l}\};Y)$. Rather than performing Gibbs sampling, we employ EAGS, as the entropy of each variable indicates the amount of information at certain position in the sequence. By generating tokens in the order of decreasing entropy, we prioritize the generation of the least informative parts of the sequence. Such prioritized generation makes a more structured sequences, leveraging the already established syntactic context.
According to Equation 1, we can compute the conditional probability for each masked position. The entropy $H$ for a given token $x_{l}$ is derived as:
$\begin{split}&H(x_{l}|Y,f)=\\ &-\sum_{m=1}^{M}p(x_{l}=v_{m}|x_{l},Y,f)\log p(x_{l}=v_{m}|x_{l},Y,f)\end{split}$
With $H$, our approach for T step-generation process can be formalized as follows:
  1. 1.
    Entropy Calculation: Compute the entropy $H(x_{l}\mid Y,f)$ for each variable $x_{l}$.
  2. 2.
    Variable Selection: Select the variable $x_{l^{*}}$ with the highest entropy for sampling
    $l^{*}=\arg\max_{l}H(x_{l}\mid Y,f)$
  3. 3.
    Sampling: Sample $x_{l^{*}}$ from its conditional distribution $p(x_{l^{*}}\mid Y,f^{*})$.
  4. 4.
    Update: Update the conditional distributions and entropy for the affected variables.
  5. 5.
    Iteration: Repeat Steps 1 through 4 until t=T.
3.2 Forward Process: Entropy-based Noise Scheduling (ENS)
To enhance the facilitation of Adaptive Gibbs Sampling in the generation process, it is essential to describe a similar process in the training phase. Therefore, we schedule the forward process of diffusion training based on the entropy $H(x_{l})$ of positions $x_{l}$ with the input sequence $[Y|X]$. Specifically, the position-wise entropy is calculated by the PLM that shares a similar language base with diffusion-trained model. Assuming the diffusion process progresses over $T$ steps, at each step $t$, we mask the $L/T$ number of positions with the lowest entropy from the set $\{x_{1},\ldots,x_{L}\}$. The masking process at step $t$ in position $l$ is described by the denoising matrix $Q_{tl}$.
$Q_{tl}=\begin{bmatrix}q_{11}&0&\cdots&0&q_{1,M}\\ 0&q_{22}&\cdots&0&q_{2,M}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&q_{M-1,M-1}&q_{M-1,M}\\ 0&0&\cdots&0&q_{MM}\\ \end{bmatrix}$
, where
$q_{mn}=\begin{cases}q_{mm}=1&\text{if }x_{l}\notin\text{MIN}([H(x_{1}),\cdots,H(x_{L})],\frac{L}{T})\\ q_{mM}=1&\text{if }x_{l}\in\text{MIN}([H(x_{1}),\cdots,H(x_{L})],\frac{L}{T})\\ 0&\text{otherwise}\end{cases}$
This entropy-based noise scheduling approach ensures that the forward process in diffusion training closely mirrors the generation process, thereby enhancing the effectiveness of Adaptive Gibbs Sampling in language generation.
3.3 Training & Inference
Training Process
Input : Context Y and Dataset D
for $BatchStep=0..N$  do
       $x\sim D$
       $\text{t}\sim\text{Randint}(0,T)$
       $f\leftarrow PLM(x|Y)$
       $\mathcal{H}\leftarrow H(x|Y,f)$
       $x^{t}\leftarrow\text{Forward}(x_{0},\mathcal{H},t)$
       $x^{t+1}\leftarrow\text{Forward}(x_{0},\mathcal{H},t+1)$
       $L_{s}=-\sum_{i}(q(x^{t}_{i})\mid x_{t+1})\log(p_{\theta}(x^{t}_{i})\mid x_{t+1})$
end for
Generation Process
Input : Sequence Length L, Total Timestep T,Trained Model M, Mask Sequence Generator $G_{M}$, and Context Y
[…]
       if  $t=T$  then // Initialize $x^{T}$
             $x^{T}\leftarrow G_{M}(L,Y)$
       else  // Denoise step
             $f\leftarrow M(x^{t},Y)$
[…]
             $x^{t-1}=\text{Sampling}(x^{t},l^{*},Y)$
       end if
end for
Algorithm 1 Training & Generation
Our whole process is in Algorithm 1. Distinct from the prevailing methodologies in diffusion models as described by Ho et al. (2020a) and Austin et al. (2023), we do not employ the PLM parameterization strategy $\widetilde{p}_{\theta}(\widetilde{z}_{0}|z_{t},t)$, which preserves the original semantic embedding spaces during the training phase. We empirically find that PLM parameterization restricts the diversity of generated responses. Instead, to ensure the completeness of sentences, we implement Minimum Bayes Risk (MBR) decoding in the final generation step. We follow the traditional diffusion loss in Equation (Ho et al., 2020b) with changing MSE to Cross Entrpy Loss.
4 Experiments
4.1 Tasks and Datasets
Question Generation We use the Quasar-T dataset (Dhingra et al., 2017) to generate questions from passages and answers. Paraphrase The QQP dataset (Wang et al., 2017) assesses paraphrase detection by treating one question as a paraphrase of another. Social Datasets The Paradetox dataset (Logacheva et al., 2022) focuses on removing profanities. The Deontology from ETHICS dataset (Hendrycks et al., 2023) evaluates ethical judgment based on scenarios. Open-ended Generation The RocStories (Mostafazadeh et al., 2016) dataset is used to generate coherent story continuations from a given context. Detailed explanations are available in Appendix B and Dataset statistics are shown in Table 8.
4.2 Baselines
Since our objective is centered on generation tasks, we compare Diffusion-EAGS with three categories of baselines with the model similar to the size of RoBERTa-base: AutoRegressive (AR) model, CDLM, and DDLM. For AR model (Vaswani et al., 2023), we employ GPT-2 (Radford et al., 2019), which is renowned for its proficiency in generation tasks. For CDLMs, our baseline includes DiffuSeq (Gong et al., 2023a) and DINOISER (Ye et al., 2024). DiffuSeq is a diffusion model specifically designed for sequence-to-sequence applications, whereas DINOISER adaptively determines the range of sampled noise scales during training. For DDLMs, we utilize DiffusionBERT (He et al., 2022) , a state-of-the-art model in the research series of DiffusionLM (Li et al., 2022), and SEDD (Lou et al., 2024), a state-of-the-art model of open-ended generation in diffusion language generation domain. For SEDD, we download the pretrained version and fine-tune on it. Experimental results on LLMs can be found in Appendix C.
4.3 Metrics
Quality metrics To measure the quality of the generated texts, we use Perplexity based-on GPT-2 Large and GPT-2 XL, SOME (Yoshimura et al., 2020), grammar metric based on the neural net, LLM-c (Lin and Chen, 2023) to measure the plausibility of the narratives, LLM-t (Koh et al., 2024) to measure toxicity, and MAUVE Pillutla et al. (2021), measuring the gap between text distributions despite divergence.
Diversity Metrics Traditional diversity metrics Self-BLEU Zhu et al. (2018) and distinct-n Li et al. (2015) are employed to evaluate the generated texts. We also adopt Vendi Score (VS) Friedman and Dieng (2023), an interpretable diversity metric, which quantifies the effective number of unique samples in a given set. Both the n-gram and embedding variations are utilized, where embedding VS interprets diversity of semantics.
4.4 Experimental Details
We employ roberta-base as encoder-based PLM. The learning rate is set at 5e-4, and a naive categorical sampling strategy is adopted. Adapted to data statistics, the maximum lengths for QG, QQP, and Paradetox is set to 64, while for Deontology set to 48. Furthermore, in accordance with the minimum construction length, the number of steps is configured to 5 and 20 size of MBR. We use 1 A100 GPU, and the batch size is set to 256.
QG
Text Quality Diversity
Model Step PPL $\downarrow$ MAUVE $\uparrow$ SOME $\uparrow$ VS(ngram) $\uparrow$ VS(emb) $\uparrow$ self-bleu $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$
GPT-2 1 124.8 0.141 0.759 4.564 3.130 0.176 0.210 0.629
DiffuSeq 20 395.0 0.149 0.730 1.555 1.274 0.901 0.170 0.564
DINOISER 2000 155.9 0.159 0.776 1.396 1.121 0.944 0.166 0.553
DiffusionBert 2000 513.6 0.150 0.712 3.040 2.209 0.566 0.392 0.759
Diffusion-EAGS 5 80.7 0.121 0.782 4.646 3.538 0.152 0.403 0.798
QQP
Model Step PPL $\downarrow$ MAUVE $\uparrow$ SOME $\uparrow$ VS(ngram) $\uparrow$ VS(emb) $\uparrow$ self-bleu $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$
GPT-2 1 66.270 0.112 0.754 3.886 2.566 0.423 0.344 0.787
DiffuSeq 2000 124.247 0.00674 0.709 1.927 1.242 0.813 0.226 0.543
DINOISER 20 79.742 0.0042 0.821 1.421 1.126 0.935 0.264 0.542
DiffusionBert 2000 500.959 0.0709 0.618 4.489 2.836 0.196 0.321 0.761
Diffusion-EAGS 5 48.106 0.683 0.824 4.006 2.390 0.338 0.421 0.832
Table 2: QG & QQP Generation
ParaDetox
Model Step PPL $\downarrow$ MAUVE $\uparrow$ SOME $\uparrow$ VS(ngram) $\uparrow$ VS(emb) $\uparrow$ self-bleu $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$
GPT-2 1 389.1 0.503 0.717 3.925 2.640 0.429 0.312 0.748
DiffuSeq 2000 $\geq 1k$ 0.683 0.703 2.059 1.465 0.841 0.410 0.820
DINOISER 20 124.8 0.255 0.767 2.287 2.174 0.981 0.211 0.486
DiffusionBert 2000 775.9 0.737 0.716 3.101 2.058 0.599 0.424 0.826
SEDD 1024 $\geq 1k$ NA 0.664 4.746 4.063 0.119 0.451 0.846
Diffusion-EAGS 5 69.5 0.773 0.796 4.755 3.659 0.126 0.475 0.834
Deontology
Step PPL $\downarrow$ MAUVE $\uparrow$ SOME $\uparrow$ VS(ngram) $\uparrow$ VS(emb) $\uparrow$ self-bleu $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$
GPT-2 1 92.0 0.131 0.860 3.665 3.126 0.425 0.474 0.874
DiffuSeq 2000 352.8 0.005 0.703 2.273 1.915 0.753 0.267 0.745
DINOISER 20 131.3 0.008 0.740 2.287 2.174 0.824 0.309 0.713
DiffusionBert 2000 295.5 0.306 0.787 4.258 3.458 0.229 0.445 0.849
Diffusion-EAGS 5 55.1 0.412 0.835 4.898 4.009 0.056 0.418 0.806
Table 3: Social Generation
5 Results
As shown in Table 234, our model consistently exhibits exceptional performance in terms of text quality while simultaneously maintaining diversity when compared to baseline models. As illustrated in Table 2, Diffusion-EAGS generates the responses with the highest PPL score for QG, and highest MAUVE and PPL score for QQP. In Table 3 Paradetox, our model demonstrates superior performance across all evaluated metrics. Such phenomenon represents that our model based on encoder-based PLMs show robustness on diverse perturbations from daily dialogues. When PPL exceeds 1,000, the model is considered to have failed in generating natural sequences, and is thus represented in gray color. In the context of Deontology, our model exceeds the baseline PPL and MAUVE scores, whereas SOME score represent the sufficient quality of text with the highest diversity score of 4.755. Additionally, in Table 4 where semantic consistency is crucial, our model significantly outperforms the original dataset’s PPL by 23 points by leveraging PLMs while maintaining a high diversity score of 4.837.
5.1 Ours vs AR model
Diffusion-EAGS demonstrates high level of text quality surpassing that of GPT-2 in Table 2 in text quality. Notably, our model effectively reflects the pattern and style of the dataset, attributed to the capability of encoder-based PLMs. Specifically, as reported in Table 3-ParaDetox, our model excels at capturing semantic information with the highest MAUVE score of 0.733, despite ParaDetox being a challenging colloquial dataset including slang, numerous abbreviations, and various perturbations. In addition, our model demonstrate robustness to such perturbations with 69.5 PPL, while GPT-2 shows a lower performance of 389.1 PPL.
As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 23, and 4. These results underscore our model’s efficacy in generating diverse responses.
Model PPL $\downarrow$ SOME $\uparrow$ LLM-c $\uparrow$ VS(ngram) $\uparrow$ self-bleu $\downarrow$
Original Data 100.6 0.895 1
GPT-2 88.5 0.856 0.88 4.722 0.124
DiffusionBert 318.2 0.783 0.72 4.735 0.088
SEDD 273.2 0.827 0.59 4.859 0.044
Diffusion-EAGS 67.3 0.844 0.87 4.837 0.058
Table 4: ROC Generation
Specifically, in the context of story generation as shown in Table 4, our model not only demonstrates effectiveness in enhancing both the quality and diversity of the text but also maintains coherence with the contextual storyline. This is further supported by the high LLM-c scores, which confirm the plausibility of the narratives produced by our model.
5.2 Ours vs CDLMs
Notably, CDLMs demonstrate a noticeable deficiency in diversity. In contrast, our model excels at producing significantly more diverse sequences. This enhanced performance can be attributed to the integration of PLM into our framework, which evidently enriches the diversity of the generated outputs. In addition, our model generally achieves higher text quality scores than CDLMs across all experiments. Diffusion-EAGS not only surpasses Diffuseq and DINOISER in PPL and SOME but also achieves a MAUVE score improvement of over 0.4, thus indicating that the generated responses accurately reflect the dataset while preserving grammatical integrity. Furthermore, our models require only a few steps, resulting in a 40% reduction of costs with higher quality and diversity.
5.3 Ours vs DDLMs
Contrary to CDLMs, DiffusionBERT exhibits limitations in text quality, as shown in Table 234. On the other hand, our model achieves significantly higher scores across all quality metrics. This observation suggests that the naive application of BERT within the diffusion process fails to fully harness the capabilities of PLMs. SEDD, an open-ended generation model shows low performance under text generation in highly constrained tasks such as ParaDetox. Even in open-ended generation settings of Table 4, Diffusion-EAGS shows higher text quality and semantic scores. Additionally, our model does not require more than 5 diffusion steps, as it can adaptively recover the absorbing state through an entropy-based generation approach. Above results further exemplify the high performance and the time efficiency of our model.
6 Analysis
6.1 Ablation Study
To explore the effectiveness of our model’s components, we conduct ablation studies focusing on three key elements: Entropy Adaptation, stepwise Gibbs Sampling, and PLM in Table 5.
The omission of EAGS initially results in a significant decline in performance, a noticeable decrease in text quality and diversity.
Dataset PPL MAUVE SOME VS(ngram) VS(emb)
Diffusion-EAGS Deont 55.1 0.412 0.835 4.898 4.009
Roc 67.3 0.87 0.844 4.837 3.999
w/o EAGS Deont 667.9 0.022 0.617 4.767 3.928
Roc 1084.9 0.035 0.613 4.874 3.957
w/o Gibbs Sampling Deont 1426.7 0.011 0.584 2.378 1.923
Roc 1293.1 0.010 0.534 1.531 1.338
w/o PLM Deont $\geq$2K 0.005 0.645 4.758 3.402
Roc $\geq$2K 0.004 0.604 4.315 2.994
Table 5: Ablation Study
We find that EAGS significantly contributes to a gradual entropy decrease, as shown in Appendix H. These observations highlight the critical role of selecting the sampling position based on given information in sequence generation. Additionally, excluding the use of the diffusion generation process, without the stepwise sampling, leads to a drastic reduction in overall performance, with PPL increasing by more than 1000 points and diversity scores dropping below 2. Consequently, our cMRF approach, integrated with the diffusion model, proves indispensable in aiding the PLM in effectively generating sequences. Furthermore, it is evident that without the incorporation of PLMs, the generation of natural sequences is unachievable.
6.2 Case Study: Keyword Based Generation
Our model operating within discrete space enables us to manipulate the output sequences using explicit instructions. To further explore this capability, we conduct the generation of sequences based on keywords positioned in the middle and at the end of masked sequences. Initially, we provide the same contextual input while varying the keywords. In the masked states, we randomly select positions, replacing them with the specified keywords. The results in Table 7 demonstrate that the generated sequences seamlessly integrate the keywords with context-specific semantics.
These findings show that our model can effectively integrate specific conditional information in an interpretable manner. Experimental results also indicates its substantial potential and suitability for diverse applications where direct controllability is crucial, such as in story generation.
6.3 Bilinguality & Low Resource
Model SacreBLEU COMET XCOMET
DisCo (Kasai et al., 2020)
w/ Easy-First 3.2806 0.2447 0.2414
w/ Mask-Predict 3.2862 0.2444 0.2414
DisCo-m
w/ Easy-First 3.7423 0.2468 0.2122
w/ Mask-Predict 3.7748 0.2466 0.2119
Diffuseq-v2 (Gong et al., 2023b) 1.90 0.3242 0.2628
SEDD
w/ from scratch 0.14 0.2375 0.2035
w/ pretrained 0.25 0.2504 0.2076
DiffusionEAGS-NLLB 20.9297 0.5720 0.6629
NLLB-naive-600M 4.1827 0.6134 0.7818
mBART-50-FT 19.6536 0.7576 0.8748
Table 6: Translation Results
Context Jake was playing with his toys. He accidentally broke his favorite one.
He cried a lot over it. His parents decided to replace it for him.
Keyword not stop Jake just could not stop crying.
Jake feel It made Jake feel So much better.
would enjoy Jake said he would enjoy the new toy
Context Neil was in Sofia, Bulgaria. He was enjoying a trip backpacking through Europe.
… He thought the food and culture in Sofia were the best.
Keyword Bulgaria! Things were looking great in Bulgaria!
Context Karen wanted to go on a trip to France. She started doing research on the trip.
She decided to book a week long trip. She left the next day for her trip.
Keyword her trip She spent almost a week there during her trip.
Table 7: Keyword-based Generation
[…]
We establish baselines using non-autoregressive (NAR) transformer models, such as Mask-and-Predict (Ghazvininejad et al., 2019a) and Easy-First (Kasai et al., 2020) - which bear resemblance to discrete diffusion models - diffusion models such as Diffuseq-v2 and SEDD, traditional translation models such as mBART-50 (Tang et al., 2020) and NLLB. For evaluation metrics, we utilize sacreBLEU (Post, 2018) and neural-net scores such as COMET (Rei et al., 2020) and XCOMET (Guerreiro et al., 2023).
Table 6 shows that predicting the target sequence without leveraging a multilingual PLM proves to be challenging. All diffusion baseline models struggle to produce correct outputs. For example, the pretrained SEDD model fails to effectively leverage conditional information, even after finetuning on the training datasets, consistent with the limitation observed in Section 5.3. Similar challenges arise in NAR transformer baselines. Despite constructing the vocabulary using the pretrained mBART-50 model (DisCo-m), the underlying issues remain. On the other hand, our proposed model, by incorporating a PLM, demonstrates promising results.
Interestingly, the output of the pretrained NLLB model (NLLB-naive-600M, not finetuned) reveal that neural network-based metrics are susceptible to the interference problem, specifically translated by other languages, even though we provide the language specific token. While such issues result in lower BLEU scores, COMET and XCOMET often interpret them as semantically coherent, indicating a potential direction for future work to improve translation evaluation metrics. Despite these phenomena, a performance gap between translation models and DDLM remains. This suggests that future research should address the semantic capabilities of diffusion models to help bridge this gap. Details are in Appendix G.
6.4 Diversity Saturation
Figure 3: Diversity graph with increasing generation numbers in ’Deontology’ dataset
Inspired by the observation that Diffusion-EAGS consistently excel in terms of diversity across all results, we delve further into the diversity capabilities of our model. We assess the diversity performance in dataset-guided generation compared to LLMs. We measure the VS for 5 to 100 generations under a single condition. Such experiment demonstrates the extent to which the model’s output diversity saturates, enabling a comparison of asymptotic diversity performance. The experiment is conducted on the ‘deontology’ dataset which allows high output diversity in its settings. Details of using LLMs are provided in Appendix E.
Figure 3 demonstrates that the diversity saturation graph for Diffusion-EAGS has a relatively steep slope, while GPT models saturate at lower values. The embedding VS of all GPT series saturates below 13. This indicates that the limitation of diversity is inherent to the architecture itself, rather than merely a factor of scale in the GPT series. In contrast, Diffusion-EAGS is capable of producing significantly more diverse textual outputs.
7 Related Works
Efforts to integrate generative flow models into sequence generation exploit the distribution shift from a source language to a target language through a series of invertible linear transformations (Ma et al., 2019; Zhang et al., 2024). However, as DDPM (Ho et al., 2020a) demonstrate the effectiveness of generating images, diffusion models have been a major topic of interest within the field of generative flow models (Song et al., 2021a, b). To apply such diffusion methodologies to NLP, there are two main streams - continuous diffusion models and discrete diffusion models.
Continuous diffusion models Diffusion-LM (Li et al., 2022) propose a method of mapping tokenized sequences to embedding dimensions guided by a pretrained classifier. DiffuSeq-v1, v2 (Gong et al., 2023a, b) apply partial noising techniques. The core of these diffusion methodologies lies in the addition and restoration of random noise to facilitate generation. However, Cold Diffusion (Bansal et al., 2022) argues that such operations do not necessarily have to be governed by stochastic randomness.
Discrete diffusion models Building on this rationale, D3PM (Austin et al., 2023) propose the discrete restoration-generation approach and DiffusionBERT He et al. (2022) adopt PLMs to DDLM. SEDDLou et al. (2024) proposes score entropy inspired by MLM loss. Recent works by Shi et al. (2024) and Sahoo et al. (2024) extend this idea to propose a simplified view of the discrete framework, and obtain better empirical results. Zheng et al. (2024) further closes the gap between discrete diffusion models and masked models. These researches make an improvement on the open ended generation task. Venkatraman et al. (2024) showcase that SEDD priors can be used to sample from a posterior distribution, namely text infilling.
8 Conclusions
In this work, we introduce Diffusion-EAGS, integrating encoder-based PLMs with diffusion models for dataset-guided generation. By leveraging step-wise diffusion models and the property of MLM, we demonstrate that MLM-based PLMs can serve as denoising function with Entropy-based Adaptive Gibbs Sampling. We also propose entropy-based noise scheduling during training for adaptive sampling. Experimental results show that Diffusion-EAGS outperforms existing baselines, achieving high diversity and precise token-level control. These results demonstrate its enhanced text generation quality and diversity, as well as wide applicability.
Limitations
While Diffusion-EAGS demonstrates significant improvements in conditional generation tasks, there are several limitations. First, the quality of the text generated by our model depends on PLMs such as BERT. Second, as our method is currently focused on text generation tasks, its applicability to text classification tasks, such as Named Entity Recognition and Part-of-Speech Tagging, remains unexplored. Future research could explore extending this method to other NLP tasks, as well as investigating the potential for reducing computational costs through more efficient entropy calculation techniques. Additionally, addressing the biases inherent in pretrained models remains a critical area of focus to ensure the robustness and fairness of generated outputs.
Ethical Statements
A notable advantage of Diffusion-EAGS is its efficient training process, which optimizes computational resources with high performance. This efficiency not only reduces the environmental impact of training PLMs but also makes advanced NLP technologies more accessible. Additionally, we have focused on output control and interpretability through the use of a discrete diffusion model. Consequently, our methodology contributes to the effective control of generated outputs in the future. Diffusion-EAGS’s strength in tasks involving conditional generation proves beneficial in the field of text generation following social guidelines, where outputs need to strongly depend on the conditions. Specifically, its high diversity performance enables active utilization in dynamically adapting text generations to meet specific social criteria.
Appendix A BERT as a MRF
Let $X=(x_{1},\ldots,x_{L})$ be a sequence of random variables $x_{l}$, each taking a value from a vocabulary $V=\{v_{1},\ldots,v_{M}\}$. These variables form a fully-connected graph, indicating mutual dependence.
To define the MRF, we focus on the full-graph clique potential:
$\phi(X)=\exp\left(\sum_{l=1}^{L}\log\phi_{l}(X)\right)$ (2)
where each log-potential $\phi_{l}(X)$ is:
$\log\phi_{l}(X)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\}),&\text{if }\text{[M]}\notin X_{l-1:l+1}\\ \\ 0,&\text{otherwise}\end{cases}$ (3)
Here, $f_{\theta}(X\backslash\{x_{l}\})$ is a function that maps the sequence $X\backslash\{x_{l}\}$ to a real-valued vector in $\mathbb{R}^{M}$, and $1h(x_{l})$ is a one-hot vector with $x_{l}$ set to 1. from this log potential, we can find probability of sequence X:
$p_{\theta}(X)=\frac{1}{Z(\theta)}\prod_{l=1}^{L}\phi_{l}(X)$
with the normalization constant $Z(\theta)$ defined as:
$Z(\theta)=\sum_{X^{\prime}}\prod_{l=1}^{L}\phi_{l}(X^{\prime})$
For a fixed $X\backslash\{x_{l}\}$, the conditional probability of $x_{l}$ is:
$p(x_{l}|X\backslash\{x_{l}\})=\frac{1}{Z(X\backslash\{x_{l}\})}\exp(1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\}))$
where $Z(X\backslash\{x_{l}\})$ is:
$Z(X\backslash\{x_{l}\})=\sum_{m=1}^{M}\exp(1h(m)^{T}f_{\theta}(X\backslash\{x_{l}\}))$
BERT uses pseudo log-likelihood learning to handle the intractability of $Z(\theta)$:
$\text{PLL}(\theta;D)=\frac{1}{|D|}\sum_{X\in D}\sum_{l=1}^{|X|}\log p(x_{l}|X\backslash\{x_{l}\})$
where $D$ is the training dataset.
Appendix B Dataset Explanations
Quasar-T QQP ParaDetox Deontology RocStories
input output input output input output input output input output
Max 63 244 104 98 35 35 24 31 76 57
Mean 14.574 31.157 13.947 13.956 15.135 13.035 13.039 12.548 42.189 13.307
Table 8: Dataset Statistics
QG The objective of Question Generation (QG) is to generate valid and fluent questions based on a given passage and a specified answer. We employ the Quasar-T dataset, introduced by Dhingra et al. (2017) in 2017, which comprises a substantial number of document-question pairs. These pairs necessitate the transformation of similar sentences into a single abstract question.
Paraphrase The objective of the Quora Question Pairs (QQP) (Wang et al., 2017) is to determine whether two questions are paraphrases of each other. We process the QQP dataset by treating one question as a paraphrase of another, a method commonly employed to assess the effectiveness of diffusion models.
Paradetox The objective of the Paradetox (Logacheva et al., 2022) is to delete the profanities in source sentence. It comprises of toxic and neutral utterances, curated from the Jigsaw, Reddit, and Twitter datasets.
Deontology The objective of Deontology (Hendrycks et al., 2023) is to to evaluate the capability of models to make ethical judgments from a deontological perspective. The dataset contains scenarios focusing on interpersonal dynamics and everyday occurrences.
Open-ended Generation We employ the RocStories dataset (Mostafazadeh et al., 2016) for open ended generation with narrative understanding tasks. This dataset contains short commonsense stories that require models to generate coherent and contextually relevant continuations. Each story comprises five sentences, where the task is to predict the fifth sentence given the first four. This setup evaluates the model’s ability to understand and generate narratives based on sequential context.
Appendix C Experimental Outputs
C.1 LLM Evaluation
Table 9: ParaDetox LLM Evalution (LLM-t)
ParaDetox Prompt
You are a specialized toxicity scoring machine that can only output integer scores.
Score the Context on a scale from 0 to 1,
where a scale of zero means "language is not perfectly toxic",
a scale of one means "language is extremely toxic".
Take a deep breath and think step-by-step.
Table 10: RocStories LLM Evalution (LLM-c)
RocStories Prompt
Scoring the naturalness in a integer scale between 0 and 1,
where a scale of zero means is not natural,
and a scale of one means natural.
Take a deep breath and think step-by-step.
C.2 Tables
We have appended the performance metrics of LLMs for reference. LLMs exhibit strong performance in tasks like QG, QQP, and Deontology, where datasets are more neat and contain fewer noises and abbreviations. Conversely, performance is lower on the Paradetox dataset, which includes a wide range of casual phrases and diverse perturbations. Provided that our model utilizes encoder-based PLMs, we observe that it demonstrates robustness to such perturbations compared to autoregressive models.
ParaDetox
PPL MAUVE SOME LLM-t VS(ngram) VS(emb) self-bleu distinct-1 distinct-2
GPT-2 389.147 0.503 0.717 0.02 3.925 2.640 0.429 0.312 0.748
GPT-3.5 104.375 0.175 0.888 0.074 3.098 1.915 0.652 0.390 0.835
GPT-4 78.979 0.125 0.879 0.18 3.214 1.906 0.592 0.412 0.841
DiffuSeq $\geq 1k$ 0.683 0.703 0.03 2.059 1.465 0.841 0.410 0.820
Diffusion-Bert 775.928 0.737 0.716 0.09 3.101 2.058 0.599 0.424 0.826
DINOISER 124.797 0.255 0.767 0.1 2.287 2.174 0.981 0.211 0.486
SEDD-small $\geq 1k$ NA 0.664 NA 4.746 4.063 0.119 0.451 0.846
Diffusion-EAGS 69.5 0.773 0.796 0.01 4.755 3.659 0.126 0.475 0.834
Table 11: ParaDetox Dataset Generation : We use LLM-evaluation approach for measuring toxicity, denoted as LLM-t
Deontology
PPL MAUVE SOME VS(ngram) VS(emb) self-bleu distinct-1 distinct-2
GPT-2 91.962 0.131 0.860 3.665 3.126 0.425 0.474 0.874
GPT-3.5 52.401 0.393 0.904 4.632 3.650 0.186 0.434 0.855
GPT-4 72.329 0.465 0.921 4.530 3.286 0.191 0.425 0.865
Table 12: Deontology Dataset Generation
QQP
PPL MAUVE SOME VS(ngram) VS(emb) self-bleu distinct-1 distinct-2
GPT-2 66.270 0.112 0.754 3.886 2.566 0.423 0.344 0.787
GPT-3.5 55.275 0.708 0.874 2.781 1.603 0.691 0.365 0.814
GPT-4 66.121 0.814 0.877 2.981 1.651 0.673 0.423 0.866
Table 13: QQP Dataset Generation
QG
PPL SOME MAUVE VS(ngram) VS(emb) self-bleu distinct-1 distinct-2
GPT-2 124.831 0.759 0.175 4.564 3.130 0.176 0.210 0.629
GPT-3.5 57.539 0.808 0.0184 3.174 2.090 0.612 0.282 0.739
GPT-4 75.435 0.822 0.0171 3.819 2.069 0.444 0.311 0.773
Table 14: QG Dataset Generation
Appendix D Well-Generated Output Examples
D.1 QG
Table 15: QG generation examples
QG Generation Output Examples of Diffusion-EAGS
Constraint $Y$ “Besides being able to hover in place, the hummingbird can also fly backwards.”
Output $X_{1}$ “What kind of bird can fly backwards?”
Output $X_{2}$ “Which bird is able to fly backwards?”
Constraint $Y$ “A marsupium or pouch is one of the features that characterise marsupials although not all have a permanent pouch and a few have none at all.”
Output $X_{1}$ “What is a pouch?”
Output $X_{2}$ “What is the smallest animal without a pouch.”
D.2 QQP
Table 16: QQP generation examples
QQP Generation Output Examples of Diffusion-EAGS
Constraint $Y$ “What are the ten best short stories written by Isaac Asimov?”
Output $X_{1}$ “What are some great most amazing stories written by Isaac Asimov?”
Output $X_{2}$ “What are the best known fiction and books of Isaac Asimov?”
Constraint $Y$ “Can we ever store energy produced in lightning?”
Output $X_{1}$ “How do we store heat energy from lightning?”
Output $X_{2}$ “How can you store energy from lightning?”
D.3 ParaDetox
Table 17: ParaDetox generation examples
ParaDetox Generation Output Examples of Diffusion-EAGS
Constraint $Y$ “this pathetic story just gets worse and worse.”
Output $X_{1}$ “this story is going to get worse due to his situation”
Output $X_{2}$ “this story continues to get worse.”
Constraint $Y$ “ fuck no!, there’s no justification for fgm.”
Output $X_{1}$ “there is no justification for it.”
Output $X_{2}$ “of course we cannot justify it.”
Table 18: As the dataset is designed to eliminate problematic words while preserving semantics, our model adheres to these guidelines.
D.4 Deontology
Table 19: Deontology generation examples
Deontology Generation Output Examples of Diffusion-EAGS
Constraint $Y$ “I am a doctor working in a hospital.”
Output $X_{1}$ “So I should know how my patients feel.”
Output $X_{2}$ “I am trained to diagnose people with complex illnesses.”
Constraint $Y$ “I am the owner of the apartment building.”
Output $X_{1}$ “I need to rent out the whole building.”
Output $X_{2}$ “So I have to rent it to others.”
D.5 RocStories
Table 20: RocStories generation examples
RocStories Generation Output Examples of Diffusion-EAGS
Constraint $Y$ “The man grew out his hair. He saw some gray hairs. He shaved his hair off. He bought some hair dye.”
Output $X_{1}$ “He wanted to look fresh and new.”
Output $X_{2}$ “His hair was dyed back to its original color.”
Constraint $Y$ “Jake was playing with his toys. He accidentally broke his favorite one. He cried a lot over it. His parents decided to replace it for him.”
Output $X_{1}$ “Jake was not very happy about it.”
Output $X_{2}$ “So he got a brand new one after all.”
Appendix E Details on Text Augmentation Using GPT models
E.1 GPT-3.5turbo ~ GPT-4-Omni
We prompt the GPT models to carry out dataset augmentation. To obtain quality responses that are similar to examples in the dataset, each generation is carried out in a 4-shot setting to leverage in-context learning, with the examples being randomly selected from the train split of the respective datasets. Furthermore, as Deshpande et al. (2023) illustrate that assigning a persona can affect the text output of LLMs to a considerable degree, and Zanella et al. (2024) show that assigning an appropriate persona can improve LLMs’ performance on the target task, albeit as automatic scorers in the anomaly detection domain, we assign the persona of a "dataset augmentation machine" to each of the LLMs in the input prompt. We observe that such persona assignment greatly lowered the number of times the LLM refused to provide a valid response when the input contain toxic content, which is relavant on toxicity datasets such as the Paradetox Dataset. This finding is in-line with the results of Deshpande et al. (2023). GPT-3.5-Turbo rejects 6.8% of the inputs on the Paradetox dataset, while GPT4, GPT4-Turbo, and GPT-4-Omni rejected none. To obtain diverse responses, all generated responses were obtained with the temperature set to 1.
The prompt template is as follows:
You are a dataset augmentation machine. Given the condition text, generate the target text.
CONDITION: <example condition 1>
TARGET: <example target(response) 1>
CONDITION: <example condition 2>
TARGET: <example target(response) 2>
CONDITION: <example condition 3>
TARGET: <example target(response) 3>
CONDITION: <example condition 4>
TARGET: <example target(response) 4>
CONDITION: <input condition>
TARGET:
Appendix F Details on CDLMs
F.1 Experimental Details
For the case of Diffuseq and Dinoiser Ye et al. (2024), we followed the official repositories to reproduce the results. Results were sampled multiple times with different seeds to evaluate the diversity. Some deviations are as follows. For max-length, we choose 64 for Paradetox, QG, and QQP, and 48 for Deontology. The values were chosen after examining the training set. As for batch size, we followed the original repositories if the parameter was provided. If not, the batch size was chosen using linear interpolation with the size of the training set. Note that unlike other benchmarks, we experimented with Diffuseq-v2 (Gong et al., 2023b) in translation task for a broader comparison with existing baselines.
F.2 Results Interpretations
Examining the results of Diffuseq, it is evident that the grammar score is comparatively lower than that of other models. This outcome is expected, as the outputs from Diffuseq frequently display inaccurate sentence structures, including duplications of words or phrases. Conversely, the outputs from Dinoiser achieve moderate grammar scores but show limited diversity. This finding, coupled with our additional experiments concerning the beam size during Dinoiser generation, suggests that Dinoiser’s performance predominantly relies on memorization.
Appendix G Details on Translation Results
G.1 Comparison Between Easy-First and Our Proposed Method
Discrete diffusion can be said to inherit ideas from NAR inference algorithm Mask-PredictGhazvininejad et al. (2019b) and Easy-First Kasai et al. (2020). Easy-First, especially, and our method are similar in how the probabilities of the predicted tokens are used for non-autoregressive inference.
The difference between the Easy-First and our method as follows: Easy-First, in each iteration, predicts tokens in each position given previous predictions on the easier positions. There is no strict unmasking process. This is in contrast to our model, which focuses on denoising masked states in accordance with the forward noising trajectory. Furthermore, the inference algorithm, as implemented in the original works Kasai et al. (2020) do not facilitate the integration of PLMs, which is a crucial component in modern NLP applications. We also bridge the gap between the diffusion framework and language modeling, a direction that have only recently began to gain traction within the research community.
We provide results on Easy-First, as well as Mask-Predict Ghazvininejad et al. (2019b) on the original DisCo architecture implementation as baselines on translations tasks in Table 6 to further elucidate the difference through empirical results.
G.2 Experimental Details
NAR Transformer & CMLM We utilized the official repository to produce obtain the results, with the default architecture, optimization, and inference configurations. We report the performance of the DisCo transformer on both the Mask-Predict and the Easy-First inference algorithms.
Diffuseq-v2 For Diffuseq-v2, we employ the vocabs of mBERT and choose 128 as max length for ende translation. Other settings are identical as in Appendix F.1.
SEDD The SEDDLou et al. (2024) model, originally designed for open-ended text generation, was adapted in this study to facilitate dataset-guided generation. To align the model’s architecture with the specific requirements of the structured dataset, several modifications were implemented in both hyperparameters and preprocessing protocols. Specifically, the input and output token lengths were constrained to a range of 64 to 128 tokens, ensuring a more appropriate fit to the dataset’s structural characteristics. Moreover, distinct special tokens were introduced to clearly differentiate between input and output sequences, thereby enhancing the model’s ability to distinguish between these components during training. Individual data entries were further demarcated by an EOS token to delineate discrete sequences within the training process.
mBART-50 & Distilled-NLLB-600M For mBART, we finetune from the checkpoint "facebook/mbart-large-50", with batch size 8, max sequence length set to 512, and with no gradient accumulation. For NLLB, we set the source language to $eng\_Latn$ and the target language to $deu\_Latn$. We employ the model "facebook/nllb-200-distilled-600M" with a batch size of 16, gradient accumulation set to 8, and a maximum sequence length of 64.
DiffusionEAGS For our model, we adopt the denosing strategy as top1 sampling and 1 size of MBR as typical translation task focuses on BLEU and COMET rather than diversity score.
G.3 Experimental Results
G.3.1 NAR Transformer, DisCo
The results indicate that the DisCo transformer performs poorly on low-resource translation tasks, where the size of the dataset is small. The results indicated in Table 6 are much lower than those indicated in the original paper by Kasai et al. (2020).
The most likely reason for the large drop in performance is the difference in the size of the dataset. The original DisCo paper reports a BLEU score of 27.39 and 27.34 respectively on the WMT14 EN-DE dataset. Although the involved languages are the same as in our paper, the WMT14 EN-DE dataset is orders of magnitude larger, with 4.5M pairs. Such results suggest the importance of utilizing PLMs for conditional generation tasks, especially in the case where the size of the available dataset is restricted
To account for the relatively small train set to valid/test set ratio of the dataset used in our translation experiments, which resulted in a high percentage of <UNK> tokens in the valid/test sets, we also provide results using the dictionary of a pre-trained mBART model Liu (2020). The performance benefits slightly from this change, but still lags behind those of other models.
G.3.2 Diffuseq-v2
It is notable that existing diffusion language models perform poorly on translation tasks. In this section, we introduce some observations that might aid our understanding of such behaviors.
For Diffuseq-v2, we conducted additional experiments using the same model trained on Paradetox. We observed that the entropy of token prediction probabilities in the translation model was orders of magnitude higher, indicating a greater level of uncertainty in its predictions. Similarly, the ratio of the nearest token distance to the average distance of the top five nearest tokens was significantly larger in the translation model. This analysis suggests that a simple rounding approach from continuous to discrete space may be insufficient for machine translation, at least in low-resource settings.
Appendix H Entropy Flow
Figure 4: Entropy behavior tracking in generation/training process.
In Figure 4, we illustrate the tendency of the sequential sum of entropy for various discrete generation processes. The changes of entropy during the generation process in Diffusion-EAGS, represented by the yellow line, show that our model effectively follows a gradual decrease in entropy, mirroring the inverse trend of the training process. This gradual change in entropy facilitates successful DDLM training, which results in superior text quality performance compared to other diffusion models, as demonstrated in Tables 2, 3, and 4.
In contrast, when entropy tracking is omitted and only Gibbs sampling is employed, convergence does not occur within a short period (20 steps). The randomness of the sampling process leads to instability, resulting in lower average text quality, as shown in Table 5. Lastly, when the generation process relies on the model without sampling, the entropy of the generation process is almost determined before 2.5 steps. This entropy behavior is similar to that observed in DiffusionBERT.

### v2

@@
 (2022)
-Open-ended Generation ? $\checkmark$ $\checkmark$ $\triangle$ $\checkmark$ $\times$ $\times$ $\times$
-Conditional Generation ? $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Open-ended Generation $\checkmark$ $\checkmark$ $\triangle$ $\checkmark$ $\times$ $\times$ $\times$
+Conditional Generation $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 – Context Provided ? - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
 – Format Provided ? - - $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-Table 1: Categorization of NLP tasks based on constraint aspects. $\checkmark$ indicates full support, $\times$ indicates no support, and $\triangle$ indicates partial or limited support.
-As diffusion models significantly enhance the quality and diversity of generated outputs in continuous domains such as images and audio (Song et al., 2021b), recent research has increasingly applied diffusion models to NLP  (Li et al., 2022; Gong et al., 2023a; Yuan et al., 2023; He et al., 2023).
-To integrate the diffusion process with traditional language models, research has predominantly followed two approaches: Continuous Diffusion Language Model (CDLM) (Yuan et al., 2023; Lovelace et al., 2023; Chen et al., 2023), which embed text in a continuous latent space, and Discrete Diffusion Language Model (DDLM) (He et al., 2023; Lou et al., 2024; Zhou et al., 2024; Shi et al., 2024; Sahoo et al., 2024; Zheng et al., 2024), which operate directly on the vocab space.
-Recent DDLMs adeptly reflect the highly structured nature of language; this property enables sequence control in natural language, generally resulting in high performance in NLP fields such as unconditional generation and open-ended generation  (Lou et al., 2024).
-However, for DDLMs to be widely adopted across various NLP domains, particularly in tasks such as close-ended conditional generation — hereafter referred to as dataset-guided generation — existing DDLMs still exhibit shortcomings in their performance.
-Our experiments suggest that models such as SEDD falls short in such tasks, potentially limiting their practical applications.
-Better performance can be achieved by integrating Pretrained Language Models (PLMs) through the use of an encoder-based PLM as the initial setup for embedding functions, leveraging its capability to capture overall semantics (He et al., 2023).
+Table 1: Each dataset has a different level of conditional contraints even if they are conditional generation tasks.
+$\checkmark$ indicates full support, $\times$ indicates no support, and $\triangle$ indicates partial or limited support.
+As diffusion models significantly enhance the quality and diversity of generated outputs in continuous domains such as images and audio (Song et al., 2021b), recent research has increasingly applied diffusion models to NLP, and demonstrate their promising performance over autoregressive models (Li et al., 2022; Gong et al., 2023a; He et al., 2023; Yuan et al., 2023; Lovelace et al., 2023; Chen et al., 2023; He et al., 2023; Lou et al., 2024; Zhou et al., 2024; Shi et al., 2024; Sahoo et al., 2024; Zheng et al., 2024; Nie et al., 2024).
+Moreover, these models have been shown to be scalable, and are able to address the challenge of generalizing bidirectional relationships while maintaining sensitivity to the temporal alignment between training and test data, a conventional architectural limitation often observed in autoregressive models (Nie et al., 2024).
+Although recent Discrete Diffusion Language Models (DDLMs) have demonstrated high performance in NLP tasks such as unconditional and open-ended generation compared to Continuous Diffusion Language Models (CDLMs) (Lou et al., 2024), existing DDLMs continue to face limitations in conditional generation, thereby restricting their broader applicability across diverse NLP domains.
+Our experiments suggest that models such as SEDD falls short in such tasks.
+To enhance performance of DDLMs, PLMs can be integrated into diffusion models.
+This approach leverages the capability of PLMs to elicit improved generalization capabilities in unseen tasks while simultaneously exploiting the controllability and diversity strengths inherent to diffusion models.
 However, integrating PLMs into diffusion models is non-trivial as PLMs typically predict masked elements in a single step, whereas diffusion models require step-wise denoising based on the overall semantics of each timestep sequence, and such gap yields limited results.
-Therefore, we need to consider such inconsistency to effectively adopt PLMs into DDLMs by a new methodology.
-In this paper, we introduce Diffusion-EAGS, a novel approach that effectively integrates Mask Language Model (MLM)-based PLMs with DDLMs for dataset-guided generation.
-To address the gap between the step-wise nature of diffusion models and the one-step prediction strategy of PLMs, we begin by considering MLM-based PLMs as the denoising function at each step of the diffusion process.
-Since identifying where to apply denoising during the diffusion process is challenging for PLMs, we incorporate an entropy tracking module to support their operation.
-For constructing entropy tracking module, we interpret each denoising step as a constrained Markov Random Field (cMRF), which enables us to adopt adaptive Gibbs sampling, which then facilitates leveraging entropy information from conditional sequences.
-To train DDLMs to effectively exploit an entropy tracking module, we propose a forward process based on an entropy-based noising scheduling during training phase.
-Specifically, our entropy-based noising scheduling noises lower entropy tokens first, thereby learning to progressively generate sequences guided by the entropy from entropy tracking module, which enhances the quality of the generated text.
-Experimental results demonstrate that Diffusion-EAGS achieves outstanding performance compared to existing autoregressive model (GPT-2), CDLMs, and DDLMs across various dataset-guided generation tasks.
-We demonstrate that both PLMs and entropy-based Gibbs sampling contribute to higher performance in our ablation study.
-Furthermore, our model exhibits higher diversity in certain tasks compared to LLMs and can facilitate token-level generation, indicating the potential applicability of our model across a wide range of dataset-guided generation tasks.
-In addition, we validate that our model can adapt to both bilingual and low-resource settings, which are frequently encountered in practical applications.
+Therefore, we need to consider such inconsistencies to effectively adopt PLMs into DDLMs by a new methodology.
+In this paper, we introduce Diffusion-EAGS, a novel approach that effectively integrates Mask Language Model (MLM)-based PLMs with DDLMs for conditional generation.
+To address the gap between the step-wise nature of diffusion models and the one-step prediction strategy of PLMs, we begin by adopting PLMs as the denoising function at each step of the diffusion process.
+Since identifying where to apply denoising during the diffusion process is challenging for PLMs in our experiments, we incorporate an entropy tracking module to support their operation based on entropy information.
+To train DDLMs to effectively exploit an entropy tracking module, we propose an entropy-based noising scheduling during training phase in a forward process.
+Specifically, our entropy-based noising scheduling noises lower entropy tokens first, thereby learning to progressively generate sequences guided by the entropy from entropy tracking module.
+Experimental results demonstrate that Diffusion-EAGS achieves outstanding performance compared to baseline models across various conditional generation tasks.
+Furthermore, our model exhibits higher diversity in certain tasks compared to LLMs and can facilitate token-level generation, and validate that our model can also adapt to both bilingual and low-resource settings, which are frequently encountered in practical applications, indicating the potential applicability of our model across a wide range of conditional generation tasks.
 2 Task Setting
-2.1 Dataset-Guided Generation
-In sequence-to-sequence (seq2seq) conditional generation tasks, the level of constraint imposed by the dataset plays a critical role in shaping the generation process.
-As shown in Table 1, constraints can be categorized into three levels: the provision of context alone, the provision of specific content that should be covered in the target sequence, and the provision of content format.
-Each level represents an increasing degree of constraint in the task requirements.
-The goal is to develop a model capable of generalizing across a wide range of downstream tasks, rather than being specialized for a specific task, while adhering to the constraints of the given dataset—a task we refer to as Dataset-Guided Generation.
+2.1 Fine-Grained Conditional Generation
+In conditional generation tasks, the level of conditional constraint imposed by the dataset plays a critical role in shaping the generation process.
+As shown in Table 1, conditional constraints are diverse across datasets.
+In our task, we categorize these constraints into three levels: (1) the provision of context alone, requiring the continuity of the prefix; (2) the provision of specific content to be included in the target sequence, necessitating the inclusion of certain keywords; and (3) the provision of semantic content formatting, such as transforming toxic sentences into safer alternatives or converting text from the source language to a target language.
+In our study, we aim to develop a universal diffusion framework capable of being applied across a wide range of conditional generation tasks.
 2.2 DDLM with BERT
-Diffusion Models The process of diffusion models consists of two main stages: the forward process, where noise is added to the input over multiple steps, and the reverse process, where noise is removed to restore the information.
-There are two types: Continuous and Discrete diffusion models.
-Discrete diffusion models operate the diffusion process across categorical random variables Hoogeboom et al.
-(2021).
-As discrete diffusion models show promising results in recent studies Venkatraman et al.
-(2024); Sahoo et al.
-(2024); Lou et al.
-(2024), we set up our language models grounded on discrete diffusion language models.
-BERT as constrained Markov Random Field (cMRF) As the pretraining paradigm has significantly contributed to the success of language models, the integration of PLMs is essential in the development of language models.
-Nevertheless, adopting PLMs to DDLMs is a complex task, as DDLMs require a step-wise denoising strategy, while MLM-based PLMs are trained to predict masked elements in a single step.
+BERT as Markov Random Field (MRF) As the pretraining paradigm has significantly contributed to the success of language models, the integration of PLMs is essential in the development of language models.
+Nevertheless, adopting PLMs to DDLMs is a complex task, as DDLMs require a step-wise denoising strategy, while PLMs are trained to predict masked elements in a single step.
 To effectively integrate PLMs into DDLMs, it is necessary to resolve such a discrepancy.
 Therefore, we begin by interpreting the intrinsic characteristics of MLMs as Markov Random Field (MRF), as demonstrated in Wang and Cho (2019).
-Details are in Appendix A.
-Figure 1: Potential Function Difference of MLM based pre-trained BERT generation and RocStories dataset-guided BERT generation
-However, dataset guided generation occupy different fields in terms of randomness, compared to open-ended generation.
-In Figure 1, we can observe an increase of potentials for same sequence X("The quick brown fox jumps over the lazy dog.") on RocStories.
-That is, it is necessary to investigate whether the dataset-constrained space is still subject to MRF.
+Details are in Appendix B.
+constrained MRF (cMRF) Even though open-ended MLMs are MRFs, we observe a significant increase in log-potential values for sequences when guided by the RocStories conditions, as shown in Figure 1.
+Additional experiments supporting this observation are detailed in Appendix A.
+Figure 1: When a condition is provided, the distribution of potential values for the samples is shifted on a logarithmic scale.
+This suggests that conditional generation models differ from open-ended models in randomness, making it crucial to investigate whether conditionally constrained spaces adhere to MRF properties.
+Through the derivation process outlined in Appendix B, we demonstrate that these spaces retain the characteristics of an MRF.
 Given a sequence of fully connected random variables $X=(x_{1},x_{2},\ldots,x_{L})$ with a set of observed variables $Y=(y_{1},y_{2},\ldots,y_{K})$, $f_{\theta}(X\backslash\{x_{l}\};Y)$ can be represented by logits from MLM given the sequence $X\backslash\{x_{l}\}$ and the observed variables $Y$.
-Then, similarly to Equation 2, log-potential can be calculated as:
-$\log\phi_{l}(X;Y)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\};Y),&\text{if }\text{[M]}\notin X_{l-1:l+1}\\ \\ 0,&\text{otherwise}\end{cases}$
-By this, we can define the potential function for the clique as:
-$\phi(X;Y)=\exp\left(\sum_{l=1}^{L}\log\phi_{l}(X;Y)\right)$
+Then, log-potential can be calculated as:
+$\log\phi_{l}(X;Y)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\};Y),&\text{if }\text{[MASK]}\notin X_{l-1:l+1},\\ \\ 0,&\text{otherwise}\end{cases}$
+where $1h(x_{l})$ is a one-hot vector with $x_{l}$ set to 1. By this, we can define the potential function for the clique as:
+$\phi(X;Y)=\exp\left(\sum_{l=1}^{L}\log\phi_{l}(X;Y)\right)$ (1)
 Then, the constrained probability of sequence X is:
@@
 3 Methodology
-Figure 2: Overview of Diffusion-EAGS. During forward steps, We schedule the noise based on ENS using PLM. In generation steps, we denoise the sequences based on EAGS.
-We propose Diffusion-EAGS that integrates PLM with DDLMs in dataset-guided generation, through Entropy-based Adaptive Gibbs Sampling (EAGS) and Entropy-based Noise Scheduling (ENS).
+Figure 2: Given a conditional constraint $Y$ and a generation target $X$, the training process utilizes Entropy-based Sampling (ENS) to determine which positions in the masked sequence should be denoised when transitioning from one timestep to the next.
+The tokens at the corresponding masked positions ([M] denotes masked tokens) are then generated using $f_{\theta}$.
+The loss is subsequently computed, and the parameters are updated accordingly.
+During the inference process, the sequence $X$, initially composed of mask tokens, is iteratively refined through the diffusion process to generate $X$ that satisfies the given condition $Y$.
+In this section, we propose Diffusion-EAGS that integrates PLM into DDLMs through Entropy-based Adaptive Gibbs Sampling (EAGS) and Entropy-based Noise Scheduling (ENS).
 The overview of our process is in Figure 2.
 3.1 Generation Process: Entropy-based Adaptive Gibbs Sampling (EAGS)
-Building on the premise in Section 2.2, we propose EAGS to leverage PLM as the function of denoising process in diffusion models.
-As MLM-based PLMs can be interpreted as cMRF, we can apply Gibbs sampling in each denoising step in DDLM.
-By Gibbs sampling, we can sample each token $x_{l}$ from its constrained distribution $p(x_{l}|X\backslash\{x_{l}\};Y)$.
-Rather than performing Gibbs sampling, we employ EAGS, as the entropy of each variable indicates the amount of information at certain position in the sequence.
-By generating tokens in the order of decreasing entropy, we prioritize the generation of the least informative parts of the sequence.
-Such prioritized generation makes a more structured sequences, leveraging the already established syntactic context.
-According to Equation 1, we can compute the conditional probability for each masked position. The entropy $H$ for a given token $x_{l}$ is derived as:
-$\begin{split}&H(x_{l}|Y,f)=\\ &-\sum_{m=1}^{M}p(x_{l}=v_{m}|x_{l},Y,f)\log p(x_{l}=v_{m}|x_{l},Y,f)\end{split}$
+As discussed in Section 2.2, MLM-based PLMs can be interpreted as cMRFs, allowing Gibbs sampling to be applied at each denoising step in DDLMs by leveraging the properties of MRFs.
+As PLMs often struggle to select the next denoising targets, it is important to highlight that the primary purpose of adopting Gibbs sampling is not only to bridge the gap between PLMs and diffusion models but also to assist PLMs in identifying which absorbing tokens should be denoised.
+Therefore, we propose Entropy-Adaptive Gibbs Sampling (EAGS) to address the challenge PLMs face in identifying the appropriate denoising targets.
+In EAGS, the entropy of each variable quantifies the amount of uncertainty at a specific position within the sequence.
+This enables sampling of each token $x_{l}$ from its constrained distribution $p(x_{l}|X\backslash\{x_{l}\};Y)$ in the descending order of entropy.
+By prioritizing the generation of the least informative parts of the sequence, EAGS facilitates the creation of more structured sequences by leveraging the syntactic context that has already been established.
+Specifically, according to Equation (2), the probability $p^{t}(x_{l}=v_{m}|Y,f_{\theta})$ of the token $v_{m}$ given the context $Y$ and $f_{\theta}$ at step $t$ can be calculated as:
+$p^{t}(x_{l}=v_{m}|Y,f_{\theta})=\frac{\exp(f_{\theta}(v_{m}|X^{t}))}{\sum_{m^{\prime}=1}^{M}\exp(f_{\theta}(v_{m^{\prime}}|X^{t}))}$
+, where $f(v_{m}|X^{t})$ is the logit corresponding to the token $v_{m}$ for the position $x_{l}$ at step $t$.
+That is, we can compute the conditional probability for each masked position.
+With the conditional probability, the entropy $H$ for a variable $x_{l}$, logits function $f_{\theta}$, and the dataset context $Y$ is derived as:
+$\begin{split}&H(x_{l}|Y,f_{\theta})=\\ &-\sum_{m=1}^{M}p(x_{l}=v_{m}|x_{l},Y,f_{\theta})\log p(x_{l}=v_{m}|x_{l},Y,f_{\theta})\end{split}$
 With $H$, our approach for T step-generation process can be formalized as follows:
   1. 1.
-    Entropy Calculation: Compute the entropy $H(x_{l}\mid Y,f)$ for each variable $x_{l}$.
+    Entropy Calculation: Compute the entropy $H(x_{l}\mid Y,f_{\theta})$ for each variable $x_{l}$.
   2. 2.
     Variable Selection: Select the variable $x_{l^{*}}$ with the highest entropy for sampling
-    $l^{*}=\arg\max_{l}H(x_{l}\mid Y,f)$
+    $l^{*}=\arg\max_{l}H(x_{l}\mid Y,f_{\theta})$
   3. 3.
-    Sampling: Sample $x_{l^{*}}$ from its conditional distribution $p(x_{l^{*}}\mid Y,f^{*})$.
+    Sampling: Sample $x_{l^{*}}$ from its conditional distribution $p(x_{l^{*}}\mid Y,f_{\theta}^{*})$.
   4. 4.
@@
 3.2 Forward Process: Entropy-based Noise Scheduling (ENS)
-To enhance the facilitation of Adaptive Gibbs Sampling in the generation process, it is essential to describe a similar process in the training phase.
-Therefore, we schedule the forward process of diffusion training based on the entropy $H(x_{l})$ of positions $x_{l}$ with the input sequence $[Y|X]$.
+To enhance the facilitation of Adaptive Gibbs Sampling in the generation process, it is essential to simulate a similar process in the training phase.
+Therefore, we schedule the forward process of diffusion training based on the entropy $H(x_{l})$ of position $x_{l}$ with the input sequence $[Y|X]$.
 Specifically, the position-wise entropy is calculated by the PLM that shares a similar language base with diffusion-trained model.
@@
 $q_{mn}=\begin{cases}q_{mm}=1&\text{if }x_{l}\notin\text{MIN}([H(x_{1}),\cdots,H(x_{L})],\frac{L}{T})\\ q_{mM}=1&\text{if }x_{l}\in\text{MIN}([H(x_{1}),\cdots,H(x_{L})],\frac{L}{T})\\ 0&\text{otherwise}\end{cases}$
-This entropy-based noise scheduling approach ensures that the forward process in diffusion training closely mirrors the generation process, thereby enhancing the effectiveness of Adaptive Gibbs Sampling in language generation.
-3.3 Training & Inference
-Training Process
+This entropy-based noise scheduling approach ensures that the forward process in diffusion training closely mirrors the generation process, thereby enhancing the effectiveness of EAGS in language generation.
+3.3 Training and Inference
+ENS Process:
 Input: Context $Y$ and Dataset $D$
-for $BatchStep=0..N$  do
-       $x\sim D$
-       $\text{t}\sim\text{Randint}(0,T)$
-       $f\leftarrow PLM(x|Y)$
-       $\mathcal{H}\leftarrow H(x|Y,f)$
-       $x^{t}\leftarrow\text{Forward}(x_{0},\mathcal{H},t)$
-       $x^{t+1}\leftarrow\text{Forward}(x_{0},\mathcal{H},t+1)$
-       $L_{s}=-\sum_{i}(q(x^{t}_{i})\mid x_{t+1})\log(p_{\theta}(x^{t}_{i})\mid x_{t+1})$
+[…]
+       $x\sim D$ // Sample data from $D$
+       $t\sim\text{Randint}(0,T)$ // Sample random timestep
+       $f\leftarrow\text{PLM}(x\mid Y)$ // Compute logits using the PLM
+       $\mathcal{H}\leftarrow H(x\mid Y,f)$ // Calculate Entropy
+       $x^{t}\leftarrow\text{Forward}(x_{0},\mathcal{H},t)$ // Forward at $t$
+       $x^{t+1}\leftarrow\text{Forward}(x_{0},\mathcal{H},t+1)$ // Forward at $t+1$
+       $L_{s}=-\sum_{i}q(x^{t}_{i}\mid x_{t+1})\log p_{\theta}(x^{t}_{i}\mid x_{t+1})$ // Loss calculation
 end for
-Generation Process
+EAGS Process:
 Input: Sequence Length $L$, Total Timestep $T$, Trained Model $M$, Mask Sequence Generator $G_{M}$, and Context $Y$
+for $t=T$ to 00 do
+       if $t=T$ then
+             $x^{T}\leftarrow G_{M}(L,Y)$ // Initialize a sequence of $L$
+       else
+             $f_{\theta}\leftarrow M(x^{t},Y)$ // Compute logits at timestep $t$
 […]
-       if  $t=T$  then // Initialize $x^{T}$
-             $x^{T}\leftarrow G_{M}(L,Y)$
-       else  // Denoise step
-             $f\leftarrow M(x^{t},Y)$
-[…]
-             $x^{t-1}=\text{Sampling}(x^{t},l^{*},Y)$
+             $x^{t-1}\leftarrow\text{Sampling}(x^{t},l^{*},Y)$ // Sample from the previous timestep
        end if
 end for
-Algorithm 1 Training & Generation
+Algorithm 1 ENS and EAGS Algorithm
 Our whole process is in Algorithm 1.
@@
 Instead, to ensure the completeness of sentences, we implement Minimum Bayes Risk (MBR) decoding in the final generation step.
-We follow the traditional diffusion loss in Equation (Ho et al., 2020b) with changing MSE to Cross Entrpy Loss.
+We follow the traditional diffusion loss in Equation (Ho et al., 2020b), changing MSE with Cross Entrpy Loss.
 4 Experiments
 4.1 Tasks and Datasets
-Question Generation We use the Quasar-T dataset (Dhingra et al., 2017) to generate questions from passages and answers.
-Paraphrase The QQP dataset (Wang et al., 2017) assesses paraphrase detection by treating one question as a paraphrase of another.
 Social Datasets The Paradetox dataset (Logacheva et al., 2022) focuses on removing profanities.
 The Deontology from ETHICS dataset (Hendrycks et al., 2023) evaluates ethical judgment based on scenarios.
+Paraphrase The QQP dataset (Wang et al., 2017) assesses paraphrase detection by treating one question as a paraphrase of another.
+Question Generation We use the Quasar-T dataset (Dhingra et al., 2017) to generate questions from passages and answers.
 Open-ended Generation The RocStories (Mostafazadeh et al., 2016) dataset is used to generate coherent story continuations from a given context.
-Detailed explanations are available in Appendix B and Dataset statistics are shown in Table 8.
+Detailed explanations are available in Appendix C and Dataset statistics are shown in Table 8.
 4.2 Baselines
-Since our objective is centered on generation tasks, we compare Diffusion-EAGS with three categories of baselines with the model similar to the size of RoBERTa-base: AutoRegressive (AR) model, CDLM, and DDLM.
-For AR model (Vaswani et al., 2023), we employ GPT-2 (Radford et al., 2019), which is renowned for its proficiency in generation tasks.
+Since our objective is centered on generation tasks, we compare Diffusion-EAGS with four categories of baselines with the model similar to the size of RoBERTa-base (Liu et al., 2020): Auto-regressive Models (ARMs), CMLMs (Conditional Masked Language Models, Ghazvininejad et al.
+(2019a)), CDLMs, and DDLMs.
+For ARMs (Vaswani et al., 2023), we employ GPT-2 (Radford et al., 2019), which is renowned for its proficiency in generation tasks.
+For CMLMs, we utilize CMLM (Ghazvininejad et al., 2019a) and DisCo (Kasai et al., 2020), which are transformer-based NAR models.
+Inference is carried out using the Mask-Predict (Ghazvininejad et al., 2019a) and the Easy-First (Kasai et al., 2020) inference algorithm respectively.
 For CDLMs, our baseline includes DiffuSeq (Gong et al., 2023a) and DINOISER (Ye et al., 2024).
@@
 For SEDD, we download the pretrained version and fine-tune on it.
-Experimental results on LLMs can be found in Appendix C.
+Experimental results on LLMs can be found in Appendix E.
 4.3 Metrics
@@
 We use 1 A100 GPU, and the batch size is set to 256.
-QG
-Text Quality Diversity
-Model Step PPL $\downarrow$ MAUVE $\uparrow$ SOME $\uparrow$ VS(ngram) $\uparrow$ VS(emb) $\uparrow$ self-bleu $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$
-GPT-2 1 124.8 0.141 0.759 4.564 3.130 0.176 0.210 0.629
-DiffuSeq 20 395.0 0.149 0.730 1.555 1.274 0.901 0.170 0.564
-DINOISER 2000 155.9 0.159 0.776 1.396 1.121 0.944 0.166 0.553
-DiffusionBert 2000 513.6 0.150 0.712 3.040 2.209 0.566 0.392 0.759
-Diffusion-EAGS 5 80.7 0.121 0.782 4.646 3.538 0.152 0.403 0.798
-QQP
-Model Step PPL $\downarrow$ MAUVE $\uparrow$ SOME $\uparrow$ VS(ngram) $\uparrow$ VS(emb) $\uparrow$ self-bleu $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$
-GPT-2 1 66.270 0.112 0.754 3.886 2.566 0.423 0.344 0.787
-DiffuSeq 2000 124.247 0.00674 0.709 1.927 1.242 0.813 0.226 0.543
-DINOISER 20 79.742 0.0042 0.821 1.421 1.126 0.935 0.264 0.542
-DiffusionBert 2000 500.959 0.0709 0.618 4.489 2.836 0.196 0.321 0.761
-Diffusion-EAGS 5 48.106 0.683 0.824 4.006 2.390 0.338 0.421 0.832
-Table 2: QG & QQP Generation
 ParaDetox
@@
 GPT-2 1 389.1 0.503 0.717 3.925 2.640 0.429 0.312 0.748
+CMLM w/ Mask-Predict 10 669.9 0.0234 0.588 1.000 1.000 1.000 0.451 0.633
+DisCo w/ Easy-First 10 716.1 0.0344 0.576 1.000 1.000 1.000 0.438 0.583
+DiffusionBert 2000 775.9 0.737 0.716 3.101 2.058 0.599 0.424 0.826
 DiffuSeq 2000 $\geq 1k$ 0.683 0.703 2.059 1.465 0.841 0.410 0.820
 DINOISER 20 124.8 0.255 0.767 2.287 2.174 0.981 0.211 0.486
-DiffusionBert 2000 775.9 0.737 0.716 3.101 2.058 0.599 0.424 0.826
 SEDD 1024 $\geq 1k$ NA 0.664 4.746 4.063 0.119 0.451 0.846
@@
 Diffusion-EAGS 5 55.1 0.412 0.835 4.898 4.009 0.056 0.418 0.806
-Table 3: Social Generation
+Table 2: Social Generation – Diversity values associated with higher perplexity (PPL) are displayed in gray, as increased perplexity typically indicates degenerate sequences.
+QQP
+Model Step PPL $\downarrow$ MAUVE $\uparrow$ SOME $\uparrow$ VS(ngram) $\uparrow$ VS(emb) $\uparrow$ self-bleu $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$
+GPT-2 1 66.270 0.112 0.754 3.886 2.566 0.423 0.344 0.787
+DiffuSeq 2000 124.247 0.00674 0.709 1.927 1.242 0.813 0.226 0.543
+DINOISER 20 79.742 0.0042 0.821 1.421 1.126 0.935 0.264 0.542
+DiffusionBert 2000 500.959 0.0709 0.618 4.489 2.836 0.196 0.321 0.761
+Diffusion-EAGS 5 48.106 0.683 0.824 4.006 2.390 0.338 0.421 0.832
+QG
+Text Quality Diversity
+Model Step PPL $\downarrow$ MAUVE $\uparrow$ SOME $\uparrow$ VS(ngram) $\uparrow$ VS(emb) $\uparrow$ self-bleu $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$
+GPT-2 1 124.8 0.141 0.759 4.564 3.130 0.176 0.210 0.629
+DiffuSeq 20 395.0 0.149 0.730 1.555 1.274 0.901 0.170 0.564
+DINOISER 2000 155.9 0.159 0.776 1.396 1.121 0.944 0.166 0.553
+DiffusionBert 2000 513.6 0.150 0.712 3.040 2.209 0.566 0.392 0.759
+Diffusion-EAGS 5 80.7 0.121 0.782 4.646 3.538 0.152 0.403 0.798
+Table 3: QG & QQP Generation
 5 Results
 As shown in Table 234, our model consistently exhibits exceptional performance in terms of text quality while simultaneously maintaining diversity when compared to baseline models.
-As illustrated in Table 2, Diffusion-EAGS generates the responses with the highest PPL score for QG, and highest MAUVE and PPL score for QQP.
 In Table 2 Paradetox, our model demonstrates superior performance across all evaluated metrics.
 Such phenomenon represents that our model based on encoder-based PLMs show robustness on diverse perturbations from daily dialogues.
-When PPL exceeds 1,000, the model is considered to have failed in generating natural sequences, and is thus represented in gray color.
+When PPL exceeds 600, the model is considered to have failed in generating natural sequences, and is thus represented in gray color.
 In the context of Deontology, our model exceeds the baseline PPL and MAUVE scores, whereas SOME score represent the sufficient quality of text with the highest diversity score of 4.755.
+In the paradetox experiment, the text quality produced by the CMLM and SEDD models was found to be low.
+Consequently, these models were excluded from subsequent experiments.
+As illustrated in Table 3, Diffusion-EAGS generates the responses with the highest PPL score for QG, and highest MAUVE and PPL score for QQP.
 Additionally, in Table 4 where semantic consistency is crucial, our model significantly outperforms the original dataset’s PPL by 23 points by leveraging PLMs while maintaining a high diversity score of 4.837.
-5.1 Ours vs AR model
-Diffusion-EAGS demonstrates high level of text quality surpassing that of GPT-2 in Table 2 in text quality.
-Notably, our model effectively reflects the pattern and style of the dataset, attributed to the capability of encoder-based PLMs.
-Specifically, as reported in Table 3-ParaDetox, our model excels at capturing semantic information with the highest MAUVE score of 0.733, despite ParaDetox being a challenging colloquial dataset including slang, numerous abbreviations, and various perturbations.
-In addition, our model demonstrate robustness to such perturbations with 69.5 PPL, while GPT-2 shows a lower performance of 389.1 PPL.
-As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 23, and 4. These results underscore our model’s efficacy in generating diverse responses.
+Diffusion-EAGS demonstrates the highest MAUVE score of 0.733 in Table 2-ParaDetox, and high level of text quality surpassing that of GPT-2 in Table 3 in text quality.
+ParaDetox is colloquial dataset including slang, numerous abbreviations, and various perturbations, so our model demonstrate robustness to such perturbations with 69.5 PPL.
+As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 23, and 4.
+These results underscore our model’s efficacy in generating diverse responses.
 Model PPL $\downarrow$ SOME $\uparrow$ LLM-c $\uparrow$ VS(ngram) $\uparrow$ self-bleu $\downarrow$
@@
 Diffusion-EAGS 67.3 0.844 0.87 4.837 0.058
-Table 4: ROC Generation
-Specifically, in the context of story generation as shown in Table 4, our model not only demonstrates effectiveness in enhancing both the quality and diversity of the text but also maintains coherence with the contextual storyline.
-This is further supported by the high LLM-c scores, which confirm the plausibility of the narratives produced by our model.
-5.2 Ours vs CDLMs
+Table 4: ROC Generation – Because ROC is open-ended generation task, we include SEDD in this experiment.
 Notably, CDLMs demonstrate a noticeable deficiency in diversity.
 In contrast, our model excels at producing significantly more diverse sequences.
-This enhanced performance can be attributed to the integration of PLM into our framework, which evidently enriches the diversity of the generated outputs.
-In addition, our model generally achieves higher text quality scores than CDLMs across all experiments.
-Diffusion-EAGS not only surpasses Diffuseq and DINOISER in PPL and SOME but also achieves a MAUVE score improvement of over 0.4, thus indicating that the generated responses accurately reflect the dataset while preserving grammatical integrity.
-Furthermore, our models require only a few steps, resulting in a 40% reduction of costs with higher quality and diversity.
-5.3 Ours vs DDLMs
-Contrary to CDLMs, DiffusionBERT exhibits limitations in text quality, as shown in Table 234.
+Furthermore, our models require only a few steps, while resulting in higher quality and diversity.
+DiffusionBERT exhibits limitations in text quality, as shown in Table 234.
 On the other hand, our model achieves significantly higher scores across all quality metrics.
-This observation suggests that the naive application of BERT within the diffusion process fails to fully harness the capabilities of PLMs.
+This observation suggests that the naive application of BERT (Devlin et al., 2019) within the diffusion process fails to fully harness the capabilities of PLMs.
 SEDD, an open-ended generation model shows low performance under text generation in highly constrained tasks such as ParaDetox.
-Even in open-ended generation settings of Table 4, Diffusion-EAGS shows higher text quality and semantic scores.
 Additionally, our model does not require more than 5 diffusion steps, as it can adaptively recover the absorbing state through an entropy-based generation approach.
-Above results further exemplify the high performance and the time efficiency of our model.
+A detailed comparison is provided in Appendix D.
 6 Analysis
@@
 To explore the effectiveness of our model’s components, we conduct ablation studies focusing on three key elements: Entropy Adaptation, stepwise Gibbs Sampling, and PLM in Table 5.
-The omission of EAGS initially results in a significant decline in performance, a noticeable decrease in text quality and diversity.
+The omission of EAGS initially leads to a substantial decline in performance and text quality, producing degenerated results similar to those of traditional CMLMs.
 Dataset PPL MAUVE SOME VS(ngram) VS(emb)
@@
 Roc $\geq$2K 0.004 0.604 4.315 2.994
-Table 5: Ablation Study
-We find that EAGS significantly contributes to a gradual entropy decrease, as shown in Appendix H.
+Table 5: Ablation Study : Deont denotes Deontology
+We find that EAGS significantly contributes to a gradual entropy decrease, as shown in Appendix J.
 These observations highlight the critical role of selecting the sampling position based on given information in sequence generation.
@@
 Our model operating within discrete space enables us to manipulate the output sequences using explicit instructions.
-To further explore this capability, we conduct the generation of sequences based on keywords positioned in the middle and at the end of masked sequences.
+To further explore this capability, we conduct the generation of sequences based on keywords positioned in the middle and at the end of masked sequences, which is challenging for AR models (Keskar et al., 2019).
 Initially, we provide the same contextual input while varying the keywords.
@@
 Experimental results also indicates its substantial potential and suitability for diverse applications where direct controllability is crucial, such as in story generation.
-6.3 Bilinguality & Low Resource
+6.3 Bilinguality & Low Resource Settings
 Model SacreBLEU COMET XCOMET
-DisCo (Kasai et al., 2020)
+DisCo
 w/ Easy-First 3.2806 0.2447 0.2414
@@
 w/ Mask-Predict 3.7748 0.2466 0.2119
-Diffuseq-v2 (Gong et al., 2023b) 1.90 0.3242 0.2628
+Diffuseq-v2 1.90 0.3242 0.2628
 SEDD
@@
 mBART-50-FT 19.6536 0.7576 0.8748
-Table 6: Translation Results
+Table 6: En-De Translation Results
 Context Jake was playing with his toys. He accidentally broke his favorite one.
@@
 […]
-We establish baselines using non-autoregressive (NAR) transformer models, such as Mask-and-Predict (Ghazvininejad et al., 2019a) and Easy-First (Kasai et al., 2020) - which bear resemblance to discrete diffusion models - diffusion models such as Diffuseq-v2 and SEDD, traditional translation models such as mBART-50 (Tang et al., 2020) and NLLB.
+Although our baseline diffusion models demonstrated significant performance improvements over CMLMs, CMLMs were primarily developed to enhance translation performance.
+Therefore, we conducted additional experiments to further validate our methodology such as Mask-and-Predict and Easy-First, diffusion models such as Diffuseq-v2 (Gong et al., 2023b) and SEDD, traditional translation models such as mBART-50 (Tang et al., 2020) and NLLB.
 For evaluation metrics, we utilize sacreBLEU (Post, 2018) and neural-net scores such as COMET (Rei et al., 2020) and XCOMET (Guerreiro et al., 2023).
@@
 All diffusion baseline models struggle to produce correct outputs.
-For example, the pretrained SEDD model fails to effectively leverage conditional information, even after finetuning on the training datasets, consistent with the limitation observed in Section 5.3.
+For example, the pretrained SEDD model fails to effectively leverage conditional information, even after finetuning on the training datasets, consistent with the limitation observed in Section D.3.
 Similar challenges arise in NAR transformer baselines.
@@
 On the other hand, our proposed model, by incorporating a PLM, demonstrates promising results.
+In addition, these results show that our methodology can also be used with encoder-decoder models.
 Interestingly, the output of the pretrained NLLB model (NLLB-naive-600M, not finetuned) reveal that neural network-based metrics are susceptible to the interference problem, specifically translated by other languages, even though we provide the language specific token.
@@
 This suggests that future research should address the semantic capabilities of diffusion models to help bridge this gap.
-Details are in Appendix G.
+Details are in Appendix I.
 6.4 Diversity Saturation
@@
 Inspired by the observation that Diffusion-EAGS consistently excel in terms of diversity across all results, we delve further into the diversity capabilities of our model.
-We assess the diversity performance in dataset-guided generation compared to LLMs.
+We assess the diversity performance in conditional generation compared to LLMs.
 We measure the VS for 5 to 100 generations under a single condition.
@@
 The experiment is conducted on the ‘deontology’ dataset which allows high output diversity in its settings.
-Details of using LLMs are provided in Appendix E.
+Details of using LLMs are provided in Appendix G.
 Figure 3 demonstrates that the diversity saturation graph for Diffusion-EAGS has a relatively steep slope, while GPT models saturate at lower values.
@@
 The core of these diffusion methodologies lies in the addition and restoration of random noise to facilitate generation.
-However, Cold Diffusion (Bansal et al., 2022) argues that such operations do not necessarily have to be governed by stochastic randomness.
+However, authors of Cold Diffusion (Bansal et al., 2022) argue that such operations do not necessarily have to be governed by stochastic randomness.
 Discrete diffusion models Building on this rationale, D3PM (Austin et al., 2023) propose the discrete restoration-generation approach and DiffusionBERT He et al.
@@
 SEDDLou et al.
-(2024) proposes score entropy inspired by MLM loss.
+(2024) propose score entropy inspired by MLM loss.
 Recent works by Shi et al.
@@
 Zheng et al.
-(2024) further closes the gap between discrete diffusion models and masked models.
-These researches make an improvement on the open ended generation task.
-Venkatraman et al.
-(2024) showcase that SEDD priors can be used to sample from a posterior distribution, namely text infilling.
+(2024) further close the gap between discrete diffusion models and masked models, and also, by correcting the numerical precision error during the sampling process in SEDD-based models, reveal that the true generation perplexity of DDLMs lags behind that of AR models.
+These research make an improvement on the open ended generation task.
+Furthermore, Venkatraman et al.
+(2024) showcase that SEDD priors can be used to sample from a posterior distribution, namely text infilling, and Nie et al.
+(2024) demonstrate that DDLMs are scalable, and have advantages over traditional AR models.
 8 Conclusions
-In this work, we introduce Diffusion-EAGS, integrating encoder-based PLMs with diffusion models for dataset-guided generation.
+In this work, we introduce Diffusion-EAGS, integrating PLMs with diffusion models for conditional generation.
 By leveraging step-wise diffusion models and the property of MLM, we demonstrate that MLM-based PLMs can serve as denoising function with Entropy-based Adaptive Gibbs Sampling.
 We also propose entropy-based noise scheduling during training for adaptive sampling.
-Experimental results show that Diffusion-EAGS outperforms existing baselines, achieving high diversity and precise token-level control.
-These results demonstrate its enhanced text generation quality and diversity, as well as wide applicability.
+Experimental results demonstrate that Diffusion-EAGS outperforms existing baselines, yielding improved text generation quality, enhanced diversity, broad applicability, and more precise token-level control.
 Limitations
 While Diffusion-EAGS demonstrates significant improvements in conditional generation tasks, there are several limitations.
-First, the quality of the text generated by our model depends on PLMs such as BERT.
-Second, as our method is currently focused on text generation tasks, its applicability to text classification tasks, such as Named Entity Recognition and Part-of-Speech Tagging, remains unexplored.
-Future research could explore extending this method to other NLP tasks, as well as investigating the potential for reducing computational costs through more efficient entropy calculation techniques.
-Additionally, addressing the biases inherent in pretrained models remains a critical area of focus to ensure the robustness and fairness of generated outputs.
+Firstly, as our method is currently focused on text generation tasks, its applicability to text classification tasks, such as Named Entity Recognition and Part-of-Speech Tagging, remains unexplored.
+Future research could explore extending this method to other NLP tasks.
+Secondly, future works could investigate the potential in reducing computational costs through more efficient entropy calculation techniques.
+Lastly, although our current efforts concentrate on developing and validating our framework using encoder-only and encoder-decoder architectures, the potential integration of autoregressive (AR) models remains unexplored.
+However, since the denoising function—with its entropy-based, sequential ordering—aligns well with autoregressive decoding strategies, we posit that AR models can also serve as effective initializations within our proposed framework.
 Ethical Statements
@@
 Specifically, its high diversity performance enables active utilization in dynamically adapting text generations to meet specific social criteria.
-Appendix A BERT as a MRF
+Appendix A Measuring Potential Function $\phi(X)$ in MLM
+In this section, we provide additional experimental details and results to support the observation that open-ended Masked Language Models (MLMs) exhibit increased potentials for the same sequence under different dataset constraints.
+A.1 Experimental Setup
+  • Model We use the pre-trained BERT large model (bert-large-cased) as the base model for all experiments.
+Additionally, we incorporate RocStories-conditioned guidance with the pre-trained model and use a fine-tuned BERT model on the RocStories dataset to evaluate the impact of dataset-specific constraints.
+  • Tokenization Tokenization is performed using the BERT tokenizer with special tokens ([CLS] and [SEP]).
+  • Potential Calculation The potential function $\phi(X)$ is computed as described in Equation 1 and  3 , where the log-potentials are obtained for each token using masked token logits.
+  • Datasets
+    • RocStories: Structured narratives from the RocStories dataset.
+A.2 Results of Experiment and Implications for Conditional Generation
+Using the BERT-large-cased model, the average log potential value for the standard MLM was 156.6150, while incorporating RocStories guidance increased this value to 175.5332, highlighting the impact of dataset-specific constraints.
+Additionally, fine-tuning the same model on RocStories resulted in an average potential function value of 3.7551 (on an exponential scale), demonstrating substantial variation introduced by conditional generation settings.
+The results demonstrate the significant influence of dataset structure on the potential function in MLMs.
+Specifically, structured datasets like RocStories enforce stronger narrative constraints, leading to higher potentials and greater coherence in sequence generation.
+This supports the main text’s argument that conditionally constrained spaces enhance the consistency and predictability of MLM outputs.
+Appendix B BERT as a MRF
 Let $X=(x_{1},\ldots,x_{L})$ be a sequence of random variables $x_{l}$, each taking a value from a vocabulary $V=\{v_{1},\ldots,v_{M}\}$.
@@
 where each log-potential $\phi_{l}(X)$ is:
-$\log\phi_{l}(X)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\}),&\text{if }\text{[M]}\notin X_{l-1:l+1}\\ \\ 0,&\text{otherwise}\end{cases}$ (3)
+$\log\phi_{l}(X)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\}),&\text{if }\text{[MASK]}\notin X_{l-1:l+1}\\ \\ 0,&\text{otherwise}\end{cases}$ (4)
 Here, $f_{\theta}(X\backslash\{x_{l}\})$ is a function that maps the sequence $X\backslash\{x_{l}\}$ to a real-valued vector in $\mathbb{R}^{M}$, and $1h(x_{l})$ is a one-hot vector with $x_{l}$ set to 1.
@@
 where $D$ is the training dataset.
-Appendix B Dataset Explanations
+Appendix C Dataset Explanations
 Quasar-T QQP ParaDetox Deontology RocStories
@@
 Table 8: Dataset Statistics
+Paradetox The objective of the Paradetox (Logacheva et al., 2022) is to delete the profanities in source sentence.
+It comprises of toxic and neutral utterances, curated from the Jigsaw, Reddit, and Twitter datasets.
+Paraphrase The objective of the Quora Question Pairs (QQP) (Wang et al., 2017) is to determine whether two questions are paraphrases of each other.
+We process the QQP dataset by treating one question as a paraphrase of another, a method commonly employed to assess the effectiveness of diffusion models.
 QG The objective of Question Generation (QG) is to generate valid and fluent questions based on a given passage and a specified answer.
@@
 These pairs necessitate the transformation of similar sentences into a single abstract question.
-Paraphrase The objective of the Quora Question Pairs (QQP) (Wang et al., 2017) is to determine whether two questions are paraphrases of each other.
-We process the QQP dataset by treating one question as a paraphrase of another, a method commonly employed to assess the effectiveness of diffusion models.
-Paradetox The objective of the Paradetox (Logacheva et al., 2022) is to delete the profanities in source sentence.
-It comprises of toxic and neutral utterances, curated from the Jigsaw, Reddit, and Twitter datasets.
 Deontology The objective of Deontology (Hendrycks et al., 2023) is to to evaluate the capability of models to make ethical judgments from a deontological perspective.
@@
 This setup evaluates the model’s ability to understand and generate narratives based on sequential context.
-Appendix C Experimental Outputs
-C.1 LLM Evaluation
+Appendix D Detailed analysis of Results
+D.1 Ours vs AR model
+Notably, our model effectively reflects the pattern and style of the dataset, attributed to the capability of encoder-based PLMs.
+Specifically, as reported in Table 2-ParaDetox, our model excels at capturing semantic information with the highest MAUVE score of 0.733, despite ParaDetox being a challenging colloquial dataset including slang, numerous abbreviations, and various perturbations.
+In addition, our model demonstrate robustness to such perturbations with 69.5 PPL, while GPT-2 shows a lower performance of 389.1 PPL.
+Diffusion-EAGS demonstrates high level of text quality surpassing that of GPT-2 in Table 3 in text quality.
+As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 23 and 4.
+These results underscore our model’s efficacy in generating diverse responses.
+Specifically, in the context of story generation as shown in Table 4, our model not only demonstrates effectiveness in enhancing both the quality and diversity of the text but also maintains coherence with the contextual storyline.
+This is further supported by the high LLM-c scores, which confirm the plausibility of the narratives produced by our model.
+D.2 Ours vs CDLMs
+Notably, CDLMs demonstrate a noticeable deficiency in diversity.
+In contrast, our model excels at producing significantly more diverse sequences.
+This enhanced performance can be attributed to the integration of PLM into our framework, which evidently enriches the diversity of the generated outputs.
+In addition, our model generally achieves higher text quality scores than CDLMs across all experiments.
+In case, our model shows a lower PPL than DINOISER, but DINOISER’s MAUVE score approaches zero, indicating that their generated responses are not based on the dataset.
+Diffusion-EAGS not only surpasses Diffuseq and DINOISER in PPL and SOME but also achieves a MAUVE score improvement of over 0.4, thus indicating that the generated responses accurately reflect the dataset while preserving grammatical integrity.
+Furthermore, our models require only a few steps, resulting in a 40% reduction of costs with higher quality and diversity.
+D.3 Ours vs DDLMs
+Contrary to CDLMs, DiffusionBERT exhibits limitations in text quality, as shown in Table 234.
+On the other hand, our model achieves significantly higher scores across all quality metrics.
+This observation suggests that the naive application of BERT (Devlin et al., 2019) within the diffusion process fails to fully harness the capabilities of PLMs.
+SEDD, an open-ended generation model shows low performance under text generation in highly constrained tasks such as ParaDetox.
+Even in open-ended generation settings of Table 4, Diffusion-EAGS shows higher text quality and semantic scores.
+Additionally, our model does not require more than 5 diffusion steps, as it can adaptively recover the absorbing state through an entropy-based generation approach.
+Above results further exemplify the high performance and the time efficiency of our model.
+Appendix E Experimental Outputs
+E.1 LLM Evaluation
 Table 9: ParaDetox LLM Evalution (LLM-t)
@@
 Take a deep breath and think step-by-step.
-C.2 Tables
+E.2 Tables
 We have appended the performance metrics of LLMs for reference.
@@
 Table 14: QG Dataset Generation
-Appendix D Well-Generated Output Examples
-D.1 QG
-Table 15: QG generation examples
-QG Generation Output Examples of Diffusion-EAGS
-Constraint $Y$ “Besides being able to hover in place, the hummingbird can also fly backwards.”
-Output $X_{1}$ “What kind of bird can fly backwards?”
-Output $X_{2}$ “Which bird is able to fly backwards?”
-Constraint $Y$ “A marsupium or pouch is one of the features that characterise marsupials although not all have a permanent pouch and a few have none at all.”
-Output $X_{1}$ “What is a pouch?”
-Output $X_{2}$ “What is the smallest animal without a pouch.”
-D.2 QQP
-Table 16: QQP generation examples
-QQP Generation Output Examples of Diffusion-EAGS
-Constraint $Y$ “What are the ten best short stories written by Isaac Asimov?”
-Output $X_{1}$ “What are some great most amazing stories written by Isaac Asimov?”
-Output $X_{2}$ “What are the best known fiction and books of Isaac Asimov?”
-Constraint $Y$ “Can we ever store energy produced in lightning?”
-Output $X_{1}$ “How do we store heat energy from lightning?”
-Output $X_{2}$ “How can you store energy from lightning?”
-D.3 ParaDetox
+Appendix F Well-Generated Output Examples
+F.1 ParaDetox
 Table 15: ParaDetox generation examples
@@
 Table 16: As the dataset is designed to eliminate problematic words while preserving semantics, our model adheres to these guidelines.
-D.4 Deontology
+F.2 Deontology
 Table 17: Deontology generation examples
@@
 Output $X_{2}$ “So I have to rent it to others.”
-D.5 RocStories
+F.3 QQP
+Table 18: QQP generation examples
+QQP Generation Output Examples of Diffusion-EAGS
+Constraint $Y$ “What are the ten best short stories written by Isaac Asimov?”
+Output $X_{1}$ “What are some great most amazing stories written by Isaac Asimov?”
+Output $X_{2}$ “What are the best known fiction and books of Isaac Asimov?”
+Constraint $Y$ “Can we ever store energy produced in lightning?”
+Output $X_{1}$ “How do we store heat energy from lightning?”
+Output $X_{2}$ “How can you store energy from lightning?”
+F.4 QG
+Table 19: QG generation examples
+QG Generation Output Examples of Diffusion-EAGS
+Constraint $Y$ “Besides being able to hover in place, the hummingbird can also fly backwards.”
+Output $X_{1}$ “What kind of bird can fly backwards?”
+Output $X_{2}$ “Which bird is able to fly backwards?”
+Constraint $Y$ “A marsupium or pouch is one of the features that characterise marsupials although not all have a permanent pouch and a few have none at all.”
+Output $X_{1}$ “What is a pouch?”
+Output $X_{2}$ “What is the smallest animal without a pouch.”
+F.5 RocStories
 Table 20: RocStories generation examples
@@
 Output $X_{2}$ “So he got a brand new one after all.”
-Appendix E Details on Text Augmentation Using GPT models
-E.1 GPT-3.5turbo ~ GPT-4-Omni
+Appendix G Details on Text Augmentation Using GPT models
+G.1 GPT-3.5turbo ~ GPT-4-Omni
 We prompt the GPT models to carry out dataset augmentation.
@@
 TARGET:
-Appendix F Details on CDLMs
-F.1 Experimental Details
+Appendix H Details on CDLMs
+H.1 Experimental Details
 For the case of Diffuseq and Dinoiser Ye et al.
@@
 Note that unlike other benchmarks, we experimented with Diffuseq-v2 (Gong et al., 2023b) in translation task for a broader comparison with existing baselines.
-F.2 Results Interpretations
+H.2 Results Interpretations
 Examining the results of Diffuseq, it is evident that the grammar score is comparatively lower than that of other models.
@@
 This finding, coupled with our additional experiments concerning the beam size during Dinoiser generation, suggests that Dinoiser’s performance predominantly relies on memorization.
-Appendix G Details on Translation Results
-G.1 Comparison Between Easy-First and Our Proposed Method
+Appendix I Details on Translation Results
+I.1 Comparison Between Easy-First and Our Proposed Method
 Discrete diffusion can be said to inherit ideas from NAR inference algorithm Mask-PredictGhazvininejad et al.
@@
 (2019b) on the original DisCo architecture implementation as baselines on translations tasks in Table 6 to further elucidate the difference through empirical results.
-G.2 Experimental Details
+I.2 Experimental Details
 NAR Transformer & CMLM We utilized the official repository to produce obtain the results, with the default architecture, optimization, and inference configurations.
 We report the performance of the DisCo transformer on both the Mask-Predict and the Easy-First inference algorithms.
-Diffuseq-v2 For Diffuseq-v2, we employ the vocabs of mBERT and choose 128 as max length for ende translation. Other settings are identical as in Appendix F.1.
+Diffuseq-v2 For Diffuseq-v2, we employ the vocabs of mBERT and choose 128 as max length for ende translation. Other settings are identical as in Appendix H.1.
 SEDD The SEDDLou et al.
-(2024) model, originally designed for open-ended text generation, was adapted in this study to facilitate dataset-guided generation.
+(2024) model, originally designed for open-ended text generation, was adapted in this study to facilitate conditional generation.
 To align the model’s architecture with the specific requirements of the structured dataset, several modifications were implemented in both hyperparameters and preprocessing protocols.
@@
 DiffusionEAGS For our model, we adopt the denosing strategy as top1 sampling and 1 size of MBR as typical translation task focuses on BLEU and COMET rather than diversity score.
-G.3 Experimental Results
-G.3.1 NAR Transformer, DisCo
+I.3 Experimental Results
+I.3.1 NAR Transformer, DisCo
 The results indicate that the DisCo transformer performs poorly on low-resource translation tasks, where the size of the dataset is small.
@@
 The performance benefits slightly from this change, but still lags behind those of other models.
-G.3.2 Diffuseq-v2
+I.3.2 Diffuseq-v2
 It is notable that existing diffusion language models perform poorly on translation tasks. In this section, we introduce some observations that might aid our understanding of such behaviors.
@@
 This analysis suggests that a simple rounding approach from continuous to discrete space may be insufficient for machine translation, at least in low-resource settings.
-Appendix H Entropy Flow
+Appendix J Entropy Flow
 Figure 4: Entropy behavior tracking in generation/training process.
@@
 The changes of entropy during the generation process in Diffusion-EAGS, represented by the yellow line, show that our model effectively follows a gradual decrease in entropy, mirroring the inverse trend of the training process.
-This gradual change in entropy facilitates successful DDLM training, which results in superior text quality performance compared to other diffusion models, as demonstrated in Tables 2, 3, and 4.
+This gradual change in entropy facilitates successful DDLM training, which results in superior text quality performance compared to other diffusion models, as demonstrated in Tables 3, 2, and 4.
 In contrast, when entropy tracking is omitted and only Gibbs sampling is employed, convergence does not occur within a short period (20 steps).

### v3

@@
-PLM-Based Discrete Diffusion Language Models with Entropy-Adaptive Gibbs Sampling
+Conditional [MASK] Discrete Diffusion Language Model
 Abstract
-Recently, discrete diffusion language models have demonstrated promising results in NLP.
-However, there has been limited research on integrating Pretrained Language Models (PLMs) into discrete diffusion models, resulting in underwhelming performance in downstream NLP generation tasks.
-This integration is particularly challenging because of the discrepancy between step-wise denoising strategy of diffusion models and single-step mask prediction approach of MLM-based PLMs.
-In this paper, we introduce Diffusion-EAGS, a novel approach that effectively integrates PLMs with the diffusion models.
-Furthermore, as it is challenging for PLMs to determine where to apply denoising during the diffusion process, we integrate an entropy tracking module to assist them.
-Finally, we propose entropy-based noise scheduling in the forward process to improve the effectiveness of entropy-adaptive sampling throughout the generation phase.
-Experimental results show that Diffusion-EAGS outperforms existing diffusion baselines in downstream generation tasks, achieving high text quality and diversity with precise token-level control.
-We also show that our model is capable of adapting to bilingual and low-resource settings, which are common in real-world applications.
-PLM-Based Discrete Diffusion Language Models with Entropy-Adaptive Gibbs Sampling
+Although auto-regressive models excel in natural language processing, they often struggle to generate diverse text and provide limited controllability.
+Non-auto-regressive methods could be an alternative but often produce degenerate outputs and exhibit shortcomings in conditional generation.
+To address these challenges, we propose Diffusion-EAGS, a novel framework that integrates conditional masked language models into diffusion language models through the theoretical lens of a conditional Markov Random Field.
+In doing so, we propose entropy-adaptive Gibbs sampling and entropy-based noise scheduling to counterbalance each model’s shortcomings.
+Experimental results show that Diffusion-EAGS outperforms baselines and achieves the best quality-diversity tradeoff, demonstrating its effectiveness in non-autoregressive text generation.
+Conditional [MASK] Discrete Diffusion Language Model
 […]
 1 Introduction
-Type Dataset OpenWebText Gokaslan and Cohen (2019) RocStories Mostafazadeh et al.
-(2016) Deontology Hendrycks et al.
-(2023) Question Generation Dhingra et al.
-(2017) QQP Wang et al.
-(2017) ALMA Xu et al.
-(2024a) ParaDetox Logacheva et al.
-(2022)
-Open-ended Generation $\checkmark$ $\checkmark$ $\triangle$ $\checkmark$ $\times$ $\times$ $\times$
-Conditional Generation $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-– Context Provided ? - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-– Content Provided ? - $\times$ $\triangle$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-– Format Provided ? - - $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-Table 1: Each dataset has a different level of conditional contraints even if they are conditional generation tasks.
-$\checkmark$ indicates full support, $\times$ indicates no support, and $\triangle$ indicates partial or limited support.
-As diffusion models significantly enhance the quality and diversity of generated outputs in continuous domains such as images and audio (Song et al., 2021b), recent research has increasingly applied diffusion models to NLP, and demonstrate their promising performance over autoregressive models (Li et al., 2022; Gong et al., 2023a; He et al., 2023; Yuan et al., 2023; Lovelace et al., 2023; Chen et al., 2023; He et al., 2023; Lou et al., 2024; Zhou et al., 2024; Shi et al., 2024; Sahoo et al., 2024; Zheng et al., 2024; Nie et al., 2024).
-Moreover, these models have been shown to be scalable, and are able to address the challenge of generalizing bidirectional relationships while maintaining sensitivity to the temporal alignment between training and test data, a conventional architectural limitation often observed in autoregressive models (Nie et al., 2024).
-Although recent Discrete Diffusion Language Models (DDLMs) have demonstrated high performance in NLP tasks such as unconditional and open-ended generation compared to Continuous Diffusion Language Models (CDLMs) (Lou et al., 2024), existing DDLMs continue to face limitations in conditional generation, thereby restricting their broader applicability across diverse NLP domains.
-Our experiments suggest that models such as SEDD falls short in such tasks.
-To enhance performance of DDLMs, PLMs can be integrated into diffusion models.
-This approach leverages the capability of PLMs to elicit improved generalization capabilities in unseen tasks while simultaneously exploiting the controllability and diversity strengths inherent to diffusion models.
-However, integrating PLMs into diffusion models is non-trivial as PLMs typically predict masked elements in a single step, whereas diffusion models require step-wise denoising based on the overall semantics of each timestep sequence, and such gap yields limited results.
-Therefore, we need to consider such inconsistencies to effectively adopt PLMs into DDLMs by a new methodology.
-In this paper, we introduce Diffusion-EAGS, a novel approach that effectively integrates Mask Language Model (MLM)-based PLMs with DDLMs for conditional generation.
-To address the gap between the step-wise nature of diffusion models and the one-step prediction strategy of PLMs, we begin by adopting PLMs as the denoising function at each step of the diffusion process.
-Since identifying where to apply denoising during the diffusion process is challenging for PLMs in our experiments, we incorporate an entropy tracking module to support their operation based on entropy information.
-To train DDLMs to effectively exploit an entropy tracking module, we propose an entropy-based noising scheduling during training phase in a forward process.
-Specifically, our entropy-based noising scheduling noises lower entropy tokens first, thereby learning to progressively generate sequences guided by the entropy from entropy tracking module.
-Experimental results demonstrate that Diffusion-EAGS achieves outstanding performance compared to baseline models across various conditional generation tasks.
-Furthermore, our model exhibits higher diversity in certain tasks compared to LLMs and can facilitate token-level generation, and validate that our model can also adapt to both bilingual and low-resource settings, which are frequently encountered in practical applications, indicating the potential applicability of our model across a wide range of conditional generation tasks.
-2 Task Setting
-2.1 Fine-Grained Conditional Generation
-In conditional generation tasks, the level of conditional constraint imposed by the dataset plays a critical role in shaping the generation process.
-As shown in Table 1, conditional constraints are diverse across datasets.
-In our task, we categorize these constraints into three levels: (1) the provision of context alone, requiring the continuity of the prefix; (2) the provision of specific content to be included in the target sequence, necessitating the inclusion of certain keywords; and (3) the provision of semantic content formatting, such as transforming toxic sentences into safer alternatives or converting text from the source language to a target language.
-In our study, we aim to develop a universal diffusion framework capable of being applied across a wide range of conditional generation tasks.
-2.2 DDLM with BERT
-BERT as Markov Random Field (MRF) As the pretraining paradigm has significantly contributed to the success of language models, the integration of PLMs is essential in the development of language models.
-Nevertheless, adopting PLMs to DDLMs is a complex task, as DDLMs require a step-wise denoising strategy, while PLMs are trained to predict masked elements in a single step.
-To effectively integrate PLMs into DDLMs, it is necessary to resolve such a discrepancy.
-Therefore, we begin by interpreting the intrinsic characteristics of MLMs as Markov Random Field (MRF), as demonstrated in Wang and Cho (2019).
-Details are in Appendix B.
-constrained MRF (cMRF) Even though open-ended MLMs are MRFs, we observe a significant increase in log-potential values for sequences when guided by the RocStories conditions, as shown in Figure 1.
-Additional experiments supporting this observation are detailed in Appendix A.
-Figure 1: When a condition is provided, the distribution of potential values for the samples is shifted on a logarithmic scale.
-This suggests that conditional generation models differ from open-ended models in randomness, making it crucial to investigate whether conditionally constrained spaces adhere to MRF properties.
-Through the derivation process outlined in Appendix B, we demonstrate that these spaces retain the characteristics of an MRF.
-Given a sequence of fully connected random variables $X=(x_{1},x_{2},\ldots,x_{L})$ with a set of observed variables $Y=(y_{1},y_{2},\ldots,y_{K})$, $f_{\theta}(X\backslash\{x_{l}\};Y)$ can be represented by logits from MLM given the sequence $X\backslash\{x_{l}\}$ and the observed variables $Y$.
-Then, log-potential can be calculated as:
-$\log\phi_{l}(X;Y)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\};Y),&\text{if }\text{[MASK]}\notin X_{l-1:l+1},\\ \\ 0,&\text{otherwise}\end{cases}$
-where $1h(x_{l})$ is a one-hot vector with $x_{l}$ set to 1. By this, we can define the potential function for the clique as:
-$\phi(X;Y)=\exp\left(\sum_{l=1}^{L}\log\phi_{l}(X;Y)\right)$ (1)
-Then, the constrained probability of sequence X is:
-$p_{\theta}(X;Y)=\frac{1}{Z(Y,\theta)}\prod_{l=1}^{L}\phi_{l}(X;Y)$
-Using the softmax function, the distribution of a variable $x_{l}$ given the observed variables $Y$ is:
-$p(x_{l}|X\backslash\{x_{l}\},Y)=\frac{\exp(1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\};Y))}{\sum_{m=1}^{M}\exp(1h(m)^{T}f_{\theta}(X\backslash\{x_{l}\};Y))}$ (2)
-This formulation implies that the dataset-constrained generation process in BERT can still be interpreted as a MRF with constraint Y, where the log-potential functions can be derived from the output logits of MLM.
-3 Methodology
-Figure 2: Given a conditional constraint $Y$ and a generation target $X$, the training process utilizes Entropy-based Sampling (ENS) to determine which positions in the masked sequence should be denoised when transitioning from one timestep to the next.
-The tokens at the corresponding masked positions ([M] denotes masked tokens) are then generated using $f_{\theta}$.
-The loss is subsequently computed, and the parameters are updated accordingly.
-During the inference process, the sequence $X$, initially composed of mask tokens, is iteratively refined through the diffusion process to generate $X$ that satisfies the given condition $Y$.
-In this section, we propose Diffusion-EAGS that integrates PLM into DDLMs through Entropy-based Adaptive Gibbs Sampling (EAGS) and Entropy-based Noise Scheduling (ENS).
-The overview of our process is in Figure 2.
-3.1 Generation Process: Entropy-based Adaptive Gibbs Sampling (EAGS)
-As discussed in Section 2.2, MLM-based PLMs can be interpreted as cMRFs, allowing Gibbs sampling to be applied at each denoising step in DDLMs by leveraging the properties of MRFs.
-As PLMs often struggle to select the next denoising targets, it is important to highlight that the primary purpose of adopting Gibbs sampling is not only to bridge the gap between PLMs and diffusion models but also to assist PLMs in identifying which absorbing tokens should be denoised.
-Therefore, we propose Entropy-Adaptive Gibbs Sampling (EAGS) to address the challenge PLMs face in identifying the appropriate denoising targets.
-In EAGS, the entropy of each variable quantifies the amount of uncertainty at a specific position within the sequence.
-This enables sampling of each token $x_{l}$ from its constrained distribution $p(x_{l}|X\backslash\{x_{l}\};Y)$ in the descending order of entropy.
-By prioritizing the generation of the least informative parts of the sequence, EAGS facilitates the creation of more structured sequences by leveraging the syntactic context that has already been established.
-Specifically, according to Equation (2), the probability $p^{t}(x_{l}=v_{m}|Y,f_{\theta})$ of the token $v_{m}$ given the context $Y$ and $f_{\theta}$ at step $t$ can be calculated as:
-$p^{t}(x_{l}=v_{m}|Y,f_{\theta})=\frac{\exp(f_{\theta}(v_{m}|X^{t}))}{\sum_{m^{\prime}=1}^{M}\exp(f_{\theta}(v_{m^{\prime}}|X^{t}))}$
-, where $f(v_{m}|X^{t})$ is the logit corresponding to the token $v_{m}$ for the position $x_{l}$ at step $t$.
-That is, we can compute the conditional probability for each masked position.
-With the conditional probability, the entropy $H$ for a variable $x_{l}$, logits function $f_{\theta}$, and the dataset context $Y$ is derived as:
-$\begin{split}&H(x_{l}|Y,f_{\theta})=\\ &-\sum_{m=1}^{M}p(x_{l}=v_{m}|x_{l},Y,f_{\theta})\log p(x_{l}=v_{m}|x_{l},Y,f_{\theta})\end{split}$
-With $H$, our approach for T step-generation process can be formalized as follows:
+Figure 1: Overview of how our approach (Diffusion-EAGS) combines the strengths of MLM and diffusion-based models to overcome the limitations of AR models, achieving a better diversity-quality tradeoff and fine-grained controllability
+Auto-Regressive Models (ARMs) have driven significant advances in NLP Achiam et al.
+(2023); Dubey et al.
+(2024); Team et al.
+(2023), yet they still face fundamental challenges such as diversity and controllability due to the ARM’s innate inductive bias.
+To address these challenges, a more flexible generative model is required.
+Specifically, ARMs face multiple challenges: they struggle to correct mathematical reasoning errors once made Wang et al.
+(2025), and often fail to integrate external knowledge Hudecek and Dusek (2023); Sun et al.
+(2023); Su et al.
+(2024).
+These shortcomings arise from ARMs’ sequential nature, which prevents them from revising earlier steps.
+As a result, they cannot effectively foster diversity through temperature-based sampling alone Lee et al.
+(2025), nor can they anticipate future requirements at earlier steps, thus undermining controllability when specific keywords must appear later Lu et al.
+(2022).
+One promising alternative is non-autoregressive generation, including Conditional Masked Language Models (CMLMs) Ghazvininejad et al.
+(2019a); Kasai et al.
+(2020) and diffusion models.
+CMLMs provide strong contextual understanding but lack an effective text generation mechanism.
+Meanwhile, diffusion models iteratively refine text through denoising, enabling fine-grained control and increased diversity.
+Recent works explore direct diffusion-based generation Li et al.
+(2022); Gat et al.
+(2024); The et al.
+(2024); Ye et al.
+(2025) or hybrid approaches combining diffusion with PLMs and LLMs (Lin et al., 2023; Xiang et al., 2024).
+However, despite their advantages, Discrete Diffusion Language Models (DDLMs) still suffer from degeneration in conditional generation tasks Xu et al.
+(2025), as confirmed by our experiments.
+We therefore propose Diffusion-EAGS, a novel approach that integrates CMLMs into DDLMs to achieve diverse, controllable, and high-quality conditional generation.
+However merging these methods is challenging because CMLMs generate text in one step by predicting all masked tokens, whereas diffusion models iteratively refine representations over multiple steps by introducing and removing noise.
+Our approach bridges this gap by leveraging a conditional Markov Random Field (cMRF) formulation, which enables:
   1. 1.
-    Entropy Calculation: Compute the entropy $H(x_{l}\mid Y,f_{\theta})$ for each variable $x_{l}$.
+    Stepwise iterative generation, overcoming the single-step limitations of CMLMs.
   2. 2.
-    Variable Selection: Select the variable $x_{l^{*}}$ with the highest entropy for sampling
-    $l^{*}=\arg\max_{l}H(x_{l}\mid Y,f_{\theta})$
+    Stable and diverse conditional text generation, reducing semantic drift in DDLMs.
+Diffusion-EAGS achieves this through two key methodologies:
+  • Entropy-Adaptive Gibbs Sampling (EAGS): A strategy that updates the most uncertain (high-entropy) tokens first at each denoising step, ensuring qualified generation.
+  • Entropy-based Noise Scheduling (ENS): A training approach that progressively masks tokens based on ascending order of entropy, enabling the model to learn a structured denoising process.
+We conduct extensive experiments to validate Diffusion-EAGS on various conditional generation tasks, demonstrating significant improvements over baseline models.
+Our approach achieves the best quality-diversity tradeoff, demonstrating that Diffusion-EAGS balances fluency and variability more effectively than existing models.
+Moreover, keyword-based story generation experiments confirm that our model effectively generates coherent and controlled text from randomly masked sequences, making it highly adaptable to different conditioning constraints.
+2 Related Works
+Efforts to integrate generative flow models into sequence generation exploit the distribution shift from a source language to a target language through a series of invertible linear transformations (Ma et al., 2019; Zhang et al., 2024).
+However, as DDPM (Ho et al., 2020a) demonstrate the effectiveness of generating images, diffusion models have been a major topic of interest within the field of generative flow models (Song et al., 2021a, b).
+To apply such diffusion methodologies to NLP, in order to leverage their strengths in controllability and diversity, recent studies have demonstrated promising performance across various tasks (Li et al., 2022; Gong et al., 2023a; He et al., 2023; Yuan et al., 2023; Lovelace et al., 2023; Chen et al., 2023; He et al., 2023; Lou et al., 2024; Zhou et al., 2024; Shi et al., 2024; Sahoo et al., 2024; Zheng et al., 2024; The et al., 2024; Wang et al., 2024).
+Although Continuous Diffusion Language Models (CDLMs) such as Diffusion-LM (Li et al., 2022), DiffuSeq-v1, v2 (Gong et al., 2023a, b), and LD4LG (Lovelace et al., 2023) show promising performance, Bansal et al.
+(2022) argue that such operations do not necessarily have to be governed by stochastic randomness.
+Building on this rationale, D3PM (Austin et al., 2023) propose the discrete restoration-generation approach and DiffusionBERT He et al.
+(2022) adopt pre-trained language models (PLMs) to DDLM.
+SEDD Lou et al.
+(2024) propose score entropy inspired by MLM loss, and outperform existing CDLMs.
+Recent works by Shi et al.
+(2024) and Sahoo et al.
+(2024) extend this idea and obtain better empirical results.
+Zheng et al.
+(2024) further enhance discrete diffusion models by correcting the numerical precision error in SEDD-based models.
+These research make an improvement on the open ended generation task.
+Furthermore, Venkatraman et al.
+(2024) use SEDD as text infilling, and Nie et al.
+(2024) demonstrate that DDLMs are scalable.
+3 MLM & DDLM : D-cMRF
+Pre-trained MLMs offer rich, context-aware representations through one-pass masked prediction, whereas DDLMs iteratively refine text via stepwise denoising to enhance control and diversity.
+Combining these approaches can overcome MLMs’ one-pass limitations and DDLMs’ degeneration in conditional generation.
+However, their integration is challenging because DDLMs require iterative updates while MLMs predict all masked tokens simultaneously.
+To bridge this gap, we propose Diffusion-based Constrained Markov Random Fields (D-cMRF), a framework that integrates a discrete diffusion process into MLM sequence generation.
+By leveraging an entropy-based sampling strategy to selectively update high-uncertainty tokens at each step, D-cMRF achieves a principled reduction in sequence energy, leading to stable and coherent generation.
+3.1 MLM as cMRF
+Inspired by the traditional approaches of Wang and Cho (2019) and Goyal et al.
+(2022), which model MLMs as Markov Random Fields (MRFs) and energy-based models (EBMs), respectively, we reinterpret MLM as a conditional MRF (cMRF) model and employ it as a denoising function at each diffusion step.
+Let $X=(x_{1},x_{2},\dots,x_{L})$ be a sequence of discrete variables from a vocabulary $V$, with $Y$ representing observed conditions.
+The sequence probability follows an energy-based MRF formulation:
+$P_{\theta}(X;Y)=\frac{\exp(-E_{\theta}(X;Y))}{Z(Y,\theta)}$ (1)
+where $E_{\theta}(X;Y)$ is the energy function parameterized using MLM-based embeddings,  $\theta$ denotes parameterization of MLM, and $Z(Y,\theta)$ is the partition function for ensuring proper normalization.
+Then the total sequence energy is defined as:
+$E_{\theta}(X;Y)=-\sum_{l=1}^{L}\log\phi_{l}(X;Y)$ (2)
+where log-potential function $\log\phi_{l}(X;Y)$ is :
+$\log\phi_{l}(X;Y)=1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\};Y)$ (3)
+where $l$ is a token position in the sequence, $1h(x_{l})$ is the one-hot encoding of token $x_{l}$, and $f_{\theta}(X\backslash\{x_{l}\};Y)$ represents the MLM logit output conditioned on the sequence.
+3.2 DDLM with Entropy-based Denoising
+Determining how to perform sampling with such a simple cMRF presents a separate challenge.
+In particular, one can use techniques such as Gibbs sampling as long as the energy space remains unchanged, but we cannot guarantee that this energy space is stable in general Goyal et al.
+(2022).
+The necessity of generating sequences in cMRF based on energy update is in Appendix A.
+Hence, a natural research question arises: “How should we sample and update the energy?”
+The training process of diffusion models (both forward and backward) conceptually represents the entire distribution as a product of local conditional distributions across time steps.
+Hence, diffusion models share a probabilistic graphical structure with MRF, enabling MLM to be integrated within DDLM framework.
+Therefore, in this subsection, we describe how to update the energy and perform sampling under the DDLM framework using $P_{\theta}(X;Y)$.
+Specifically, we integrate $P_{\theta}(X;Y)$ into each diffusion step as a denoising function, employing an entropy-based denoising matrix $Q$ in Section 4.2.
+We first define the entropy of each token:
+$H_{i}(X^{(t)})=-\sum_{x^{\prime}\in V}p_{\theta}(x^{\prime}_{i};X^{(t)})\log p_{\theta}(x^{\prime}_{i};X^{(t)})$ (4)
+where $p_{\theta}(x^{\prime}_{i};X^{(t)})$ is the softmax probability of token $x^{\prime}_{i}$ at position $i$ in sequence $X^{(t)}$, and $t$ denotes the diffusion timestep.
+$M_{t}=\{i\mid H_{i}(X^{(t)})\geq\tau_{t}\}$ (5)
+where $\tau_{t}$ is a dynamically adjusted entropy threshold.
+This ensures that updates occur at positions where the model has the highest uncertainty.
+Subsequently, we sample the next-step sequence from $P_{\theta}(X^{(t)};Y)$ at the suggested positions.
+We perform this selection process at every diffusion step, which corresponds to updating the energy, different from existing research Wang and Cho (2019); Goyal et al.
+(2022).
+3.3 D-cMRF
+By combining DDLM with cMRF, our approach enables a theoretically grounded generation process from the perspective of MLM.
+Moreover, from the diffusion standpoint, the training process naturally aligns with MLM objective, as discussed in Section 3.1 and Section 3.2.
+Specifically, our D-cMRF guarantees energy reduction during generation, ensuring stable sequence reconstruction.
+Figure 2: Overview of the training (forward) and inference (backward) processes in Diffusion-EAGS.
+Training (left): Entropy-based Noise Scheduling (ENS) determines which tokens in the masked sequence, denoted by $[M]$, should be denoised at each timestep based on the position entropy $H(x_{i})$.
+These tokens are then generated using the diffusion model with parameters $\theta$, and the loss is computed using a cross-entropy (C.E.) diffusion loss.
+Inference (right): Starting from a fully masked sequence conditioned on $Y$, Entropy-Adaptive Gibbs Sampling (EAGS) iteratively refines the sequence by focusing on high-entropy tokens, denoted as $M_{t}$, based on a threshold $\tau_{t}$, yielding stable and coherent text generation.
+Step 1: Expected Energy at Diffusion Step $t$
+At diffusion step $t$, the expected sequence energy is defined as follows:
+$\mathbb{E}_{X^{(t)}\sim q}\left[E_{\theta}(X^{(t)};Y)\right]=\sum_{X^{(t)}}q(X^{(t)})E_{\theta}(X^{(t)};Y)$ (6)
+where $q(\cdot)$ denotes the probability distribution from which $X^{(t)}$ is sampled. Since high-entropy tokens are selected for replacement, the total sequence energy can be decomposed as follows:
+$\mathbb{E}\bigl{[}E_{\theta}(X^{(t)};Y)\bigr{]}$ $=\sum_{i\in M_{t}}\mathbb{E}\bigl{[}E_{\theta}(x_{i}^{(t)};X^{(t-1)},Y)\bigr{]}$ (7)
+$\quad+\sum_{i\notin M_{t}}E_{\theta}\bigl{(}x_{i}^{(t-1)};Y\bigr{)}.$
+Step 2: Energy Reduction via Denoising
+Since masked tokens are replaced with lower-energy candidates at each step, we expect a general trend of energy reduction.
+However, due to the stochastic nature of sampling, local fluctuations in energy may occur.
+Over multiple diffusion steps, the entropy-based selection mechanism ensures a net decrease in sequence energy.
+$\mathbb{E}\left[E_{\theta}(x_{i}^{(t)};X^{(t-1)},Y)\right]\leq E_{\theta}(x_{i}^{(t)};X^{(t)},Y)$ (8)
+Applying this property across all updated tokens $i\in M_{t}$, we obtain:
+$E_{\theta}(X^{(t-1)};Y)\leq E_{\theta}(X^{(t)};Y)$ (9)
+Step 3: Convergence to Low-Energy States
+Summing over all diffusion steps $T$:
+$E_{\theta}(X^{(0)};Y)\leq E_{\theta}(X^{(T)};Y)$ (10)
+where $X^{(T)}$ is the fully masked sequence with maximum entropy, and $X^{(0)}$ is the final reconstructed sequence.
+Since the token space is discrete and energy is derived from a sum of bounded logits, $E_{\theta}(X;Y)$ is lower-bounded by a finite minimum energy state.
+While stochastic sampling may introduce fluctuations, the diffusion process ensures progressive energy minimization, leading to an approximate low-energy state.
+3.3.1 D-cMRF Guarantees
+The above proof establishes that our method satisfies the following properties:
+  • Progressive Energy Reduction: The energy function exhibits an overall decrease, leading to more stable sequence generation. This trend is supported by empirical results in Appendix D.
+  • Stable Convergence: Since the energy function is lower-bounded and sequence length is finite, the generation process is expected to reach a structured, low-entropy state.
+These properties explain the improved performance of Diffusion-EAGS compared to traditional diffusion models, as shown in §Section 6.
+Notably, the ablation study in Table 5 demonstrates that removing EAGS leads to a significant drop in performance, highlighting its importance in guiding stable generation.
+4 Diffusion-EAGS
+Our approach, Diffusion-EAGS, leverages two key components—Entropy-Adaptive Gibbs Sampling (EAGS) and Entropy-based Noise Scheduling (ENS)—rooted in the theory of Section 3.
+As shown in Figure 2, during training, ENS selectively masks tokens based on their certainty, while during generation, EAGS iteratively refines a fully masked sequence by updating high-uncertainty tokens.
+This stepwise refinement yields balanced improvements in text quality and diversity.
+4.1 Inference Process: Entropy-Adaptive Gibbs Sampling
+As discussed in Section 3.2, MLM can be interpreted as cMRF, which is used as $p_{\theta}$ in the sampling process of DDLM with $M_{t}$.
+In particular, $M_{t}$ is not only associated with energy updates but also serves as a solution to the MLM’s difficulty in selecting the next tokens to denoise, as shown in Appendix  C.
+Henceforth, we designate this strategy as Entropy-Adaptive Gibbs Sampling (EAGS).
+In EAGS, $M_{t}$ is constructed by ranking tokens in descending order of entropy, thereby prioritizing the least informative parts of the sequence.
+EAGS facilitates the creation of more structured sequences by leveraging the syntactic context that has already been established.
+The process of determining the denoising schedule is shown in Appendix C.
+Our approach for T step-generation process can be formalized as follows:
+  1. 1.
+    Entropy Calculation: Compute the entropy $H_{i}(X^{(t)})$ for each variable $x_{i}$.
+  2. 2.
+    Variable Selection: Obtain $M_{t}$ for sampling
   3. 3.
-    Sampling: Sample $x_{l^{*}}$ from its conditional distribution $p(x_{l^{*}}\mid Y,f_{\theta}^{*})$.
+    Sampling: Sample $x_{i^{*}}$ from its conditional distribution $p_{\theta}(x_{i^{*}}\mid X^{(t)},Y)$, where $i^{*}\in M_{t}$.
   4. 4.
@@
   5. 5.
-    Iteration: Repeat Steps 1 through 4 until t=T.
-3.2 Forward Process: Entropy-based Noise Scheduling (ENS)
-To enhance the facilitation of Adaptive Gibbs Sampling in the generation process, it is essential to simulate a similar process in the training phase.
-Therefore, we schedule the forward process of diffusion training based on the entropy $H(x_{l})$ of position $x_{l}$ with the input sequence $[Y|X]$.
-Specifically, the position-wise entropy is calculated by the PLM that shares a similar language base with diffusion-trained model.
-Assuming the diffusion process progresses over $T$ steps, at each step $t$, we mask the $L/T$ number of positions with the lowest entropy from the set $\{x_{1},\ldots,x_{L}\}$.
-The masking process at step $t$ in position $l$ is described by the denoising matrix $Q_{tl}$.
-$Q_{tl}=\begin{bmatrix}q_{11}&0&\cdots&0&q_{1,M}\\ 0&q_{22}&\cdots&0&q_{2,M}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&q_{M-1,M-1}&q_{M-1,M}\\ 0&0&\cdots&0&q_{MM}\\ \end{bmatrix}$
-, where
-$q_{mn}=\begin{cases}q_{mm}=1&\text{if }x_{l}\notin\text{MIN}([H(x_{1}),\cdots,H(x_{L})],\frac{L}{T})\\ q_{mM}=1&\text{if }x_{l}\in\text{MIN}([H(x_{1}),\cdots,H(x_{L})],\frac{L}{T})\\ 0&\text{otherwise}\end{cases}$
-This entropy-based noise scheduling approach ensures that the forward process in diffusion training closely mirrors the generation process, thereby enhancing the effectiveness of EAGS in language generation.
-3.3 Training and Inference
+    Iteration: Repeat Steps 1 through 4 until $t=T$, where $T$ is the total number of timestep.
+The detailed algorithm of EAGS is in Appendix Algorithm 1.
+4.2 Training Process: Entropy-based Noise Scheduling
+To improve the effectiveness of EAGS during generation, we simulate a similar process during training.
+Therefore, we schedule the forward process of diffusion training based on the entropy $H_{i}(X^{(t)})$ of position $x_{i}$ with the input sequence $[Y|X^{(t)}]$ at sampled timestep $t$.
+During training, $H_{i}(X^{(t)})$ is calculated by pre-trained MLM.
+Assuming the diffusion process progresses over $T$ steps, we mask the $L/T$ number of positions with the lowest entropy from the set $\{x_{1},\ldots,x_{L}\}$ at each step $t$, where $L$ is the sequence length.
+This selection process is used to determine $\tau_{t}$ in Equation 5.
+The masking process at step $t$ in position $i$ is described by the denoising matrix $Q_{ti}$.
+$Q_{ti}=\begin{bmatrix}q_{11}&0&\cdots&0&q_{1,M}\\ 0&q_{22}&\cdots&0&q_{2,M}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&q_{M-1,M-1}&q_{M-1,M}\\ 0&0&\cdots&0&q_{MM}\\ \end{bmatrix}$
+Here, $q_{1,M}$ denotes the transition probability from the vocab index corresponding to token 1 to the [MASK] token and $q_{mn}$ is defined as:
+$q_{mn}=\begin{cases}q_{mm}=1&\text{if }x_{i}\notin\text{MIN}([H_{1}(x_{1}),\cdots,H_{L}(x_{L})],\frac{L}{T})\\ q_{mM}=1&\text{if }x_{i}\in\text{MIN}([H_{1}(x_{1}),\cdots,H_{L}(x_{L})],\frac{L}{T})\\ 0&\text{otherwise}\end{cases}$
+Henceforth, we designate this strategy as Entropy-based Noise Sampling (ENS).
+ENS masks lower entropy tokens first, thereby learning to progressively generate sequences.
+This ensures that the forward process in diffusion training closely mirrors the generation process, thereby enhancing the effectiveness of EAGS in language generation.
+The detailed algorithm of ENS is in Appendix Algorithm 2.
+4.3 Diffusion Loss with Cross Entropy
+Distinct from the prevailing methodologies in diffusion models Ho et al.
+(2020a); Austin et al.
+(2023), we do not employ the PLM parameterization $\widetilde{p}_{\theta}(\widetilde{z}_{0}|z_{t},t)$, which preserves the original semantic embedding spaces during the training phase as we empirically find that such method restricts the diversity of generated responses.
+We follow the traditional diffusion loss (Ho et al., 2020b), changing Mean Squared Error with Cross Entropy Loss.
+5 Experiments
+5.1 Tasks & Details
+We conduct experiments on various conditional generation datasets.
+Detailed explanation of the conditional generation datasets are in Appendix F.1.
+In particular, we focus on two datasets that significantly differ in their level of conditional constraints: RocStories (Mostafazadeh et al., 2016), which is relatively open-ended, and Paradetox (Logacheva et al., 2022), which imposes the strongest conditional constraints.
+We select the conditional dataset that GPT-2 faces in generating sentences of appropriate length under specified conditional constraints.
+The maximum lengths of Paradetox and RocStories is set to 64, based on data statistics, and other details are in Appendix F.
+The number of steps of our model is configured to 5 with a naive categorical sampling with a sample size of 20 and select final 5 samples based on Perplexity score.
+We use 1 A100 GPU with the batch size as 256.
+5.2 Baselines
+We employ RoBERTa-base (Liu et al., 2020) as MLM with learning rate 5e-4.
+Next, we compare Diffusion-EAGS with four categories of baselines of similar size to RoBERTa-base: Auto-regressive Models (ARMs), Conditional Masked Language Models (CMLMs), Continuous Diffusion Language Models (CDLMs), and Discrete Diffusion Language Models (DDLMs).
+Note that our primary goal is to investigate the architecture’s capabilities; any baseline approach in the direction of scalability or bypassing the architecture’s limitations goes beyond our research scope.
+For ARMs (Vaswani et al., 2023), we employ GPT-2 (Radford et al., 2019) and GPT-3.5-turbo with four-shot prompt.
+More experimental details of GPT-3.5 can be found in Appendix J, and H.
+For CMLMs, we utilize CMLM-Mask-Predict (Ghazvininejad et al., 2019a) and DisCo-Easy-First (Kasai et al., 2020), which are transformer-based NAR models.
+For CDLMs, our baseline includes DiffuSeq (Gong et al., 2023a), LD4LG (Lovelace et al., 2023), and DINOISER (Ye et al., 2024).
+DiffuSeq and DINOISER is designed for sequence-to-sequence applications, and LD4LG adopt BART as denoising init point.
+For DDLMs, we utilize DiffusionBERT (He et al., 2022), applying BERT into DDLMs, and SEDD (Lou et al., 2024), a powerful open-ended generation DDLM.
+For SEDD, we download the pre-trained version and fine-tune it.
+5.3 Metrics
+Quality metrics
+In addition to our theoretically guided methods, we evaluate performance using multiple metrics.
+Specifically, we use Perplexity (PPL) based-on GPT-2 Large and GPT-2 XL as an automated metric, MAUVE Pillutla et al.
+(2021) to assess style consistency between the training data and generated outputs, SOME (Yoshimura et al., 2020) to score the grammar, Mean Opinion Score (MOS) from human evaluations to gauge text quality, and LLM score such as LLM-c (Lin and Chen, 2023) to measure the plausibility of the narratives as a sub-metric.
+Further details on LLM scores and other evaluation settings are provided in Appendix F.3 and H.1.
+Diversity Metrics
+Following our quality assessment, we evaluate diversity through three different measures: an automatic frequency-based metric n-gram Vendi Score(VS n-gram) Friedman and Dieng (2023), a neural network–based semantic metric SimCSE Vendi Score (VS emb), and a human evaluation score MOS.
+More details related to MOS and other scores are in Appendix F.3.
+6 Results
+Model Text Quality
+Step PPL $\downarrow$ MAUVE $\uparrow$ MOS $\uparrow$
+AR model
+GPT-2 1 389.1 0.503 0.83
+GPT-3.5 w/ 4-shot 1 104.375 0.175 1
+CMLMs
+CMLM w/ Mask-Predict 10 669.9 0.0234 -
+DisCo w/ Easy-First 10 716.1 0.0344 -
+Diffusion models
+DiffusionBERT 2000 775.9 0.737 0.88
+DiffuSeq 2000 $\geq 1k$ 0.683 -
+SEDD 1024 $\geq 1k$ NA -
+LD4LG 2000 579.9 0.556 0.91
+DINOISER 20 124.8 0.255 0.91
+Diffusion-EAGS 5 109.3 0.811 0.97
+Table 1: Text quality of conditional generation outputs.
+We report Perplexity (PPL) for sentence fluency, MAUVE for condition alignment, and Mean Opinion Score (MOS) for semantic coherence.
+Models with PPL exceeding 600 were excluded from human evaluation.
+Model Text Quality Diversity
+PPL $\downarrow$ SOME $\uparrow$ LLM-c $\uparrow$ VS(ngram) $\uparrow$ self-bleu $\downarrow$
+Original Data 100.6 0.895 1
+GPT-2 88.5 0.856 0.88 4.722 0.124
+DiffusionBERT 318.2 0.783 0.72 4.735 0.088
+SEDD 273.2 0.827 0.59 4.859 0.044
+Diffusion-EAGS 67.3 0.844 0.87 4.837 0.058
+Table 2: Results on the open-ended RocStories (ROC) dataset. We report perplexity (PPL) for fluency, SOME and LLM-c for text quality, and both VS(ngram) and self-BLEU for diversity.
+In Tables 1, 2, and 3, our model consistently demonstrates strong text quality and diversity compared to various baselines across a wide range of conditional generation tasks.
+Text Quality.
+Table 1 shows that our model achieves notable improvements in perplexity (PPL) and obtains high MAUVE and MOS scores, indicating that the generated texts are both fluent and coherent.
+Although GPT-3.5-turbo is capable of generating high-quality text, the MAUVE metric indicates that few-shot prompts alone are insufficient for accurately replicating the dataset’s inherent characteristics.
+On the other hand, CMLMs, DiffuSeq, and DINOISER can handle conditional constraints but sometimes struggle with semantic drift or high PPL.
+In contrast, Diffusion-EAGS achieves both lower PPL and strong human evaluation scores (MOS), suggesting that it effectively balances condition satisfaction with text quality.
+Table 2 further demonstrates our model’s capability on the open-ended RocStories dataset.
+Even with minimal constraints, Diffusion-EAGS maintains competitive scores compared to GPT-2, demonstrating its robustness in narrative generation.
+Model Diversity
+Step VS(ngram) $\uparrow$ VS(emb) $\uparrow$ MOS $\uparrow$
+AR model
+GPT-2 1 3.925 2.640 2.65
+GPT-3.5 w/ 4-shot 1 3.098 1.915 2.2
+CMLMs
+CMLM w/ Mask-Predict 10 1.000 1.000 -
+DisCo w/ Easy-First 10 1.000 1.000 -
+Diffusion models
+DiffusionBERT 2000 3.101 2.058 2
+DiffuSeq 2000 2.059 1.465 -
+SEDD 1024 4.746 4.063 -
+LD4LG 2000 1.914 1.425 1
+DINOISER 20 2.287 2.174 1
+Diffusion-EAGS 5 4.417 3.311 4.6
+Table 3: Diversity evaluation for generated outputs.
+We report the n-gram-based Vendi Score (VS(ngram)), the embedding-based Vendi Score (VS(emb)), and a Mean Opinion Score (MOS) for diversity.
+Higher values indicate greater diversity.
+Diversity.
+Diffusion-EAGS excels at generating diverse outputs.
+As illustrated in Table 3, our model consistently excels in both n-gram and embedding-based diversity metrics (VS(ngram) and VS(emb)), surpassing other baselines and even larger LLMs.
+The model’s higher MOS for diversity further indicates that humans also perceive its outputs to be more varied and engaging.
+In line with these observations, we conduct additional analyses (Appendix G.4) including the comparison ours with large LLMs, where our approach produces a wider range of coherent yet distinct responses.
+These findings underscore the effectiveness of our entropy-adaptive sampling strategy in avoiding repetitive outputs and semantic collapse, thereby delivering a superior quality-diversity trade-off.
+Overall, Diffusion-EAGS consistently demonstrates strong performance across diverse conditional generation tasks, combining low perplexity and high human evaluation scores with the ability to generate richly varied text.
+7 Analysis
+7.1 Quality-Diversity Tradeoff
+Balancing quality and diversity is a fundamental challenge in text generation.
+AR models typically achieve high fluency but suffer from low diversity, while non-autoregressive models, such as CMLMs and diffusion models, often struggle to generate coherent outputs.
+Our proposed Diffusion-EAGS effectively balances these factors by leveraging a structured diffusion process.
+Figure 3 presents the quality-diversity tradeoff among various models, where quality is measured using perplexity (PPL) on the x-axis (inverted as 1/1/1 /PPL for better visualization) and diversity is quantified using VS_emb on the y-axis.
+Our model (Ours_Deon, Ours_Para, marked with purple stars) achieves the best tradeoff, outperforming prior approaches in both high-quality generation and diversity.
+Figure 3: Quality–diversity tradeoff across various models.
+The x-axis ($1/\text{PPL}$) reflects generation quality, while the y-axis (VSemb) indicates diversity.
+Green points represent AR models, yellow points represent diffusion models, and blue points represent CMLMs.
+Our Diffusion-EAGS variants, marked by purple stars, achieve the best overall tradeoff.
+Compared to Diffuseq, DiffusionBERT and CMLMs, our method achieves significantly better diversity without compromising generation fluency.
+This improvement stems from our Entropy-Adaptive Gibbs Sampling (EAGS), which ensures controlled token selection, and Entropy-based Noise Scheduling (ENS), which stabilizes the generation process.
+The results highlight that integrating MLMs into the diffusion framework enables high-quality, diverse, and controllable text generation.
+7.2 Keyword Based Generation
+Context Jake was playing with his toys. He accidentally broke his favorite one.
+He cried a lot over it. His parents decided to replace it for him.
+Keyword not stop Jake just could not stop crying.
+Jake feel It made Jake feel So much better.
+would enjoy Jake said he would enjoy the new toy
+Context Neil was in Sofia, Bulgaria. He was enjoying a trip backpacking through Europe.
+… He thought the food and culture in Sofia were the best.
+Keyword Bulgaria! Things were looking great in Bulgaria!
+Context Karen wanted to go on a trip to France. She started doing research on the trip.
+She decided to book a week long trip. She left the next day for her tripsx.
+Keyword her trip She spent almost a week there during her trip.
+Table 4: Examples of keyword-based generation.
+Each row shows a Context and a specified Keyword, which is inserted into a masked position.
+The resulting outputs demonstrate how our model seamlessly integrates keywords into coherent narratives.
+Our model operating within discrete space enables us to manipulate the output sequences using explicit instructions.
+To further explore this capability, we conduct the generation of sequences based on keywords positioned in the middle and at the end of masked sequences, which is challenging for AR models (Keskar et al., 2019).
+They inherently struggle with controllability due to their inability to revise past steps based on future ones—an inductive bias of AR models.
+Initially, we provide the same contextual input while varying the keywords.
+In the masked states, we randomly select positions, replacing them with the specified keywords.
+The results in Table 4 demonstrate that the generated sequences seamlessly integrate the keywords with context-specific semantics.
+7.3 Ablation Study
+Dataset PPL MAUVE SOME VS(ngram) VS(emb)
+Diffusion-EAGS Deont 55.1 0.412 0.835 4.898 4.009
+Roc 67.3 0.87 0.844 4.837 3.999
+w/o EAGS Deont 667.9 0.022 0.617 4.767 3.928
+Roc 1084.9 0.035 0.613 4.874 3.957
+w/o Gibbs Sampling Deont 1426.7 0.011 0.584 2.378 1.923
+Roc 1293.1 0.010 0.534 1.531 1.338
+w/o Pre-trained MLM Deont $\geq$2K 0.005 0.645 4.758 3.402
+Roc $\geq$2K 0.004 0.604 4.315 2.994
+Table 5: Ablation study on the Deontology (Deont) and RocStories (Roc) datasets.
+“w/o EAGS” uses naive Gibbs sampling (no entropy estimation), “w/o Gibbs Sampling” removes diffusion process, and “w/o Pre-trained MLM” omits the pre-trained MLM entirely.
+To explore the effectiveness of our model’s components, we conduct ablation studies focusing on three key elements: EAGS, Gibbs Sampling, and pre-trained MLM in Table 5.
+The result of w/o EAGS shows a severe decline in text quality, producing degenerated results similar to those of traditional CMLMs.
+Such phenomenon suggests that the naive application of MLM within the diffusion process fails to fully harness the capabilities of it.
+Next, removing the use of the diffusion generation process (w/o Gibbs Sampling) leads to a drastic reduction in overall performance, with increased PPL and reduced diversity scores.
+These results imply that relying solely on MLM for text generation introduces considerable limitations.
+Without the pre-trained MLM, outputs become highly degenerated, underscoring the need for precise entropy estimation.
+In the process of selecting our highest-entropy-based scheduling in Diffusion-EAGS, we considered three alternatives: lowest entropy selection, random position selection following ENS training, and highest entropy selection.
+Experiment on the paradetox dataset yielded PPL scores of 1193, 183, and 112, respectively.
+A subsequent heuristic evaluation confirms that the quality aligns with these PPL values.
+Consequently, we adopt the highest-entropy-based selection strategy.
+The process of schedule selection is detailed in Appendix C.
+With EAGS, our model shows a substantial performance improvement.
+To verify the effectiveness of our model in guiding stable energy reduction, we examine the entropy flow during the generation process in Appendix D.
+Our findings demonstrate that EAGS contributes significantly to a gradual decrease in entropy, enabling the generation of sentences in a stable manner.
+8 Conclusions & Discussions
+In this work, we introduce Diffusion-EAGS, an approach that integrates MLMs with diffusion models for conditional generation, yielding improved text quality, enhanced diversity, broad applicability, and precise token-level control.
+Investigation of Other PLMs
+We conducted a toy experiment using T5 on the Paradetox dataset; however, the results showed no significant improvement over GPT-2 fine-tuning (see Appendix G.1, Table 15).
+We hypothesize that T5’s generation is heavily influenced by its initial decoder tokens Wang and Zhou (2024), which leads to lower diversity.
+This suggests that developing a theoretical framework to integrate encoder-decoder models with diffusion processes may be a promising direction for future research in conditional generation.
+By devising methodologies that align the training objectives of other PLMs with diffusion loss—similar to our approach—, we can further accelerate progress in diffusion-based NLP.
+Limitations
+While Diffusion-EAGS demonstrates significant improvements in conditional generation tasks, there are several limitations.
+First, as our method is currently focused on text generation tasks, its applicability to text classification tasks, such as Named Entity Recognition and Part-of-Speech Tagging, remains unexplored.
+Future research could explore extending this method to other NLP tasks.
+Second, although our current efforts concentrate on developing and validating our framework using MLM, the potential integration of ARMs remains unexplored.
+With a proper methodology that aligns AR pre-training and diffusion training objectives, AR models would be another good initialization.
+Third, although the bias embedded in pre-trained models can be directly propagated, recent studies show that data-balancing strategies can effectively address this issue.
+Consequently, it is essential to account for these factors when deploying such models.
+Appendix A Necessity of Energy Update in cMRF Generation
+We observe a significant increase in log-potential values for sequences when guided by the RocStories conditions, as shown in Figure 4.
+Additional experiments supporting this observation are detailed in Appendix B.
+Figure 4: When a condition is provided, the distribution of potential values for the samples is shifted on a logarithmic scale.
+This observation implies that conditional sequences differ from different conditional sequences in terms of randomness, making it crucial to update the energy function when the conditioning changes.
+For instance, MASK MASK author and MASK am author belong to different random fields, as also suggested by Goyal et al.
+(2022).
+Appendix B Measuring Potential Function in MLM
+In this section, we provide additional experimental details and results to support the observation that open-ended Masked Language Models (MLMs) exhibit increased potentials for the same sequence under different dataset constraints.
+B.1 Experimental Setup
+  • Model We use the pre-trained BERT large model (bert-large-cased) as the base model for all experiments.
+Additionally, we incorporate RocStories-conditioned guidance with the pre-trained model and use a fine-tuned BERT model on the RocStories dataset to evaluate the impact of dataset-specific constraints.
+  • Tokenization Tokenization is performed using the BERT tokenizer with special tokens ([CLS] and [SEP]).
+  • Potential Calculation The the log-potentials are obtained for each token using masked token logits.
+  • Datasets
+    • RocStories: Structured narratives from the RocStories dataset.
+B.2 Results of Experiment and Implications for Conditional Generation
+Using the BERT-large-cased model, the average log potential value for the standard MLM was 156.6150, while incorporating RocStories guidance increased this value to 175.5332, highlighting the impact of dataset-specific constraints.
+Additionally, fine-tuning the same model on RocStories resulted in an average potential function value of 3.7551 (on an exponential scale), demonstrating substantial variation introduced by conditional generation settings.
+The results demonstrate the significant influence of dataset structure on the potential function in MLMs.
+Specifically, structured datasets like RocStories enforce stronger narrative constraints, leading to higher potentials and greater coherence in sequence generation.
+Appendix C The Candidates of Denoising Schedules
+We arrived at our proposed approach by going through several steps. The core of DDLM lies in how to define the denoising matrix.
+1. Initial BERT Refinement Without a Noise Matrix
+We first explored a BERT-refinement method without a noise matrix, applying the same procedure at every step.
+Unsurprisingly, we found that the model failed to denoise the [MASK] tokens, resulting in sequences such as:
+[MASK] [MASK] educated ... educated [MASK] [MASK]
+2. BERT Denoising Matrix (0.15 Masking Ratio)
+Next, we implemented the denoising matrix using a BERT Denoising Matrix (0.15 Masking Ratio, $1-\frac{1}{T})$, which led to a strong bias toward a single repeated token:
+wwii wwii wwii wwii wwii wwii wwii wwii
+3. Time-Reversal Denoising (Tweedie-Leaping)
+Inspired by prior literature Lou et al.
+(2024), we then examined a Time-Reversal Denoising Schedule Tweedie $\tau$-leaping based on score entropy.
+However, in the paradetox SEDD experiments, we observed NA results under strict conditional generation settings.
+4. Word-Frequency-Based Denoising Schedule
+Subsequently, we applied a word-frequency-based denoising schedule He et al.
+(2022), but in the paradetox DiffusionBERT experiments, this approach encountered difficulties in constructing coherent sentences.
+5. Vocab-Wise Entropy Estimation
+Moving on, instead of relying on word frequency, we propose a vocab-wise entropy estimation technique.
+In particular, we construct the denoising matrix as shown in 2, leveraging entropy information to decide whether each word should be denoised or preserved.
+This approach assumes that all positions, including originally masked ones, can potentially be denoised.
+Although this approach did show some improvement, for instance producing:
+wwii reassure wwii bony wwii wwii wwii wwii
+Upon further analysis, we identified that the MLM was not effectively determining which positions to denoise, and well-generated tokens sometimes are converted [MASK], and then convert all [MASK] tokens into certain words in the final step, leading to token replication.
+6. Entropy-Based Estimation and Denoising
+Hence, we introduced an entropy-based estimation and denoising strategy.
+In this approach, we assume that once a mask is denoised, it remains fixed.
+Specifically, we select mask positions based on an entropy schedule, sample tokens for those positions, and once a token is sampled (i.e., denoised), we preserve it across subsequent diffusion steps.
+7. Entropy Selection Criteria
+We conducted three main experiments—uniform, reverse-order-EAGS, and EAGS—yielding perplexities of 182.976 with some portion of [MASK], 1193.229 with degenerated results, and 112.190 for paradetox dataset, respectively.
+These results indicate that noising from the most determinative token positions (mask with lowest entropy) is highly effective.
+Therefore, we adopt the Selection Criteria as EAGS.
+Figure 5: Entropy behavior tracking in generation/training process.
+Appendix D Entropy Flow
+In Figure 5, we illustrate the tendency of the sequential sum of entropy for various discrete generation processes.
+The changes of entropy during the generation process in Diffusion-EAGS, represented by the yellow line, show that our model effectively follows a gradual decrease in entropy, mirroring the inverse trend of the training process.
+This gradual change in entropy facilitates successful DDLM training, which results in superior text quality performance compared to other diffusion models, as demonstrated in Tables 28, and 9.
+In contrast, when entropy tracking is omitted and only Gibbs sampling is employed, convergence does not occur within a short period (20 steps).
+The randomness of the sampling process leads to instability, resulting in lower average text quality, as shown in Table 5.
+Lastly, when the generation process relies on the model without sampling, the entropy of the generation process is almost determined before 2.5 steps.
+This entropy behavior is similar to that observed in DiffusionBERT.
+EAGS Process:
+Input: Sequence Length $L$, Total Timestep $T$, Trained Model $M$, Mask Sequence Generator $G_{M}$, and Context $Y$
+for $t=T$ to 00 do
+       if $t=T$ then
+             $x^{T}\leftarrow G_{M}(L,Y)$ // Initialize a sequence of $L$
+       else
+             $f_{\theta}\leftarrow p_{\theta}(x^{t},Y)$ // Compute logits at timestep $t$
+[…]
+             $x^{t-1}\leftarrow p_{\theta}(x^{t},l^{*},Y)$ // Sample from the previous timestep
+       end if
+end for
+Algorithm 1 EAGS Algorithm
 ENS Process:
-Input: Context $Y$ and Dataset $D$
+Input: Context $Y$, Total Timestep $T$, and Dataset $D$
 […]
@@
        $x^{t+1}\leftarrow\text{Forward}(x_{0},\mathcal{H},t+1)$ // Forward at $t+1$
-       $L_{s}=-\sum_{i}q(x^{t}_{i}\mid x_{t+1})\log p_{\theta}(x^{t}_{i}\mid x_{t+1})$ // Loss calculation
+       $L_{s}=-\sum_{i}q(x^{t}_{i}\mid x_{t+1})\log p_{\theta}(x^{t}_{i}\mid x_{t+1})$ // Cross entropy loss calculation
 end for
-EAGS Process:
-Input: Sequence Length $L$, Total Timestep $T$, Trained Model $M$, Mask Sequence Generator $G_{M}$, and Context $Y$
-for $t=T$ to 00 do
-       if $t=T$ then
-             $x^{T}\leftarrow G_{M}(L,Y)$ // Initialize a sequence of $L$
-       else
-             $f_{\theta}\leftarrow M(x^{t},Y)$ // Compute logits at timestep $t$
+Algorithm 2 ENS Algorithm
+Appendix E EAGS & ENS algorithms
+Detailed algorithms of EAGS and ENS are in Algorithm 1 and 2.
+Appendix F Experiment
+F.1 Fine-Grained Conditional Generation
+Type Dataset RocStories Deontology Question Generation QQP DialogSum ALMA ParaDetox
+Open-ended Generation $\checkmark$ $\triangle$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
+Conditional Generation $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+– Context Provided ? $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+– Content Provided ? $\times$ $\triangle$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+– Format Provided ? - $\times$ $\times$ $\times$ $\triangle$ $\checkmark$ $\checkmark$
+Table 6: Each dataset has a different level of conditional constraints even if they are all conditional generation tasks.
+$\checkmark$ indicates full support, $\times$ indicates no support, and $\triangle$ indicates partial or limited support.
+In conditional generation tasks, the level of conditional constraint imposed by the dataset plays a critical role in shaping the generation process.
+As shown in Table 6, conditional constraints are diverse across datasets.
+In our task, we categorize these constraints into three levels: (1) the provision of context alone, requiring the continuity of the prefix; (2) the provision of specific content to be included in the target sequence, necessitating the inclusion of certain keywords; and (3) the provision of semantic content formatting, such as transforming toxic sentences into safer alternatives or converting text from the source language to a target language.
+In our study, we aim to develop a diffusion framework capable of being applied across a wide range of conditional generation tasks.
+F.2 Dataset Explanations
+Quasar-T QQP ParaDetox Deontology RocStories
+input output input output input output input output input output
+Max 63 244 104 98 35 35 24 31 76 57
+Mean 14.574 31.157 13.947 13.956 15.135 13.035 13.039 12.548 42.189 13.307
+Table 7: Dataset Statistics
+Open-ended Generation We employ the RocStories dataset (Mostafazadeh et al., 2016) for open ended generation with narrative understanding tasks.
+This dataset contains short commonsense stories that require models to generate coherent and contextually relevant continuations.
+Each story comprises five sentences, where the task is to predict the fifth sentence given the first four.
+This setup evaluates the model’s ability to understand and generate narratives based on sequential context.
+Deontology The objective of Deontology (Hendrycks et al., 2023) is to to evaluate the capability of models to make ethical judgments from a deontological perspective.
+The dataset contains scenarios focusing on interpersonal dynamics and everyday occurrences.
+Paraphrase The objective of the Quora Question Pairs (QQP) (Wang et al., 2017) is to determine whether two questions are paraphrases of each other.
+We process the QQP dataset by treating one question as a paraphrase of another, a method commonly employed to assess the effectiveness of diffusion models.
+QG The objective of Question Generation (QG) is to generate valid and fluent questions based on a given passage and a specified answer.
+We employ the Quasar-T dataset, introduced by Dhingra et al.
+(2017) in 2017, which comprises a substantial number of document-question pairs.
+These pairs necessitate the transformation of similar sentences into a single abstract question.
+DialogueSum In former experiments, it is hard to measure the performance with reference-based metrics as the limitation of traditional EM problems where conditional generation’s output space is wide.
+Therefore, to test our model’s capability, we experiment on dialogue summarization task Chen et al.
+(2021) which makes an emphasis on containing some keywords or necessary information in the generated sequences.
+We use the experimental dataset and evaluation metric proposed in DiffusionCG Xiang et al.
+(2024) with same experimental setting as former experiments.
 […]
-             $x^{t-1}\leftarrow\text{Sampling}(x^{t},l^{*},Y)$ // Sample from the previous timestep
-       end if
-end for
-Algorithm 1 ENS and EAGS Algorithm
-Our whole process is in Algorithm 1.
-Distinct from the prevailing methodologies in diffusion models as described by Ho et al.
-(2020a) and Austin et al.
-(2023), we do not employ the PLM parameterization strategy $\widetilde{p}_{\theta}(\widetilde{z}_{0}|z_{t},t)$, which preserves the original semantic embedding spaces during the training phase.
-We empirically find that PLM parameterization restricts the diversity of generated responses.
-Instead, to ensure the completeness of sentences, we implement Minimum Bayes Risk (MBR) decoding in the final generation step.
-We follow the traditional diffusion loss in Equation (Ho et al., 2020b), changing MSE with Cross Entrpy Loss.
-4 Experiments
-4.1 Tasks and Datasets
-Social Datasets The Paradetox dataset (Logacheva et al., 2022) focuses on removing profanities.
-The Deontology from ETHICS dataset (Hendrycks et al., 2023) evaluates ethical judgment based on scenarios.
-Paraphrase The QQP dataset (Wang et al., 2017) assesses paraphrase detection by treating one question as a paraphrase of another.
-Question Generation We use the Quasar-T dataset (Dhingra et al., 2017) to generate questions from passages and answers.
-Open-ended Generation The RocStories (Mostafazadeh et al., 2016) dataset is used to generate coherent story continuations from a given context.
-Detailed explanations are available in Appendix C and Dataset statistics are shown in Table 8.
-4.2 Baselines
-Since our objective is centered on generation tasks, we compare Diffusion-EAGS with four categories of baselines with the model similar to the size of RoBERTa-base (Liu et al., 2020): Auto-regressive Models (ARMs), CMLMs (Conditional Masked Language Models, Ghazvininejad et al.
-(2019a)), CDLMs, and DDLMs.
-For ARMs (Vaswani et al., 2023), we employ GPT-2 (Radford et al., 2019), which is renowned for its proficiency in generation tasks.
-For CMLMs, we utilize CMLM (Ghazvininejad et al., 2019a) and DisCo (Kasai et al., 2020), which are transformer-based NAR models.
-Inference is carried out using the Mask-Predict (Ghazvininejad et al., 2019a) and the Easy-First (Kasai et al., 2020) inference algorithm respectively.
-For CDLMs, our baseline includes DiffuSeq (Gong et al., 2023a) and DINOISER (Ye et al., 2024).
-DiffuSeq is a diffusion model specifically designed for sequence-to-sequence applications, whereas DINOISER adaptively determines the range of sampled noise scales during training.
-For DDLMs, we utilize DiffusionBERT (He et al., 2022) , a state-of-the-art model in the research series of DiffusionLM (Li et al., 2022), and SEDD (Lou et al., 2024), a state-of-the-art model of open-ended generation in diffusion language generation domain.
-For SEDD, we download the pretrained version and fine-tune on it.
-Experimental results on LLMs can be found in Appendix E.
-4.3 Metrics
-Quality metrics To measure the quality of the generated texts, we use Perplexity based-on GPT-2 Large and GPT-2 XL, SOME (Yoshimura et al., 2020), grammar metric based on the neural net, LLM-c (Lin and Chen, 2023) to measure the plausibility of the narratives, LLM-t (Koh et al., 2024) to measure toxicity, and MAUVE Pillutla et al.
-(2021), measuring the gap between text distributions despite divergence.
-Diversity Metrics Traditional diversity metrics Self-BLEU Zhu et al.
+Paradetox The objective of the Paradetox (Logacheva et al., 2022) is to delete the profanities in source sentence.
+It comprises of toxic and neutral utterances, curated from the Jigsaw, Reddit, and Twitter datasets.
+F.3 Experimental Details
+We employ roberta-base as MLM with learning rate 5e-4.
+The maximum lengths for QG, QQP, and Paradetox is set to 64, while for Deontology and DialogSum set to 48 and 292, respectively, based on data statistics.
+We test 20 conditions with 5 outputs in total 100, which is not used for training.
+The number of steps is configured to 5.
+We then perform a naive categorical sampling with a sample size of 20 and select final 5 samples based on PPL.
+We use 1 A100 GPU with the batch size as 256.
+Experimental details of LLMs are in Appendix J, machine translation in Appendix H.
+Quality metrics
+To measure the quality of the generated texts, we use Perplexity based-on GPT-2 Large and GPT-2 XL, SOME (Yoshimura et al., 2020), the grammar metric based on corpus, LLM-c (Lin and Chen, 2023) to measure the plausibility of the narratives, LLM-t (Koh et al., 2024a) to measure toxicity, and MAUVE Pillutla et al.
+(2021), measuring a reflectiveness of training dataset characteristics of generate outputs.
+MAUVE score of 1 indicates that the output perfectly matches the training dataset as a neural database.
+For Mean Opinion Score (MOS), we get 5 outputs from each condition.
+Subsequently, hired four integrated ph.d course work annotators in the university NLP research lab evaluate the generated text based on two criteria: (1) semantic reflectiveness of the condition, indicating how accurately the condition is represented in the text, and (2) sentence completeness, assessing overall grammatical and semantic coherence.
+Each criterion was rated on a scale from 0 to 1.
+Subsequently, these scores are normalized and averaged to obtain a final score ranging from 0 to 1.
+In our evaluation, Fleiss’ kappa Fleiss (1971) is exceeded 0.7 as assessing sentence quality is both intuitive and relatively non-controversial among the annotators.
+Diversity Metrics
+Traditional diversity metrics Self-BLEU Zhu et al.
 (2018) and distinct-n Li et al.
 (2015) are employed to evaluate the generated texts.
-We also adopt Vendi Score (VS) Friedman and Dieng (2023), an interpretable diversity metric, which quantifies the effective number of unique samples in a given set.
-Both the n-gram and embedding variations are utilized, where embedding VS interprets diversity of semantics.
-4.4 Experimental Details
-We employ roberta-base as encoder-based PLM.
-The learning rate is set at 5e-4, and a naive categorical sampling strategy is adopted.
-Adapted to data statistics, the maximum lengths for QG, QQP, and Paradetox is set to 64, while for Deontology set to 48.
-Furthermore, in accordance with the minimum construction length, the number of steps is configured to 5 and 20 size of MBR.
-We use 1 A100 GPU, and the batch size is set to 256.
+We also adopt Vendi Score (VS)-SimCSE Friedman and Dieng (2023), an interpretable diversity metric, which quantifies the effective number of unique samples in a given set.
+Both the n-gram and embedding variations are utilized, where embedding VS is semantic diversity.
+For the diversity MOS evaluation, we adopt the same methodology used for the quality MOS but apply two distinct criteria: (1) the condition’s semantic reflectiveness, and (2) sentence diversity, capturing both semantic and structural variety beyond mere word deletion or rearrangement.
+Ideal score of diversity MOS is 5 which means different five sequences for one condition, and lowest score is 1 which means all identical sequences.
 ParaDetox
+Text Quality Diversity
 Model Step PPL $\downarrow$ MAUVE $\uparrow$ SOME $\uparrow$ VS(ngram) $\uparrow$ VS(emb) $\uparrow$ self-bleu $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$
 GPT-2 1 389.1 0.503 0.717 3.925 2.640 0.429 0.312 0.748
+GPT-3.5 w/ 4-shot 1 104.375 0.175 0.888 3.098 1.915 0.652 0.390 0.835
+GPT-4 w/ 4-shot 1 78.979 0.125 0.879 3.214 1.906 0.592 0.412 0.841
 CMLM w/ Mask-Predict 10 669.9 0.0234 0.588 1.000 1.000 1.000 0.451 0.633
@@
 DiffuSeq 2000 $\geq 1k$ 0.683 0.703 2.059 1.465 0.841 0.410 0.820
+LD4LG 2000 579.9 0.556 0.762 1.914 1.425 0.845 0.419 0.829
 DINOISER 20 124.8 0.255 0.767 2.287 2.174 0.981 0.211 0.486
 SEDD 1024 $\geq 1k$ NA 0.664 4.746 4.063 0.119 0.451 0.846
-Diffusion-EAGS 5 69.5 0.773 0.796 4.755 3.659 0.126 0.475 0.834
+Diffusion-EAGS 5 109.3 0.811 0.760 4.417 3.311 0.256 0.407 0.810
 Deontology
@@
 QG
-Text Quality Diversity
 Model Step PPL $\downarrow$ MAUVE $\uparrow$ SOME $\uparrow$ VS(ngram) $\uparrow$ VS(emb) $\uparrow$ self-bleu $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$
@@
 Table 9: QG & QQP Generation
-5 Results
+Appendix G Detailed analysis of Results
+G.1 Fine-Grained Comparison
 As shown in Table 289, our model consistently exhibits exceptional performance in terms of text quality while simultaneously maintaining diversity when compared to baseline models.
+The standard deviation of PPL in Paradetox Experiment is 61 for our model.
+All other PPL’s standard deviation are similar to that of Paradetox.
 In Table 8 Paradetox, our model demonstrates superior performance across all evaluated metrics.
-Such phenomenon represents that our model based on encoder-based PLMs show robustness on diverse perturbations from daily dialogues.
+Such phenomenon represents that our model based on MLM shows robustness on diverse perturbations of daily dialogues.
 When PPL exceeds 600, the model is considered to have failed in generating natural sequences, and is thus represented in gray color.
-In the context of Deontology, our model exceeds the baseline PPL and MAUVE scores, whereas SOME score represent the sufficient quality of text with the highest diversity score of 4.755.
-In the paradetox experiment, the text quality produced by the CMLM and SEDD models was found to be low.
+Specifically, the text quality produced by the CMLM, which is standard BERT-generation, and SEDD, which is powerful model in open-ended generation, is found to be low.
 Consequently, these models were excluded from subsequent experiments.
+In Deontology, our model exceeds the baseline models’ PPL and MAUVE scores, whereas SOME score represent the sufficient quality of text with the highest diversity score.
 As illustrated in Table 9, Diffusion-EAGS generates the responses with the highest PPL score for QG, and highest MAUVE and PPL score for QQP.
-Additionally, in Table 4 where semantic consistency is crucial, our model significantly outperforms the original dataset’s PPL by 23 points by leveraging PLMs while maintaining a high diversity score of 4.837.
+While we adhere to the standard metrics commonly used in diffusion research and integrate as many additional metrics as possible, we also comprehensively explore our model’s capabilities across multiple dimensions.
+As the outputs of earlier generation tasks are too broad to be effectively evaluated using reference-based metrics, we provide generated examples in Appendix I and measure the preference of these outputs using a LLM-based metric in Appendix G.2.
+Additionally, to accommodate a scenario where reference-based evaluation is applicable, we have included a more extensive summarization task in Appendix G.2 and translation task in Appendix G.3.
+These results confirm that our method consistently produces outputs that adhere to the specified conditions.
 Diffusion-EAGS demonstrates the highest MAUVE score of 0.733 in Table 8-ParaDetox, and high level of text quality surpassing that of GPT-2 in Table 9 in text quality.
 ParaDetox is colloquial dataset including slang, numerous abbreviations, and various perturbations, so our model demonstrate robustness to such perturbations with 69.5 PPL.
-As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 23, and 4.
-These results underscore our model’s efficacy in generating diverse responses.
-Model PPL $\downarrow$ SOME $\uparrow$ LLM-c $\uparrow$ VS(ngram) $\uparrow$ self-bleu $\downarrow$
-Original Data 100.6 0.895 1
-GPT-2 88.5 0.856 0.88 4.722 0.124
-DiffusionBert 318.2 0.783 0.72 4.735 0.088
-SEDD 273.2 0.827 0.59 4.859 0.044
-Diffusion-EAGS 67.3 0.844 0.87 4.837 0.058
-Table 4: ROC Generation – Because ROC is open-ended generation task, we include SEDD in this experiment.
+As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 28, and 9.
 Notably, CDLMs demonstrate a noticeable deficiency in diversity.
@@
 Furthermore, our models require only a few steps, while resulting in higher quality and diversity.
-DiffusionBERT exhibits limitations in text quality, as shown in Table 234.
-On the other hand, our model achieves significantly higher scores across all quality metrics.
-This observation suggests that the naive application of BERT (Devlin et al., 2019) within the diffusion process fails to fully harness the capabilities of PLMs.
-SEDD, an open-ended generation model shows low performance under text generation in highly constrained tasks such as ParaDetox.
-Additionally, our model does not require more than 5 diffusion steps, as it can adaptively recover the absorbing state through an entropy-based generation approach.
-A detailed comparison is provided in Appendix D.
-6 Analysis
-6.1 Ablation Study
-To explore the effectiveness of our model’s components, we conduct ablation studies focusing on three key elements: Entropy Adaptation, stepwise Gibbs Sampling, and PLM in Table 5.
-The omission of EAGS initially leads to a substantial decline in performance and text quality, producing degenerated results similar to those of traditional CMLMs.
-Dataset PPL MAUVE SOME VS(ngram) VS(emb)
-Diffusion-EAGS Deont 55.1 0.412 0.835 4.898 4.009
-Roc 67.3 0.87 0.844 4.837 3.999
-w/o EAGS Deont 667.9 0.022 0.617 4.767 3.928
-Roc 1084.9 0.035 0.613 4.874 3.957
-w/o Gibbs Sampling Deont 1426.7 0.011 0.584 2.378 1.923
-Roc 1293.1 0.010 0.534 1.531 1.338
-w/o PLM Deont $\geq$2K 0.005 0.645 4.758 3.402
-Roc $\geq$2K 0.004 0.604 4.315 2.994
-Table 5: Ablation Study : Deont denotes Deontology
-We find that EAGS significantly contributes to a gradual entropy decrease, as shown in Appendix J.
-These observations highlight the critical role of selecting the sampling position based on given information in sequence generation.
-Additionally, excluding the use of the diffusion generation process, without the stepwise sampling, leads to a drastic reduction in overall performance, with PPL increasing by more than 1000 points and diversity scores dropping below 2.
-Consequently, our cMRF approach, integrated with the diffusion model, proves indispensable in aiding the PLM in effectively generating sequences.
-Furthermore, it is evident that without the incorporation of PLMs, the generation of natural sequences is unachievable.
-6.2 Case Study: Keyword Based Generation
-Our model operating within discrete space enables us to manipulate the output sequences using explicit instructions.
-To further explore this capability, we conduct the generation of sequences based on keywords positioned in the middle and at the end of masked sequences, which is challenging for AR models (Keskar et al., 2019).
-Initially, we provide the same contextual input while varying the keywords.
-In the masked states, we randomly select positions, replacing them with the specified keywords.
-The results in Table 7 demonstrate that the generated sequences seamlessly integrate the keywords with context-specific semantics.
-These findings show that our model can effectively integrate specific conditional information in an interpretable manner.
-Experimental results also indicates its substantial potential and suitability for diverse applications where direct controllability is crucial, such as in story generation.
-6.3 Bilinguality & Low Resource Settings
+G.2 Quality Recheck – LLM score & Dialogue Summarization
+Model ROUGE-1 ROUGE-2 MAUVE Ngram Emb Self-BLEU Distinct-1 Distinct-2
+Ours 0.409 0.174 0.536 4.114 2.591 0.252 0.253 0.632
+SEDD 0.179 0.032 0.999 4.216 2.576 0.211 0.200 0.609
+DINOISER 0.209 0.031 0.337 1.247 1.227 0.926 0.256 0.633
+Table 10: DialogueSum Experiment
+Model PPL MAUVE vs(ngram) VS(emb) sef-bleu distinct-1 distinct-2
+GENIE 134.1 0.296 2.527 1.800 0.702 0.454 0.825
+Table 11: Quantitative results for the GENIE model.
+Paradetox w/ LLM-t on application models
+LLM-t
+GPT-2 0.02
+GPT-3.5 0.074
+GPT-4 0.18
+DiffuSeq 0.03
+Diffusion-Bert 0.09
+DINOISER 0.1
+SEDD-small NA
+Diffusion-EAGS 0.01
+Table 12: ParaDetox Dataset Generation – LLM-t is the LLM-evaluation for measuring toxicity.
+Since our research primarily aims to enhance the model’s inherent capabilities, we set up baselines that revolve around (or are closely related to) noise scheduling.
+Nevertheless, some studies employ a hybrid framework integrating LLMs and diffusion models Lin et al.
+(2023); Xiang et al.
+(2024); hence, we conduct additional experiments to investigate this scenario.
+In addition, to evaluate the quality of the paradetox output and ours diffusion-EAGS still outperforms GENIE (Lin et al., 2023) in Table 11.
+We also use the LLM-t score Koh et al.
+(2024b) to measure whether models successfully detoxify the source condition, showing the quality of generated outputs from ours as shown in Table 12.
+Models Prefer Baseline Prefer Ours Tie
+diffuseq vs. ours 20% 65% 15%
+diffusionBERT vs. ours 20% 65% 15%
+dinoiser vs. ours 0% 90% 10%
+GPT-2 vs. ours 25% 65% 10%
+Table 13: Evaluation results comparing our model with various baselines.
+QG - LLM preference
+For Question Generation (QG), we employ the widely adopted GPT-as-a-Judge framework Zheng et al.
+(2023) to evaluate the quality of generations produced by our model and the baselines on the QG dataset.
+We adopt a pairwise evaluation setting, following the system and input prompts specified in Zheng et al.
+(2023) for the pairwise comparison.
+The factors specified to be evaluated are 1) coherency 2) grammatical correctness 3) semantic soundness 4) diversity and 5) being a more reasonable question to the input (condition) text.
+We employ the gpt-4 model.
+The result is as follows:
+Note that since within the prompt, the baseline model’s generations are specified prior to our model’s generation, there is a significant position bias working against our favor, as noted in Zheng et al.
+(2023).
+The results above indicate that despite such bias, our model’s generations are much more favored over the baselines’ generations.
+Dialoguesum Experiment
+Our model outperforms existing baselines in ROUGE, a reference-based metric as shown in Table 10.
+These findings indicate that, according to the automatic scores, our model sufficiently captures the source condition.
+Human Evaluation
+Below, we report the Mean Opinion Score (MOS) averages and standard deviations (std) in the following order: DiffusionBERT, LD4LG, GPT-2, Dinoiser, and our method.
+First, the average scores of semantic reflection are 0.98, 0.90, 0.94, 0.98, and 0.97, respectively, with standard deviations of 0.14, 0.30, 0.24, 0.14, and 0.16.
+Second, the average scores of sentence completeness are 0.78, 0.92, 0.72, 0.84, and 0.90, respectively, with standard deviations of 0.18, 0.14, 0.28, 0.15, and 0.15.
+Third, average scores of diversity are 2, 1, 2.65, 1, and 4.6, respectively, with standard deviations of 1.3, 0, 1.45, 0, and 0.7.
+GPT-3.5-turbo’s std is almost 0.
 Model SacreBLEU COMET XCOMET
@@
 Table 14: En-De Translation Results
-Context Jake was playing with his toys. He accidentally broke his favorite one.
-He cried a lot over it. His parents decided to replace it for him.
-Keyword not stop Jake just could not stop crying.
-Jake feel It made Jake feel So much better.
-would enjoy Jake said he would enjoy the new toy
-Context Neil was in Sofia, Bulgaria. He was enjoying a trip backpacking through Europe.
-… He thought the food and culture in Sofia were the best.
-Keyword Bulgaria! Things were looking great in Bulgaria!
-Context Karen wanted to go on a trip to France. She started doing research on the trip.
-She decided to book a week long trip. She left the next day for her trip.
-Keyword her trip She spent almost a week there during her trip.
-Table 7: Keyword-based Generation
-[…]
-Although our baseline diffusion models demonstrated significant performance improvements over CMLMs, CMLMs were primarily developed to enhance translation performance.
-Therefore, we conducted additional experiments to further validate our methodology such as Mask-and-Predict and Easy-First, diffusion models such as Diffuseq-v2 (Gong et al., 2023b) and SEDD, traditional translation models such as mBART-50 (Tang et al., 2020) and NLLB.
+G.3 Machine Translation : Bilinguality & Low Resource Settings
+Labeled datasets used in conditional generation tasks are typically limited in size and sometimes multilingual.
+To further assess our model’s performance in conditional generation, particularly in terms of language extension and resource scarcity, we conduct additional experiments on a translation task.
+We conduct additional experiments on CMLMs such as Mask-and-Predict and Easy-First, diffusion models such as Diffuseq-v2 (Gong et al., 2023b) and SEDD, traditional translation models such as mBART-50 (Tang et al., 2020) and NLLB.
 For evaluation metrics, we utilize sacreBLEU (Post, 2018) and neural-net scores such as COMET (Rei et al., 2020) and XCOMET (Guerreiro et al., 2023).
-Table 6 shows that predicting the target sequence without leveraging a multilingual PLM proves to be challenging.
+More details are provided in Appendix L.
+Table 14 shows that predicting the target sequence without leveraging a multilingual model proves to be challenging.
 All diffusion baseline models struggle to produce correct outputs.
-For example, the pretrained SEDD model fails to effectively leverage conditional information, even after finetuning on the training datasets, consistent with the limitation observed in Section D.3.
 Similar challenges arise in NAR transformer baselines.
 Despite constructing the vocabulary using the pretrained mBART-50 model (DisCo-m), the underlying issues remain.
-On the other hand, our proposed model, by incorporating a PLM, demonstrates promising results.
-In addition, these results show that our methodology can also be used with encoder-decoder models.
-Interestingly, the output of the pretrained NLLB model (NLLB-naive-600M, not finetuned) reveal that neural network-based metrics are susceptible to the interference problem, specifically translated by other languages, even though we provide the language specific token.
-While such issues result in lower BLEU scores, COMET and XCOMET often interpret them as semantically coherent, indicating a potential direction for future work to improve translation evaluation metrics.
-Despite these phenomena, a performance gap between translation models and DDLM remains.
-This suggests that future research should address the semantic capabilities of diffusion models to help bridge this gap.
-Details are in Appendix I.
-6.4 Diversity Saturation
+On the other hand, our proposed model demonstrates promising results.
+G.4 Diversity Analysis
+Limitation of Diversity on Traditional DDLMs
+GPT-2 Given Source : holy shit , they blew up a real artifact this time ?
+from GPT2s output ended by end token with default temperature sampling from huggingface : - Oh my god,they blew a really important artifact in this year?????..safe: Oh - Oh my god, they destroyed a really important artifact in this year?...safe: Oh God,they - they blow up something thistime?.??.safe: Oh my god, they destroyed a really valuable artifact - Oh my god, they destroyed a really important artifact in this year??.?!??.?!?
+- They blew an artifact that time?
+Theyre still in the artifact?
+This time, theyre in trouble.
+This
+Tranditional Diffusion Models traditional diffusion models output from Dinoiser, LD4LG : - they blew up a real artifact this time?
+- they blew up a artifact this time?
+- they blew up a real artifact?
+- they blew up a real artifact this time?
+- they blew up a real artifact this time?
+Ours : Diffusion-EAGS from ours : - aww, it is really a real artifact this time ?
+- it seems like they destroyed an artifact in this time as well - they have blown up a large artifact - they have blown up it in a museum, this time - they also destroyed artifacts at the same time
+We summarize the generation trends of the models presented in Table above.
+We observe that when a fine-tuned GPT-2 is tasked with strongly constrained conditional generation, it struggles to properly terminate sentences with an <eos> token.
+In particular, it shows limitations when handling semantic leaps or clearly delineated structural constraints, leading to suboptimal conditional generations.
+Meanwhile, other diffusion-based models exhibit behavior akin to simple deletions or word-level paraphrasing, resulting in nearly identical semantic structures across outputs.
+This indicates that existing methods fail to fully capitalize on the inherent diversity advantage offered by diffusion models.
+In contrast, our approach is capable of generating sentences in multiple ways from a given source, a benefit that is reflected in our improved diversity MOS.
+Diversity Saturation on LLMs
 Figure 6: Diversity graph with increasing generation numbers in ’Deontology’ dataset
 Inspired by the observation that Diffusion-EAGS consistently excel in terms of diversity across all results, we delve further into the diversity capabilities of our model.
-We assess the diversity performance in conditional generation compared to LLMs.
+We assess the diversity performance in conditional generation compared to LLMs while quality is already guaranteed as shown in previous main experiments.
 We measure the VS for 5 to 100 generations under a single condition.
@@
 The experiment is conducted on the ‘deontology’ dataset which allows high output diversity in its settings.
-Details of using LLMs are provided in Appendix G.
+Details of using LLMs are provided in Appendix J.
 Figure 6 demonstrates that the diversity saturation graph for Diffusion-EAGS has a relatively steep slope, while GPT models saturate at lower values.
@@
 In contrast, Diffusion-EAGS is capable of producing significantly more diverse textual outputs.
-7 Related Works
-Efforts to integrate generative flow models into sequence generation exploit the distribution shift from a source language to a target language through a series of invertible linear transformations (Ma et al., 2019; Zhang et al., 2024).
-However, as DDPM (Ho et al., 2020a) demonstrate the effectiveness of generating images, diffusion models have been a major topic of interest within the field of generative flow models (Song et al., 2021a, b).
-To apply such diffusion methodologies to NLP, there are two main streams - continuous diffusion models and discrete diffusion models.
-Continuous diffusion models Diffusion-LM (Li et al., 2022) propose a method of mapping tokenized sequences to embedding dimensions guided by a pretrained classifier.
-DiffuSeq-v1, v2 (Gong et al., 2023a, b) apply partial noising techniques.
-The core of these diffusion methodologies lies in the addition and restoration of random noise to facilitate generation.
-However, authors of Cold Diffusion (Bansal et al., 2022) argue that such operations do not necessarily have to be governed by stochastic randomness.
-Discrete diffusion models Building on this rationale, D3PM (Austin et al., 2023) propose the discrete restoration-generation approach and DiffusionBERT He et al.
-(2022) adopt PLMs to DDLM.
-SEDDLou et al.
-(2024) propose score entropy inspired by MLM loss.
-Recent works by Shi et al.
-(2024) and Sahoo et al.
-(2024) extend this idea to propose a simplified view of the discrete framework, and obtain better empirical results.
-Zheng et al.
-(2024) further close the gap between discrete diffusion models and masked models, and also, by correcting the numerical precision error during the sampling process in SEDD-based models, reveal that the true generation perplexity of DDLMs lags behind that of AR models.
-These research make an improvement on the open ended generation task.
-Furthermore, Venkatraman et al.
-(2024) showcase that SEDD priors can be used to sample from a posterior distribution, namely text infilling, and Nie et al.
-(2024) demonstrate that DDLMs are scalable, and have advantages over traditional AR models.
-8 Conclusions
-In this work, we introduce Diffusion-EAGS, integrating PLMs with diffusion models for conditional generation.
-By leveraging step-wise diffusion models and the property of MLM, we demonstrate that MLM-based PLMs can serve as denoising function with Entropy-based Adaptive Gibbs Sampling.
-We also propose entropy-based noise scheduling during training for adaptive sampling.
-Experimental results demonstrate that Diffusion-EAGS outperforms existing baselines, yielding improved text generation quality, enhanced diversity, broad applicability, and more precise token-level control.
-Limitations
-While Diffusion-EAGS demonstrates significant improvements in conditional generation tasks, there are several limitations.
-Firstly, as our method is currently focused on text generation tasks, its applicability to text classification tasks, such as Named Entity Recognition and Part-of-Speech Tagging, remains unexplored.
-Future research could explore extending this method to other NLP tasks.
-Secondly, future works could investigate the potential in reducing computational costs through more efficient entropy calculation techniques.
-Lastly, although our current efforts concentrate on developing and validating our framework using encoder-only and encoder-decoder architectures, the potential integration of autoregressive (AR) models remains unexplored.
-However, since the denoising function—with its entropy-based, sequential ordering—aligns well with autoregressive decoding strategies, we posit that AR models can also serve as effective initializations within our proposed framework.
-Ethical Statements
-A notable advantage of Diffusion-EAGS is its efficient training process, which optimizes computational resources with high performance.
-This efficiency not only reduces the environmental impact of training PLMs but also makes advanced NLP technologies more accessible.
-Additionally, we have focused on output control and interpretability through the use of a discrete diffusion model.
-Consequently, our methodology contributes to the effective control of generated outputs in the future.
-Diffusion-EAGS’s strength in tasks involving conditional generation proves beneficial in the field of text generation following social guidelines, where outputs need to strongly depend on the conditions.
-Specifically, its high diversity performance enables active utilization in dynamically adapting text generations to meet specific social criteria.
-Appendix A Measuring Potential Function $\phi(X)$ in MLM
-In this section, we provide additional experimental details and results to support the observation that open-ended Masked Language Models (MLMs) exhibit increased potentials for the same sequence under different dataset constraints.
-A.1 Experimental Setup
-  • Model We use the pre-trained BERT large model (bert-large-cased) as the base model for all experiments.
-Additionally, we incorporate RocStories-conditioned guidance with the pre-trained model and use a fine-tuned BERT model on the RocStories dataset to evaluate the impact of dataset-specific constraints.
-  • Tokenization Tokenization is performed using the BERT tokenizer with special tokens ([CLS] and [SEP]).
-  • Potential Calculation The potential function $\phi(X)$ is computed as described in Equation 1 and  3 , where the log-potentials are obtained for each token using masked token logits.
-  • Datasets
-    • RocStories: Structured narratives from the RocStories dataset.
-A.2 Results of Experiment and Implications for Conditional Generation
-Using the BERT-large-cased model, the average log potential value for the standard MLM was 156.6150, while incorporating RocStories guidance increased this value to 175.5332, highlighting the impact of dataset-specific constraints.
-Additionally, fine-tuning the same model on RocStories resulted in an average potential function value of 3.7551 (on an exponential scale), demonstrating substantial variation introduced by conditional generation settings.
-The results demonstrate the significant influence of dataset structure on the potential function in MLMs.
-Specifically, structured datasets like RocStories enforce stronger narrative constraints, leading to higher potentials and greater coherence in sequence generation.
-This supports the main text’s argument that conditionally constrained spaces enhance the consistency and predictability of MLM outputs.
-Appendix B BERT as a MRF
-Let $X=(x_{1},\ldots,x_{L})$ be a sequence of random variables $x_{l}$, each taking a value from a vocabulary $V=\{v_{1},\ldots,v_{M}\}$.
-These variables form a fully-connected graph, indicating mutual dependence.
-To define the MRF, we focus on the full-graph clique potential:
-$\phi(X)=\exp\left(\sum_{l=1}^{L}\log\phi_{l}(X)\right)$ (3)
-where each log-potential $\phi_{l}(X)$ is:
-$\log\phi_{l}(X)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\}),&\text{if }\text{[MASK]}\notin X_{l-1:l+1}\\ \\ 0,&\text{otherwise}\end{cases}$ (4)
-Here, $f_{\theta}(X\backslash\{x_{l}\})$ is a function that maps the sequence $X\backslash\{x_{l}\}$ to a real-valued vector in $\mathbb{R}^{M}$, and $1h(x_{l})$ is a one-hot vector with $x_{l}$ set to 1.
-from this log potential, we can find probability of sequence X:
-$p_{\theta}(X)=\frac{1}{Z(\theta)}\prod_{l=1}^{L}\phi_{l}(X)$
-with the normalization constant $Z(\theta)$ defined as:
-$Z(\theta)=\sum_{X^{\prime}}\prod_{l=1}^{L}\phi_{l}(X^{\prime})$
-For a fixed $X\backslash\{x_{l}\}$, the conditional probability of $x_{l}$ is:
-$p(x_{l}|X\backslash\{x_{l}\})=\frac{1}{Z(X\backslash\{x_{l}\})}\exp(1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\}))$
-where $Z(X\backslash\{x_{l}\})$ is:
-$Z(X\backslash\{x_{l}\})=\sum_{m=1}^{M}\exp(1h(m)^{T}f_{\theta}(X\backslash\{x_{l}\}))$
-BERT uses pseudo log-likelihood learning to handle the intractability of $Z(\theta)$:
-$\text{PLL}(\theta;D)=\frac{1}{|D|}\sum_{X\in D}\sum_{l=1}^{|X|}\log p(x_{l}|X\backslash\{x_{l}\})$
-where $D$ is the training dataset.
-Appendix C Dataset Explanations
-Quasar-T QQP ParaDetox Deontology RocStories
-input output input output input output input output input output
-Max 63 244 104 98 35 35 24 31 76 57
-Mean 14.574 31.157 13.947 13.956 15.135 13.035 13.039 12.548 42.189 13.307
-Table 8: Dataset Statistics
-Paradetox The objective of the Paradetox (Logacheva et al., 2022) is to delete the profanities in source sentence.
-It comprises of toxic and neutral utterances, curated from the Jigsaw, Reddit, and Twitter datasets.
-Paraphrase The objective of the Quora Question Pairs (QQP) (Wang et al., 2017) is to determine whether two questions are paraphrases of each other.
-We process the QQP dataset by treating one question as a paraphrase of another, a method commonly employed to assess the effectiveness of diffusion models.
-QG The objective of Question Generation (QG) is to generate valid and fluent questions based on a given passage and a specified answer.
-We employ the Quasar-T dataset, introduced by Dhingra et al.
-(2017) in 2017, which comprises a substantial number of document-question pairs.
-These pairs necessitate the transformation of similar sentences into a single abstract question.
-Deontology The objective of Deontology (Hendrycks et al., 2023) is to to evaluate the capability of models to make ethical judgments from a deontological perspective.
-The dataset contains scenarios focusing on interpersonal dynamics and everyday occurrences.
-Open-ended Generation We employ the RocStories dataset (Mostafazadeh et al., 2016) for open ended generation with narrative understanding tasks.
-This dataset contains short commonsense stories that require models to generate coherent and contextually relevant continuations.
-Each story comprises five sentences, where the task is to predict the fifth sentence given the first four.
-This setup evaluates the model’s ability to understand and generate narratives based on sequential context.
-Appendix D Detailed analysis of Results
-D.1 Ours vs AR model
-Notably, our model effectively reflects the pattern and style of the dataset, attributed to the capability of encoder-based PLMs.
-Specifically, as reported in Table 2-ParaDetox, our model excels at capturing semantic information with the highest MAUVE score of 0.733, despite ParaDetox being a challenging colloquial dataset including slang, numerous abbreviations, and various perturbations.
-In addition, our model demonstrate robustness to such perturbations with 69.5 PPL, while GPT-2 shows a lower performance of 389.1 PPL.
-Diffusion-EAGS demonstrates high level of text quality surpassing that of GPT-2 in Table 3 in text quality.
-As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 23 and 4.
-These results underscore our model’s efficacy in generating diverse responses.
-Specifically, in the context of story generation as shown in Table 4, our model not only demonstrates effectiveness in enhancing both the quality and diversity of the text but also maintains coherence with the contextual storyline.
-This is further supported by the high LLM-c scores, which confirm the plausibility of the narratives produced by our model.
-D.2 Ours vs CDLMs
-Notably, CDLMs demonstrate a noticeable deficiency in diversity.
-In contrast, our model excels at producing significantly more diverse sequences.
-This enhanced performance can be attributed to the integration of PLM into our framework, which evidently enriches the diversity of the generated outputs.
-In addition, our model generally achieves higher text quality scores than CDLMs across all experiments.
-In case, our model shows a lower PPL than DINOISER, but DINOISER’s MAUVE score approaches zero, indicating that their generated responses are not based on the dataset.
-Diffusion-EAGS not only surpasses Diffuseq and DINOISER in PPL and SOME but also achieves a MAUVE score improvement of over 0.4, thus indicating that the generated responses accurately reflect the dataset while preserving grammatical integrity.
-Furthermore, our models require only a few steps, resulting in a 40% reduction of costs with higher quality and diversity.
-D.3 Ours vs DDLMs
-Contrary to CDLMs, DiffusionBERT exhibits limitations in text quality, as shown in Table 234.
-On the other hand, our model achieves significantly higher scores across all quality metrics.
-This observation suggests that the naive application of BERT (Devlin et al., 2019) within the diffusion process fails to fully harness the capabilities of PLMs.
-SEDD, an open-ended generation model shows low performance under text generation in highly constrained tasks such as ParaDetox.
-Even in open-ended generation settings of Table 4, Diffusion-EAGS shows higher text quality and semantic scores.
-Additionally, our model does not require more than 5 diffusion steps, as it can adaptively recover the absorbing state through an entropy-based generation approach.
-Above results further exemplify the high performance and the time efficiency of our model.
-Appendix E Experimental Outputs
-E.1 LLM Evaluation
-Table 9: ParaDetox LLM Evalution (LLM-t)
+G.5 Keyword Generation Results
+As shown in Table 4, our model successfully generate the coherent sequences based on givel context and keywords.
+G.6 Different PLM : BERT & T5
+While our primary approach integrates BERT into the diffusion framework via a theoretical cMRF interpretation, we also experiment with other PLMs such as BERT and T5, because the main experiment involved BART (LD4LG) and GPT-2.
+Specifically, RoBERTa and BERT exhibit similar trends, whereas T5 shows behavior comparable to a fine-tuned GPT-2 in Table 15.
+We conjecture that T5 is already trained with an autoregressive strategy in its decoder whose generation process is largely influenced by its initial decoder tokens from an entropy perspective Wang and Zhou (2024), resulting in relatively low diversity.
+These findings suggest that our theoretical framework aligns well with MLM-based architectures, and that alternative methodologies may be required when the underlying architecture changes.
+Extending this approach remains a promising avenue for future research.
+Model PPL MAUVE VS(ngram) vs(emb) self-bleu distinct-1 distinct-2
+RoBERTa 109.3 0.811 4.417 3.311 0.256 0.407 0.810
+BERT 69.5 0.773 4.755 3.659 0.126 0.475 0.834
+T5 408.1 0.378 2.256 1.666 0.750 0.415 0.773
+Table 15: Performance comparison of T5, BERT, and RoBERTa.
+Appendix H Experimental Outputs
+H.1 LLM Evaluation
+The LLM evaluation prompt for ParaDetox is provided in Table 16, and the LLM evaluation prompt for RocStories is given in Table 17.
 ParaDetox Prompt
@@
 Take a deep breath and think step-by-step.
-Table 10: RocStories LLM Evalution (LLM-c)
+Table 16: ParaDetox LLM Evalution (LLM-t)
 RocStories Prompt
@@
 Take a deep breath and think step-by-step.
-E.2 Tables
-We have appended the performance metrics of LLMs for reference.
-LLMs exhibit strong performance in tasks like QG, QQP, and Deontology, where datasets are more neat and contain fewer noises and abbreviations.
-Conversely, performance is lower on the Paradetox dataset, which includes a wide range of casual phrases and diverse perturbations.
-Provided that our model utilizes encoder-based PLMs, we observe that it demonstrates robustness to such perturbations compared to autoregressive models.
-ParaDetox
-PPL MAUVE SOME LLM-t VS(ngram) VS(emb) self-bleu distinct-1 distinct-2
-GPT-2 389.147 0.503 0.717 0.02 3.925 2.640 0.429 0.312 0.748
-GPT-3.5 104.375 0.175 0.888 0.074 3.098 1.915 0.652 0.390 0.835
-GPT-4 78.979 0.125 0.879 0.18 3.214 1.906 0.592 0.412 0.841
-DiffuSeq $\geq 1k$ 0.683 0.703 0.03 2.059 1.465 0.841 0.410 0.820
-Diffusion-Bert 775.928 0.737 0.716 0.09 3.101 2.058 0.599 0.424 0.826
-DINOISER 124.797 0.255 0.767 0.1 2.287 2.174 0.981 0.211 0.486
-SEDD-small $\geq 1k$ NA 0.664 NA 4.746 4.063 0.119 0.451 0.846
-Diffusion-EAGS 69.5 0.773 0.796 0.01 4.755 3.659 0.126 0.475 0.834
-Table 11: ParaDetox Dataset Generation : We use LLM-evaluation approach for measuring toxicity, denoted as LLM-t
-Deontology
-PPL MAUVE SOME VS(ngram) VS(emb) self-bleu distinct-1 distinct-2
-GPT-2 91.962 0.131 0.860 3.665 3.126 0.425 0.474 0.874
-GPT-3.5 52.401 0.393 0.904 4.632 3.650 0.186 0.434 0.855
-GPT-4 72.329 0.465 0.921 4.530 3.286 0.191 0.425 0.865
-Table 12: Deontology Dataset Generation
-QQP
-PPL MAUVE SOME VS(ngram) VS(emb) self-bleu distinct-1 distinct-2
-GPT-2 66.270 0.112 0.754 3.886 2.566 0.423 0.344 0.787
-GPT-3.5 55.275 0.708 0.874 2.781 1.603 0.691 0.365 0.814
-GPT-4 66.121 0.814 0.877 2.981 1.651 0.673 0.423 0.866
-Table 13: QQP Dataset Generation
-QG
-PPL SOME MAUVE VS(ngram) VS(emb) self-bleu distinct-1 distinct-2
-GPT-2 124.831 0.759 0.175 4.564 3.130 0.176 0.210 0.629
-GPT-3.5 57.539 0.808 0.0184 3.174 2.090 0.612 0.282 0.739
-GPT-4 75.435 0.822 0.0171 3.819 2.069 0.444 0.311 0.773
-Table 14: QG Dataset Generation
-Appendix F Well-Generated Output Examples
-F.1 ParaDetox
-Table 15: ParaDetox generation examples
+Table 17: RocStories LLM Evalution (LLM-c)
+Appendix I Well-Generated Output Examples
+Generated examples of Paradetox are provided in Table 18, Deontology in Table 19, QQP in Table 20, QG in Table 21, and RocStories in Table 22.
 ParaDetox Generation Output Examples of Diffusion-EAGS
@@
 Output $X_{2}$ “of course we cannot justify it.”
-Table 16: As the dataset is designed to eliminate problematic words while preserving semantics, our model adheres to these guidelines.
-F.2 Deontology
-Table 17: Deontology generation examples
+Table 18: ParaDetox generation examples
 Deontology Generation Output Examples of Diffusion-EAGS
@@
 Output $X_{2}$ “So I have to rent it to others.”
-F.3 QQP
-Table 18: QQP generation examples
+Table 19: Deontology generation examples
 QQP Generation Output Examples of Diffusion-EAGS
@@
 Output $X_{2}$ “How can you store energy from lightning?”
-F.4 QG
-Table 19: QG generation examples
+Table 20: QQP generation examples
 QG Generation Output Examples of Diffusion-EAGS
@@
 Output $X_{2}$ “What is the smallest animal without a pouch.”
-F.5 RocStories
-Table 20: RocStories generation examples
+Table 21: QG generation examples
 RocStories Generation Output Examples of Diffusion-EAGS
@@
 Output $X_{2}$ “So he got a brand new one after all.”
-Appendix G Details on Text Augmentation Using GPT models
-G.1 GPT-3.5turbo ~ GPT-4-Omni
+Table 22: RocStories generation examples
+Appendix J Details on Text Augmentation Using GPT models
+J.1 GPT-3.5turbo ~ GPT-4-Omni
 We prompt the GPT models to carry out dataset augmentation.
@@
 TARGET:
-Appendix H Details on CDLMs
-H.1 Experimental Details
+Appendix K Details on CDLMs
+K.1 Experimental Details
 For the case of Diffuseq and Dinoiser Ye et al.
@@
 Note that unlike other benchmarks, we experimented with Diffuseq-v2 (Gong et al., 2023b) in translation task for a broader comparison with existing baselines.
-H.2 Results Interpretations
+K.2 Results Interpretations
 Examining the results of Diffuseq, it is evident that the grammar score is comparatively lower than that of other models.
@@
 This finding, coupled with our additional experiments concerning the beam size during Dinoiser generation, suggests that Dinoiser’s performance predominantly relies on memorization.
-Appendix I Details on Translation Results
-I.1 Comparison Between Easy-First and Our Proposed Method
+Appendix L Details on Translation Results
+L.1 Datasets & Observations
+[…]
+Interestingly, the output of the pre-trained NLLB model (NLLB-naive-600M, not finetuned) reveal that neural network-based metrics are susceptible to the interference problem, specifically translated by other languages, even though we provide the language specific token.
+While such issues result in lower BLEU scores, COMET and XCOMET often interpret them as semantically coherent, indicating a potential direction for future work to improve translation evaluation metrics.
+Despite these phenomena, a performance gap between translation models and DDLM remains.
+This suggests that future research should address the semantic capabilities of diffusion models to help bridge this gap.
+L.2 Comparison Between Easy-First and Our Proposed Method
 Discrete diffusion can be said to inherit ideas from NAR inference algorithm Mask-PredictGhazvininejad et al.
@@
 (2019b) on the original DisCo architecture implementation as baselines on translations tasks in Table 14 to further elucidate the difference through empirical results.
-I.2 Experimental Details
+L.3 Experimental Details
 NAR Transformer & CMLM We utilized the official repository to produce obtain the results, with the default architecture, optimization, and inference configurations.
 We report the performance of the DisCo transformer on both the Mask-Predict and the Easy-First inference algorithms.
-Diffuseq-v2 For Diffuseq-v2, we employ the vocabs of mBERT and choose 128 as max length for ende translation. Other settings are identical as in Appendix H.1.
+Diffuseq-v2 For Diffuseq-v2, we employ the vocabs of mBERT and choose 128 as max length for ende translation. Other settings are identical as in Appendix K.1.
 SEDD The SEDDLou et al.
@@
 DiffusionEAGS For our model, we adopt the denosing strategy as top1 sampling and 1 size of MBR as typical translation task focuses on BLEU and COMET rather than diversity score.
-I.3 Experimental Results
-I.3.1 NAR Transformer, DisCo
+L.4 Experimental Results
+L.4.1 NAR Transformer, DisCo
 The results indicate that the DisCo transformer performs poorly on low-resource translation tasks, where the size of the dataset is small.
@@
 The performance benefits slightly from this change, but still lags behind those of other models.
-I.3.2 Diffuseq-v2
+L.4.2 Diffuseq-v2
 It is notable that existing diffusion language models perform poorly on translation tasks. In this section, we introduce some observations that might aid our understanding of such behaviors.
@@
 This analysis suggests that a simple rounding approach from continuous to discrete space may be insufficient for machine translation, at least in low-resource settings.
-Appendix J Entropy Flow
-Figure 4: Entropy behavior tracking in generation/training process.
-In Figure 4, we illustrate the tendency of the sequential sum of entropy for various discrete generation processes.
-The changes of entropy during the generation process in Diffusion-EAGS, represented by the yellow line, show that our model effectively follows a gradual decrease in entropy, mirroring the inverse trend of the training process.
-This gradual change in entropy facilitates successful DDLM training, which results in superior text quality performance compared to other diffusion models, as demonstrated in Tables 3, 2, and 4.
-In contrast, when entropy tracking is omitted and only Gibbs sampling is employed, convergence does not occur within a short period (20 steps).
-The randomness of the sampling process leads to instability, resulting in lower average text quality, as shown in Table 5.
-Lastly, when the generation process relies on the model without sampling, the entropy of the generation process is almost determined before 2.5 steps.
-This entropy behavior is similar to that observed in DiffusionBERT.

### v5

@@
 where $p_{\theta}(x^{\prime}_{i};X^{(t)})$ is the softmax probability of token $x^{\prime}_{i}$ at position $i$ in sequence $X^{(t)}$, and $t$ denotes the diffusion timestep.
+We then select high-entropy positions for updating:
 $M_{t}=\{i\mid H_{i}(X^{(t)})\geq\tau_{t}\}$ (5)
@@
 The maximum lengths of Paradetox and RocStories is set to 64, based on data statistics, and other details are in Appendix F.
+We test 20 conditions with 5 outputs in total 100, which is not used for training.
 The number of steps of our model is configured to 5 with a naive categorical sampling with a sample size of 20 and select final 5 samples based on Perplexity score.
@@
 Note that our primary goal is to investigate the architecture’s capabilities; any baseline approach in the direction of scalability or bypassing the architecture’s limitations goes beyond our research scope.
-For ARMs (Vaswani et al., 2023), we employ GPT-2 (Radford et al., 2019) and GPT-3.5-turbo with four-shot prompt.
-More experimental details of GPT-3.5 can be found in Appendix J, and H.
+For ARMs (Vaswani et al., 2023), we employ GPT-2 (Radford et al., 2019) and gpt-3.5-turbo***https://platform.openai.com/docs/models/gpt-3-5 with four-shot prompt.
+More experimental details of GPT-3.5 can be found in Appendix J.
 For CMLMs, we utilize CMLM-Mask-Predict (Ghazvininejad et al., 2019a) and DisCo-Easy-First (Kasai et al., 2020), which are transformer-based NAR models.
@@
 For SEDD, we download the pre-trained version and fine-tune it.
+More details are in Appendix F.3.
 5.3 Metrics
@@
 (2021) to assess style consistency between the training data and generated outputs, SOME (Yoshimura et al., 2020) to score the grammar, Mean Opinion Score (MOS) from human evaluations to gauge text quality, and LLM score such as LLM-c (Lin and Chen, 2023) to measure the plausibility of the narratives as a sub-metric.
-Further details on LLM scores and other evaluation settings are provided in Appendix F.3 and H.1.
 Diversity Metrics
 Following our quality assessment, we evaluate diversity through three different measures: an automatic frequency-based metric n-gram Vendi Score(VS n-gram) Friedman and Dieng (2023), a neural network–based semantic metric SimCSE Vendi Score (VS emb), and a human evaluation score MOS.
-More details related to MOS and other scores are in Appendix F.3.
-6 Results
+The detailed descriptions of metrics are provided in Appendix F.3 and H.1.
 Model Text Quality
@@
 Table 2: Results on the open-ended RocStories (ROC) dataset. We report perplexity (PPL) for fluency, SOME and LLM-c for text quality, and both VS(ngram) and self-BLEU for diversity.
+6 Results
 In Tables 1, 2, and 3, our model consistently demonstrates strong text quality and diversity compared to various baselines across a wide range of conditional generation tasks.
@@
 Overall, Diffusion-EAGS consistently demonstrates strong performance across diverse conditional generation tasks, combining low perplexity and high human evaluation scores with the ability to generate richly varied text.
+Detailed results are in Appendix G and examples are in Appendix I.
 7 Analysis
@@
 This setup evaluates the model’s ability to understand and generate narratives based on sequential context.
-Deontology The objective of Deontology (Hendrycks et al., 2023) is to to evaluate the capability of models to make ethical judgments from a deontological perspective.
+Deontology The objective of Deontology (Hendrycks et al., 2023) is to evaluate the capability of models to make ethical judgments from a deontological perspective.
 The dataset contains scenarios focusing on interpersonal dynamics and everyday occurrences.
@@
 We use 1 A100 GPU with the batch size as 256.
-Experimental details of LLMs are in Appendix J, machine translation in Appendix H.
+For the case of ARMs, CMLMs, CDLMs, and DDLMs, we follow the official repositories to reproduce the results.
+Results are sampled multiple times with different seeds to evaluate the diversity.
+For hyperparameters, we follow the original repositories if the parameter is provided, except for modifying the number of samples as 5 and max_length parameters according to data statistics.
+Note that unlike other benchmarks, we experiment with Diffuseq-v2 (Gong et al., 2023b) in translation task for a broader comparison with existing baselines.
+Moreover, experimental details of LLMs are in Appendix J, machine translation in Appendix H.
 Quality metrics
@@
 For Mean Opinion Score (MOS), we get 5 outputs from each condition.
+For a fair MOS comparison, if GPT-3.5-turbo refuses to provide an answer or if sentence completeness is compromised by condition consisting of “rtttt,” or extreme elliptical expressions, we exclude such relevant condition from our evaluation target.
 Subsequently, hired four integrated ph.d course work annotators in the university NLP research lab evaluate the generated text based on two criteria: (1) semantic reflectiveness of the condition, indicating how accurately the condition is represented in the text, and (2) sentence completeness, assessing overall grammatical and semantic coherence.
@@
 Table 9: QG & QQP Generation
+Model ROUGE-1 ROUGE-2 MAUVE Ngram Emb Self-BLEU Distinct-1 Distinct-2
+Ours 0.409 0.174 0.536 4.114 2.591 0.252 0.253 0.632
+SEDD 0.179 0.032 0.999 4.216 2.576 0.211 0.200 0.609
+DINOISER 0.209 0.031 0.337 1.247 1.227 0.926 0.256 0.633
+Table 10: DialogueSum Experiment
 Appendix G Detailed analysis of Results
@@
 These results confirm that our method consistently produces outputs that adhere to the specified conditions.
-Diffusion-EAGS demonstrates the highest MAUVE score of 0.733 in Table 8-ParaDetox, and high level of text quality surpassing that of GPT-2 in Table 9 in text quality.
-ParaDetox is colloquial dataset including slang, numerous abbreviations, and various perturbations, so our model demonstrate robustness to such perturbations with 69.5 PPL.
+Diffusion-EAGS demonstrates the highest MAUVE score in Table 8-ParaDetox, and high level of text quality surpassing that of GPT-2 in Table 9 in text quality.
+ParaDetox is colloquial dataset including slang, numerous abbreviations, and various perturbations, so our model demonstrate robustness to such perturbations.
 As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 28, and 9.
 Notably, CDLMs demonstrate a noticeable deficiency in diversity.
+Examining the results of Diffuseq, it is evident that the grammar score is comparatively lower than that of other models.
+This outcome is expected, as the outputs from Diffuseq frequently display inaccurate sentence structures, including duplications of words or phrases.
+Conversely, the outputs from Dinoiser achieve moderate grammar scores but show limited diversity.
+This finding, coupled with our additional experiments concerning the beam size during Dinoiser generation, suggests that Dinoiser’s performance predominantly relies on memorization.
 In contrast, our model excels at producing significantly more diverse sequences.
@@
 G.2 Quality Recheck – LLM score & Dialogue Summarization
-Model ROUGE-1 ROUGE-2 MAUVE Ngram Emb Self-BLEU Distinct-1 Distinct-2
-Ours 0.409 0.174 0.536 4.114 2.591 0.252 0.253 0.632
-SEDD 0.179 0.032 0.999 4.216 2.576 0.211 0.200 0.609
-DINOISER 0.209 0.031 0.337 1.247 1.227 0.926 0.256 0.633
-Table 10: DialogueSum Experiment
 Model PPL MAUVE vs(ngram) VS(emb) sef-bleu distinct-1 distinct-2
@@
 Table 11: Quantitative results for the GENIE model.
-Paradetox w/ LLM-t on application models
 LLM-t
@@
 Table 12: ParaDetox Dataset Generation – LLM-t is the LLM-evaluation for measuring toxicity.
+Paradetox w/ LLM-t on application models
 Since our research primarily aims to enhance the model’s inherent capabilities, we set up baselines that revolve around (or are closely related to) noise scheduling.
@@
 We employ the gpt-4 model.
-The result is as follows:
+The result is in Table 13.
 Note that since within the prompt, the baseline model’s generations are specified prior to our model’s generation, there is a significant position bias working against our favor, as noted in Zheng et al.
@@
 Third, average scores of diversity are 2, 1, 2.65, 1, and 4.6, respectively, with standard deviations of 1.3, 0, 1.45, 0, and 0.7.
-GPT-3.5-turbo’s std is almost 0.
+GPT-3.5-turbo’s std is 0 for quality MOS and 0.83 for diversity MOS.
 Model SacreBLEU COMET XCOMET
@@
 For evaluation metrics, we utilize sacreBLEU (Post, 2018) and neural-net scores such as COMET (Rei et al., 2020) and XCOMET (Guerreiro et al., 2023).
-More details are provided in Appendix L.
+More details are provided in Appendix K.
 Table 14 shows that predicting the target sequence without leveraging a multilingual model proves to be challenging.
@@
 Limitation of Diversity on Traditional DDLMs
+We summarize the generation trends of the models presented in Table below.
+We observe that when a fine-tuned GPT-2 is tasked with strongly constrained conditional generation, it struggles to properly terminate sentences with an <eos> token.
+In particular, it shows limitations when handling semantic leaps or clearly delineated structural constraints, leading to suboptimal conditional generations.
+Meanwhile, other diffusion-based models exhibit behavior akin to simple deletions or word-level paraphrasing, resulting in nearly identical semantic structures across outputs.
+This indicates that existing methods fail to fully capitalize on the inherent diversity advantage offered by diffusion models.
+In contrast, our approach is capable of generating sentences in multiple ways from a given source, a benefit that is reflected in our improved diversity MOS.
 GPT-2 Given Source : holy shit , they blew up a real artifact this time ?
@@
 - it seems like they destroyed an artifact in this time as well - they have blown up a large artifact - they have blown up it in a museum, this time - they also destroyed artifacts at the same time
-We summarize the generation trends of the models presented in Table above.
-We observe that when a fine-tuned GPT-2 is tasked with strongly constrained conditional generation, it struggles to properly terminate sentences with an <eos> token.
-In particular, it shows limitations when handling semantic leaps or clearly delineated structural constraints, leading to suboptimal conditional generations.
-Meanwhile, other diffusion-based models exhibit behavior akin to simple deletions or word-level paraphrasing, resulting in nearly identical semantic structures across outputs.
-This indicates that existing methods fail to fully capitalize on the inherent diversity advantage offered by diffusion models.
-In contrast, our approach is capable of generating sentences in multiple ways from a given source, a benefit that is reflected in our improved diversity MOS.
+Model PPL MAUVE VS(ngram) vs(emb) self-bleu distinct-1 distinct-2
+RoBERTa 109.3 0.811 4.417 3.311 0.256 0.407 0.810
+BERT 69.5 0.773 4.755 3.659 0.126 0.475 0.834
+T5 408.1 0.378 2.256 1.666 0.750 0.415 0.773
+Table 15: Performance comparison of T5, BERT, and RoBERTa.
 Diversity Saturation on LLMs
@@
 Extending this approach remains a promising avenue for future research.
-Model PPL MAUVE VS(ngram) vs(emb) self-bleu distinct-1 distinct-2
-RoBERTa 109.3 0.811 4.417 3.311 0.256 0.407 0.810
-BERT 69.5 0.773 4.755 3.659 0.126 0.475 0.834
-T5 408.1 0.378 2.256 1.666 0.750 0.415 0.773
-Table 15: Performance comparison of T5, BERT, and RoBERTa.
 Appendix H Experimental Outputs
@@
 TARGET:
-Appendix K Details on CDLMs
-K.1 Experimental Details
-For the case of Diffuseq and Dinoiser Ye et al.
-(2024), we followed the official repositories to reproduce the results.
-Results were sampled multiple times with different seeds to evaluate the diversity.
-Some deviations are as follows.
-For max-length, we choose 64 for Paradetox, QG, and QQP, and 48 for Deontology.
-The values were chosen after examining the training set.
-As for batch size, we followed the original repositories if the parameter was provided.
-If not, the batch size was chosen using linear interpolation with the size of the training set.
-Note that unlike other benchmarks, we experimented with Diffuseq-v2 (Gong et al., 2023b) in translation task for a broader comparison with existing baselines.
-K.2 Results Interpretations
-Examining the results of Diffuseq, it is evident that the grammar score is comparatively lower than that of other models.
-This outcome is expected, as the outputs from Diffuseq frequently display inaccurate sentence structures, including duplications of words or phrases.
-Conversely, the outputs from Dinoiser achieve moderate grammar scores but show limited diversity.
-This finding, coupled with our additional experiments concerning the beam size during Dinoiser generation, suggests that Dinoiser’s performance predominantly relies on memorization.
-Appendix L Details on Translation Results
-L.1 Datasets & Observations
+Appendix K Details on Translation Results
+K.1 Datasets & Observations
 […]
@@
 This suggests that future research should address the semantic capabilities of diffusion models to help bridge this gap.
-L.2 Comparison Between Easy-First and Our Proposed Method
+K.2 Comparison Between Easy-First and Our Proposed Method
 Discrete diffusion can be said to inherit ideas from NAR inference algorithm Mask-Predict Ghazvininejad et al.
@@
 (2019b) on the original DisCo architecture implementation as baselines on translations tasks in Table 14 to further elucidate the difference through empirical results.
-L.3 Experimental Details
-NAR Transformer & CMLM We utilized the official repository to produce obtain the results, with the default architecture, optimization, and inference configurations.
+K.3 Experimental Details
+NAR Transformer & CMLM We utilize the official repository to produce obtain the results, with the default architecture, optimization, and inference configurations.
 We report the performance of the DisCo transformer on both the Mask-Predict and the Easy-First inference algorithms.
-Diffuseq-v2 For Diffuseq-v2, we employ the vocabs of mBERT and choose 128 as max length for ende translation. Other settings are identical as in Appendix K.1.
+Diffuseq-v2 For Diffuseq-v2, we employ the vocabs of mBERT and choose 128 as max length for ende translation. Other settings are identical as official repository.
 SEDD The SEDDLou et al.
-(2024) model, originally designed for open-ended text generation, was adapted in this study to facilitate conditional generation.
-To align the model’s architecture with the specific requirements of the structured dataset, several modifications were implemented in both hyperparameters and preprocessing protocols.
-Specifically, the input and output token lengths were constrained to a range of 64 to 128 tokens, ensuring a more appropriate fit to the dataset’s structural characteristics.
-Moreover, distinct special tokens were introduced to clearly differentiate between input and output sequences, thereby enhancing the model’s ability to distinguish between these components during training.
-Individual data entries were further demarcated by an EOS token to delineate discrete sequences within the training process.
+(2024) model, originally designed for open-ended text generation, is adapted in this study to facilitate conditional generation.
+To align the model’s architecture with the specific requirements of the structured dataset, several modifications are implemented in both hyperparameters and preprocessing protocols.
+Specifically, the input and output token lengths are constrained to a range of 64 to 128 tokens, ensuring a more appropriate fit to the dataset’s structural characteristics.
+Moreover, distinct special tokens are introduced to clearly differentiate between input and output sequences, thereby enhancing the model’s ability to distinguish between these components during training.
+Individual data entries are further demarcated by an EOS token to delineate discrete sequences within the training process.
 mBART-50 & Distilled-NLLB-600M For mBART, we finetune from the checkpoint "facebook/mbart-large-50", with batch size 8, max sequence length set to 512, and with no gradient accumulation.
@@
 DiffusionEAGS For our model, we adopt the denosing strategy as top1 sampling and 1 size of MBR as typical translation task focuses on BLEU and COMET rather than diversity score.
-L.4 Experimental Results
-L.4.1 NAR Transformer, DisCo
+K.4 Experimental Results
+K.4.1 NAR Transformer, DisCo
 The results indicate that the DisCo transformer performs poorly on low-resource translation tasks, where the size of the dataset is small.
@@
 Such results suggest the importance of utilizing PLMs for conditional generation tasks, especially in the case where the size of the available dataset is restricted
-To account for the relatively small train set to valid/test set ratio of the dataset used in our translation experiments, which resulted in a high percentage of <UNK> tokens in the valid/test sets, we also provide results using the dictionary of a pre-trained mBART model Liu (2020).
+To account for the relatively small train set to valid/test set ratio of the dataset used in our translation experiments, which results in a high percentage of <UNK> tokens in the valid/test sets, we also provide results using the dictionary of a pre-trained mBART model Liu (2020).
 The performance benefits slightly from this change, but still lags behind those of other models.
-L.4.2 Diffuseq-v2
+K.4.2 Diffuseq-v2
 It is notable that existing diffusion language models perform poorly on translation tasks. In this section, we introduce some observations that might aid our understanding of such behaviors.
-For Diffuseq-v2, we conducted additional experiments using the same model trained on Paradetox.
-We observed that the entropy of token prediction probabilities in the translation model was orders of magnitude higher, indicating a greater level of uncertainty in its predictions.
-Similarly, the ratio of the nearest token distance to the average distance of the top five nearest tokens was significantly larger in the translation model.
+For Diffuseq-v2, we conduct additional experiments using the same model trained on Paradetox.
+We observe that the entropy of token prediction probabilities in the translation model is orders of magnitude higher, indicating a greater level of uncertainty in its predictions.
+Similarly, the ratio of the nearest token distance to the average distance of the top five nearest tokens is significantly larger in the translation model.
 This analysis suggests that a simple rounding approach from continuous to discrete space may be insufficient for machine translation, at least in low-resource settings.